### v3

# Mamba-in-Mamba: Centralized Mamba-Cross-Scan in Tokenized Mamba Model for Hyperspectral Image Classification

Abstract
Hyperspectral image (HSI) classification is pivotal in the remote sensing (RS) field, particularly with the advancement of deep learning techniques. Sequential models, adapted from the natural language processing (NLP) field—such as Recurrent Neural Networks (RNNs) and Transformers—have been tailored to this task, offering unique viewpoints. However, they face several problems: 1) RNNs struggle with aggregating central features and are sensitive to interfering pixels; 2) Transformers require extensive computational resources and often underperform with limited HSI training samples. To address the issues associated with RNNs and Transformers, recent advances have introduced State Space Models (SSM) and Mamba, known for their lightweight and parallel scanning capabilities within linear sequence processing, providing a balance between the two. However, the application of existing Mamba-based models to HSI classification has not been fully explored, and the suitability of these models for this task remains an open question. Furthermore, the current methods for converting images into sequence data in Mamba-based models are simplistic and inefficient. In response, this study introduces the innovative Mamba-in-Mamba (MiM) architecture for HSI classification, marking the pioneering deployment of Mamba in this field. The MiM model includes: 1) A novel centralized Mamba-Cross-Scan (MCS) mechanism for transforming images into sequence data; 2) A Tokenized Mamba (T-Mamba) encoder that incorporates a Gaussian Decay Mask (GDM), a Semantic Token Learner (STL), and a Semantic Token Fuser (STF) for enhanced feature generation and concentration; and 3) A Weighted MCS Fusion (WMF) module coupled with a Multi-Scale Loss Design to improve decoding efficiency. Experimental results from four public HSI datasets with fixed and disjoint training-testing samples demonstrate that our method outperforms existing baselines and state-of-the-art approaches, highlighting its efficacy and potential in HSI applications.
Index Terms:
Hyperspectral image classification, deep learning, multiscanning, state space model, Mamba.
I Introduction
Hyperspectral imaging (HSI) is a sophisticated remote sensing (RS) technique that captures images across numerous dense and continuous spectral bands. This method spans the electromagnetic spectrum from visible to infrared wavelengths, often utilizing satellites or unmanned aerial vehicles (UAVs) [1]. Unlike traditional imaging methods, HSI provides detailed spectral information for each pixel, enabling a more precise characterization of materials [2] and objects on the Earth’s surface [3]. The comprehensive data obtained from HSI makes it invaluable in various applications including agriculture [4], environmental monitoring [5], geology [6], urban planning [7], and defense and security [8]. Given the broad utility of this technology, accurate HSI classification, which assigns a label to each pixel based on its spectral signature, is critical. This classification task is not only essential for enhancing the effectiveness of these applications but has also become a focus of research within the RS community [9].
Figure 1: Conceptual diagram of the Mamba-in-Mamba (MiM) model. Initially, four distinct types of centralized Mamba-Cross-Scan (MCS) are employed to convert a cropped HSI patch into sequential data. It is then fed into four shared Tokenized Mamba (T-Mamba) encoders designed to process sequences effectively. To adjust the impact of each MCS type, a Weighted MCS Fusion (WMF) is conducted, ensuring a balanced integration of the features extracted from each MCS type. Finally, the outputs from cascaded down-sampling-driven MiM layers are fed into a decoder for classification.
In recent years, the deployment of deep learning-based models for HSI classification has seen extensive growth [10]. These models typically employ a ‘patch-wise’ learning framework, where a local patch around each pixel is processed to assign semantic labels [11, 12]. Notably, sequential models like Recurrent Neural Networks (RNNs) [13, 14, 15, 16, 17] and Transformers [18, 19, 20, 21, 22, 23, 24, 25, 26, 27, 28], originally developed for natural language processing (NLP), have been adapted to HSI classification. They have demonstrated impressive results by utilizing many-to-one or many-to-many schemes to capture contextual features within pixels or partitioned windows of one HSI patch, resulting in a comprehensive feature representation. Despite their success, both types of models face shortcomings that warrant further investigation to enhance their practicability and efficiency.
First, RNNs process an HSI patch by flattening it into a sequence and employing a many-to-one learning framework. This approach selects the output at the last step as the representative feature for the entire patch. However, RNNs inherently favor the latter steps of the input sequence, resulting in a feature representation that is biased towards the final processed steps. This bias can adversely affect model training, particularly when using the cross-entropy criterion to compare the predicted label with the true label of the central pixel. Although some strategies use attention-based weighted summations to balance the outputs across all steps, the attention values often remain skewed towards later pixels, diminishing the influence of the central pixel. Furthermore, RNNs also struggle to handle large HSI patches effectively due to their sequential processing nature, which limits parallel computation across pixels. Larger patches convert into longer sequences, increasing the model’s exposure to interfering pixels with labels differing from the central pixel. This discrepancy can significantly degrade training performance as the model becomes more sensitive to peripheral pixel variations rather than focusing on the central context.
Second, to address the inflexibility of RNNs in HSI classification, Transformer-based methods have been introduced, utilizing parallel self-attention to enhance performance [35]. While these methods overcome some of RNNs’ limitations through parallel processing and self-attention mechanisms, they hold their own set of shortcomings: 1) The self-attention mechanism scales quadratically with the size of HSI patches, leading to significant computational overhead. This becomes especially problematic when larger patches are used to capture more spatial or contextual information around the center pixel, making the approach resource-intensive and less suited for hardware-constrained environments; 2) Transformers often have a large number of parameters, increasing the risk of overfitting, especially with the limited training data typically available in HSI classification; 3) Unlike RNNs or convolutional neural networks (CNNs) [36, 37, 38], which inherently capture recurrent features or local spatial features, Transformers do not have this inductive bias. While positional embeddings are used to introduce some spatial context, their effectiveness is limited without substantial training samples, potentially reducing the model’s ability to accurately capture spatial contexts in HSIs. Although some research [39, 40, 41, 42] has explored hybrid approaches that integrate RNNs or CNNs with Transformers to leverage complementary strengths. However, these methods may still increase the computational burden and complexity, potentially leading to suboptimal performance.
Third, the use of sequential models for image interpretation presents challenges, particularly in how images are arranged into sequences due to the non-causal nature of visual data: 1) Many approaches rely on a simple ‘Raster’ scan to flatten the image, which can result in discontinuities, or ‘jump connections,’ when the scan reaches the boundaries of the image; 2) Previous works [43, 44] have demonstrated that a continuous scan approach yields better results in HSI classification. It can maintain the highest spatial correlation by ensuring continuity in scanning across pixels. It also preserves multi-directional information inherent in the image’s direction-sensitive properties. Despite these advantages, existing one [45] often processes scans one by one in a unidirectional manner, then aggregate the results. This method does not necessarily prioritize the central pixel’s features during model training and then increase the feature redundancy and computational burden.
In response to three scenarios discussed, two critical needs emerge: 1) Alternative Base Model: There is a need for an alternative model that enhances parallel computing or incorporates attention mechanisms, offering greater flexibility than RNNs. This model should also demand fewer resources and exhibit less complexity, proving lighter and more efficient than Transformers, and 2) Enhanced Scanning Method: A refined scanning method is required that supports continuous, multi-directional processing. This method should provide image-aware cognition without being redundant, effectively capturing the spatial correlations and dynamics inherent in image data.
Recently, structured state space sequence (S4) models [46] and the Mamba architecture [47, 48, 49, 50] are brought into the spotlight. These models are adept at handling long-sequence data and can scale linearly with sequence length. The Mamba architecture, in particular, builds on S4 by adding a selective-scan mechanism that adaptively attentions on relevant inputs, showing superior computational efficiency compared to Transformers. Although the Mamba-based methods has shown promising results in some visual tasks [51, 52, 54, 53, 55, 56] (e.g. Vision Mamba (ViM), Visual Mamba (VMamba), etc), its potential for HSI classification has yet to be fully explored. Therefore, there is a critical need to develop a suitable vision-based Mamba model specifically designed for HSI classification.
To this end, in this study, we introduce a innovative approach—the Mamba-in-Mamba (MiM) architecture—specifically designed for HSI classification. This method marks the pioneering application of the ViM in this task. The MiM model comprises several innovative components: 1) Centralized Mamba-Cross-Scan (MCS): This new scanning mechanism transforms images into diverse half-directional sequences, offering a more efficient, lightweight, and rational approach, 2) Tokenized Mamba (T-Mamba) Encoder: Integrates a Gaussian Decay Mask (GDM), a Semantic Token Learner (STL), and a Semantic Token Fuser (STF) to significantly enhance representative feature generation and concentration, and 3) Weighted MCS Fusion (WMF) module: Works in conjunction with a Multi-Scale Loss Design to boost decoding efficiency and overall model performance. The conceptual framework of this study is illustrated in Fig. 1.
The contributions of this study are summarized as follows:
  • 1)
    We explore the potential applications of the improved ViM architecture for HSI classification task thereby achieving the better performances and efficiency.
  • 2)
    To fully explore the ability of ViM, we propose a novel T-Mamba encoder with GDM, STL, and STF to make it more suitable for the HSI classification.
  • 3)
    We propose centralized MCS to comprehensively and continuously scan the HSI patch into several sequences diversely, which can fully and lightly understand the HSI patch within the highlights on the center pixel.
  • 4)
    The proposal shows very competitive and even SOTA performance on four datasets with fixed and disjoint training-testing samples, demonstrating the feasibility.
The rest of the paper is organized as follows: Section II presents the preliminary knowledge of Mamba, Section III details the methodology, Section IV discusses the experimental setup and results, and Section V concludes the study.
II Preliminary
This section provides a foundational concepts associated with State Space Models (SSM) and the Mamba model, presenting the knowledge necessary to proceed this study.
II-A Structured State Space Sequence Model (S4)
The SSM-based models, represent a recent class of sequence models within the field of deep learning. They are broadly related to RNNs, CNNs, and classical state space models.
Inspired by continuous system theory [57], SSMs are commonly considered linear time-invariant systems that map a signal $\mathbf{x}(t)\in\mathbb{R}^{1\times\mathcal{D}}$ to $\mathbf{y}(t)\in\mathbb{R}^{1\times\mathcal{D}}$ through a latent state $\mathbf{h}(t)\in\mathbb{R}^{1\times\mathcal{N}}$. Here, $t$ represents the time step, $\mathcal{D}$ is the initial dimension, and $\mathcal{N}$ is the dimension of the state.
Mathematically, SSMs are typically formulated as linear ordinary differential equations (ODEs) [58], as follows:
$\mathbf{h}^{\prime}(t)$ $=\mathbf{A}\,\mathbf{h}(t)+\mathbf{B}\,\mathbf{x}(t),$ (1a)
$\mathbf{y}(t)$ $=\mathbf{C}\,\mathbf{h}(t)+\mathbf{D}\,\mathbf{x}(t),$ (1b)
where $\mathbf{A}\in\mathbb{R}^{\mathcal{N}\times\mathcal{N}}$ is the evolution parameter, $\mathbf{B}\in\mathbb{R}^{\mathcal{N}\times\mathcal{D}}$ correctly aligns dimensions for input-to-state transformation, and $\mathbf{C}\in\mathbb{R}^{\mathcal{D}\times\mathcal{N}}$ is the projection parameter. $\mathbf{D}\in\mathbb{R}^{1\times\mathcal{D}}$, which typically supports skip connections, ensures that $\mathbf{y}(t)\in\mathbb{R}^{1\times\mathcal{D}}$.
S4 is a discrete version of a continuous system, developed to address the significant challenges posed when integrating continuous systems into deep learning algorithms. A timescale parameter $\mathbf{\Delta}\in\mathbb{R}^{1\times\mathcal{D}}$ is used to transform the continuous parameters $\mathbf{A}$ and $\mathbf{B}$ into discrete parameters $\mathbf{\overline{A}}\in\mathbb{R}^{\mathcal{N}\times\mathcal{N}}$ and $\mathbf{\overline{B}}\in\mathbb{R}^{\mathcal{N}\times\mathcal{D}}$. The transformation commonly employs a zero-order hold (ZOH) method [^1] https://en.wikipedia.org/wiki/Zero-order_hold, defined as:
$\mathbf{\overline{A}}$ $=\exp(\mathbf{\Delta}\mathbf{A}),$ (2a)
$\mathbf{\overline{B}}$ $=(\mathbf{\Delta}\mathbf{A})^{-1}(\exp(\mathbf{\Delta}\mathbf{A})-\mathbf{I})\odot\mathbf{\Delta}\mathbf{B},$ (2b)
Therefore, the discretized versions of Eqs. 1a and 1b, can be reformulated not as a function-to-function mapping, but rather as a sequence-to-sequence model, as shown in Fig. 2:
$\mathbf{h}(t)$ $=\mathbf{\overline{A}}\,\mathbf{h}(t-1)+\mathbf{\overline{B}}\,\mathbf{x}(t),$ (3a)
$\mathbf{y}(t)$ $=\mathbf{C}\,\mathbf{h}(t)+\mathbf{D}\,\mathbf{x}(t).$ (3b)
Finally, the models compute the output through a global convolution operation as follows:
$\mathbf{\overline{K}}$ $=\Big{(}\mathbf{C}\,\mathbf{\overline{B}},\mathbf{C}\,\mathbf{\overline{A}}\,\mathbf{\overline{B}},\,\ldots,\,\mathbf{C}\,\mathbf{\overline{A}}^{L-1}\,\mathbf{\overline{B}}\Big{)},$ (4a)
$\mathbf{y}$ $=\mathbf{x}\star\mathbf{\overline{K}},$ (4b)
where $L$ is the length of the input sequence $\mathbf{x}$, and $\mathbf{\overline{K}}\in\mathbb{R}^{\mathcal{L}}$ is a structured SSM kernel. $\star$ denotes the convolutional operation.
The notable work [59] states that the matrices $\mathbf{A}$, $\mathbf{B}$, and $\mathbf{C}$ in SSM share a crucial attribute known as Linear Time Invariance (LTI). It means that they are constant and remain unchanged for each token generated by the SSM. In essence, regardless of the input sequence presented to the SSM, the values of $\mathbf{A}$, $\mathbf{B}$, and $\mathbf{C}$ remain fixed, resulting in a static representation that is not content-aware. Additionally, the matrices $\mathbf{A}$, $\mathbf{B}$, and $\mathbf{C}$ are independent of the input, as their dimensions $\mathcal{N}$ and $\mathcal{D}$ do not vary, further emphasizing their time-invariant nature.
Figure 2: The graphic illustration of S4 model. $\mathbf{x}(t)$ is a input feature (i.e. scalar or vector) at time step $t$. $\mathbf{h}(t)$ represents the latent state feature. $\mathbf{\overline{A}}$, $\mathbf{\overline{B}}$, $\mathbf{C}$, and $\mathbf{D}$ represents the transformation matrices.
II-B Selective Scan-Based S4 Model (S6)
To address the limitations identified in the S4 model, Mamba enhances the performance of the S4 model by introducing a selective scan mechanism as the core S4 operator, termed as S6 model. It allows the continuous parameters to vary with the input, thereby enhancing selective information processing across sequences.
It modifies the S4 model by introducing adaptive parameterization mechanisms for the matrices $\mathbf{B}$, $\mathbf{C}$, and the timescale parameter $\mathbf{\Delta}$ as follows:
$\mathbf{B}$ $={Projection}_{B}(\mathbf{x}),$ (5a)
$\mathbf{C}$ $={Projection}_{C}(\mathbf{x}),$ (5b)
$\mathbf{\Delta}$ $=\tau_{\Delta}({Parameter}+{Projection}_{\Delta}(\mathbf{x})),$ (5c)
where $\mathbf{x}\in\mathbb{R}^{\mathcal{B}\times\mathcal{L}\times\mathcal{D}}$. Here, $\mathcal{B}$ represents the batch size, $\mathcal{L}$ denotes the sequence length, and $\mathcal{D}$ is the sequence dimension. The functions ${Projection}_{B}()$ and ${Projection}_{C}()$ perform linear projections from the $\mathcal{D}$-dimensional space to the $\mathcal{N}$-dimensional space, resulting in $\mathbf{B}\in\mathbb{R}^{\mathcal{B}\times\mathcal{L}\times\mathcal{N}}$ and $\mathbf{C}\in\mathbb{R}^{\mathcal{B}\times\mathcal{L}\times\mathcal{N}}$. Additionally, ${Projection}_{\Delta}()$ transforms $\mathbf{x}$ to a size of ××$\mathbb{R}^{\mathcal{B}\times\mathcal{L}\times 1}$, which is then broadcast to the $\mathcal{D}$-dimension. $\tau_{\Delta}$, typically a softplus activation [60], adjusts the scale, resulting in a new $\mathbf{\Delta}$ of size ××$\mathbb{R}^{\mathcal{B}\times\mathcal{L}\times\mathcal{D}}$.
In a summary, Mamba enhances the S4 model by making the matrices $\mathbf{B}$, $\mathbf{C}$, and the step size $\mathbf{\Delta}$ input-dependent, incorporating the sequence length and batch size to ensure that each input token is processed with distinct matrices. Moreover, these dynamic matrices selectively filter which components to maintain or omit in the hidden state, drawing an analogy to the ‘recurrence’ mechanism in RNNs. This selective processing is implemented through a ‘parallel scan’ operation [^2] https://developer.nvidia.com/gpugems/gpugems3/part-vi-gpu-computing/chapter-39-parallel-prefix-sum-scan-cuda in CUDA, which segments the sequences, processes them iteratively, and combines them efficiently. This integration of dynamic matrices and parallel processing exemplifies how Mamba captures the fast and adaptable nature of sequence handling, thereby enhancing the model’s adaptability.
Algorithm 1 Pseudo procedure of the proposed method
1:  Input: Input HSI data $\mathbf{X}\in\mathbb{R}^{\mathcal{H}\times\mathcal{W}\times\mathcal{C}_{0}}$, ground truth $\mathbf{L}\in\mathbb{R}^{\mathcal{H}\times\mathcal{W}}$; patch size $p$; fixed training samples.
2:  Output: Predicted labels of the testing samples.
3:  Initialization: Optimizer: AdamW; Learning rate: 0.0005; Criterion: CrossEntropy; Epochs: 300; Batch size: 64.
4:  Procedure:
5:  Apply principal component analysis (PCA) on $\mathbf{X}\in\mathbb{R}^{\mathcal{H}\times\mathcal{W}\times\mathcal{C}}$ for dimension reduction, getting $\mathcal{C}_{1}$.
6:  Crop patch $\textbf{X}^{(i,\,j)}$ centered at pixel $\textbf{x}^{(i,\,j)}$ from samples.
7:  Perform depth-wise convolution on $\textbf{X}^{(i,\,j)}$ from feature dimension $\mathcal{C}_{1}$ to $\mathcal{C}_{2}$, generating local features.
8:  Perform point-wise convolution on $\textbf{X}^{(i,\,j)}$ then, serving as dynamic positional embedding.
9:  Apply four types of MCS on $\textbf{X}^{(i,\,j)}$, getting sequences, $\mathbf{S}_{1}^{(i,\,j)}$, $\mathbf{S}_{2}^{(i,\,j)}$, $\mathbf{S}_{3}^{(i,\,j)}$, and $\mathbf{S}_{4}^{(i,\,j)}$ (Eqs. 14a-19a).
10:  MiM model:
11:  for $m=1$ to $4$ in $\mathbf{S}_{m}^{(i,\,j)}$ do
12:     Split $\mathbf{S}_{m}^{(i,\,j)}$ into sub-sequences, $\mathbf{s}_{f}$ and $\mathbf{s}_{b}$ (Eq. 21).
13:     T-Mamba encoder:
14:     for Both forward pass $\mathbf{s}_{f}$ and backward pass $\mathbf{s}_{b}$ do
15:        Feed sub-sequences $\mathbf{s}_{f}$, $\mathbf{s}_{b}$ into respective Conv1d and S6 model, obtaining $\widetilde{\mathbf{s}_{f}}$ and $\widetilde{\mathbf{s}_{b}}$ (Eqs. 22a-23b).
16:        Apply GDMs to $\widetilde{\mathbf{s}_{f}}$ and $\widetilde{\mathbf{s}_{b}}$, resulting in $\widehat{\mathbf{s}_{f}}$ and $\widehat{\mathbf{s}_{b}}$ (Eqs. 24a and 24b).
17:        Merge two features $\widehat{\mathbf{s}_{f}}$ and $\widehat{\mathbf{s}_{b}}$, getting $\widehat{\mathbf{s}}$ (Eq. 25).
18:        Deploy STL on $\widehat{\mathbf{s}}$ to generate down-sampled semantic tokens $\mathbf{u}$ (Eq. 26).
19:        Apply STF on $\mathbf{u}$ and original feature $\mathbf{z}$ to generate enhanced tokens $\mathbf{\widehat{u}}$ (Eq. 27).
20:        Apply a residual connection between $\mathbf{\widehat{u}}$ and pooled $\mathbf{S}_{m}^{(i,\,j)}$, producing current output $\mathbf{o}_{m}^{(i,\,j)}$ (Eq. 28).
21:        Activate $\mathbf{o}_{m}^{(i,\,j)}$ using the Tanh function.
22:     end for
23:  end for
24:  Merge four outputs, $\mathbf{o}_{1}^{(i,\,j)}$, $\mathbf{o}_{2}^{(i,\,j)}$, $\mathbf{o}_{3}^{(i,\,j)}$, and $\mathbf{o}_{4}^{(i,\,j)}$, by WMF for weighted fusion, recorded as current scale ($p$) feature $\mathbf{O}_{p}^{(i,\,j)}$ (Eq. 49).
25:  Iterate MiM models on $\mathbf{o}_{1}^{(i,\,j)}$, $\mathbf{o}_{2}^{(i,\,j)}$, $\mathbf{o}_{3}^{(i,\,j)}$, and $\mathbf{o}_{4}^{(i,\,j)}$, until the patch size $p$ is reduced to $1$ (Eq. 52).
26:  Train the model using a decoder (MLP) and a multi-scale loss function (Eq. 53).
II-C Mamba-Based Works
As a cutting-edge foundational model, Mamba has recently gained popularity. There have been various implementations exploring its applicability to diverse tasks.
Zhu et al. [52] pioneered the ViM architecture for general vision tasks such as object detection and semantic segmentation, employing bidirectional scans for precise image patch positioning prior to processing with the Mamba encoder. Building on this framework, Chen et al. [53] enhanced the Mamba with a dynamic multi-path activation mechanism, significantly improving performance across various RS image classification datasets. Concurrently, Liu et al. [54] developed VMamba, which utilizes 2D-Selective-Scan (SS2D) to transform non-causal visual images into ordered patch sequences for applications in image classification and object detection. Further extending this methodology, Chen et al. [55] applied VMamba to RS change detection tasks, encompassing binary change detection, semantic change detection, and building damage assessment, with VMamba serving as the sequential encoder to model spatio-temporal relationships. Additionally, Zhao et al. [61] introduced RS-Mamba for dense prediction in extensive RS images, incorporating the Omni-directional Selective Scan Module (OSSM) for feature aggregation. In the HSI classification domain, Jia et al. [62] adapted the standard SSM-based model, SpectralMamba, for spectral domain processing, while Yang et al. [63] combined Mamba’s spectral processing with CNN-based spatial feature extraction for HSI classification. Moreover, Sheng et al. [64] proposed a lightweight spectral-spatial Mamba-convolution model for HSI classification, utilizing spatial unidirectional and spectral bidirectional scanning for enhanced feature alignment.
Despite these advancements, to our knowledge, the application of ViM with multi-directional scanning for HSI classification has not yet been fully developed or explored. Identifying an optimized ViM architecture for HSI remains an open question. Thus, we aim to explore this gap and propose a tailored HSI-driven ViM architecture.
Figure 3: The overall structure. It begins with a cropped HSI patch after PCA and then goes through a depth-wise convolution and a point-wise convolution for local feature generation and dynamic positional embedding. It is then divided into four types of sequences by MCS, as shown in Fig. 5. Each scan type is independently processed by a specially designed T-Mamba encoder without sharing parameters, as detailed in Fig. 6, to extract contextual features. The outputs from the four T-Mamba encoders are combined through a WMF module with dynamic weighted scan fusion. Concurrently, fused feature at current scale is recorded for final multi-scale decoding. This process is repeated, with the feature sizes being progressively down-sampled until they reach a spatial size of $1$. Finally, the accumulated multi-scale features from each iteration are compiled to feed into a decoder for model training.
III Methodology
This section firstly presents (A) methodological overview briefly. Then, it presents the detailed contents: (B) Centralized MCS, (C) T-Mamba Encoder, (D) WMF Module, (E) Multi-Scale Loss Design. The overall structure and the pseudo algorithm is shown in Fig. 3 and Algorithm 1, respectively.
III-A Methodological Overview
Our method is composed of several parts:
  1. 1.
    We introduce a centralized MCS mechanism that transforms an HSI patch into four types of sequences, each featuring two directional scans.
  2. 2.
    The T-Mamba encoder is engineered to efficiently manage two directional scans. It utilizes the GDM, a downsample-driven STL, and an attention-driven STF to facilitate the effective learning of representative tokens.
  3. 3.
    The MiM model processes the four types of sequences generated by the MCS across four T-Mamba encoders, setting the foundation for subsequent feature fusion.
  4. 4.
    A WMF is developed to dynamically assign different weights to the outputs from the MiM model. Meanwhile, a scale feature, derived from the fused outputs, is used to represent this phase for later decoding.
  5. 5.
    Due to the down-sampling-like operation by the MiM model, we iterate the aforementioned procedures until the spatial size of the feature reduces to 1.
  6. 6.
    The accumulated multi-scale features, are prepared for the decoder. Each feature is linked to a specific loss function to optimize performance during training.
Detailed explanations for each component are provided in the subsequent subsections.
Figure 4: The conceptual difference between (a) self-attention mechanism in Transformers, and (b) proposed MCS mechanism with T-Mamba in overall structure for recognizing global contextual information in the patch.
III-B Centralized Mamba-Cross-Scan (MCS)
The original HSI symbolized as $\mathbf{\mathbf{X}}\in\mathbb{R}^{\mathcal{H}\times\mathcal{W}\times\mathcal{C}_{0}}$ can be defined as follows:
$\centering\mathbf{\mathbf{X}}=\Big{\{}\mathbf{\mathbf{x}}^{(i,j)}\,\Big{|}\,i=0,1,...,\mathcal{H}-1,\,j=0,1,...,\mathcal{W}-1\Big{\}},\@add@centering$ (6)
where $\mathbf{\mathbf{x}}^{(i,j)}\in\mathbb{R}^{\mathcal{C}_{0}}$ signifies the spectral signature at the spatial position $(i,j)$. Consequently,
$\centering\mathbf{\mathbf{X}}=\left(\begin{matrix}\mathbf{\mathbf{x}}^{(0,0)}&\mathbf{\mathbf{x}}^{(0,1)}&\dots&\mathbf{\mathbf{x}}^{(0,\mathcal{W}-1)}\\ \mathbf{\mathbf{x}}^{(1,0)}&\mathbf{\mathbf{x}}^{(1,1)}&\dots&\mathbf{\mathbf{x}}^{(1,\mathcal{W}-1)}\\ \vdots&\vdots&\ddots&\vdots\\ \mathbf{\mathbf{x}}^{(\mathcal{H}-1,0)}&\mathbf{\mathbf{x}}^{(\mathcal{H}-1,1)}&\dots&\mathbf{\mathbf{x}}^{(\mathcal{H}-1,\mathcal{W}-1)}.\end{matrix}\right).\@add@centering$ (7)
Designed as a patch-wise framework, the input for the neural model is a cropped patch from the original HSI. For one cropped HSI patch $\mathbf{X}^{(i,\,j)}$ centered at $\mathbf{x}^{(i,\,j)}$ with patch size $p$, an odd number, the patch can be represented as:
$\mathbf{X}^{(i,\,j)}=\Big{\{}\,\mathbf{x}^{(i+\alpha,\,j+\beta)}\,\Big{|}\,\alpha,\beta=-\frac{p-1}{2},\ldots,\frac{p-1}{2}\Big{\}},$ (8)
where $\alpha,\beta\in\mathbb{Z}$ and $i+\alpha$ and $j+\beta$ record the position.
Taking $p=5$ as an example, the cropped HSI patch $\mathbf{X}^{(i,\,j)}$ can be represented as:
$\centering\mathbf{X}^{(i,\,j)}=\left(\begin{matrix}\mathbf{x}^{(i-2,\,j-2)}&\mathbf{x}^{(i-2,\,j-1)}&\dots&\mathbf{x}^{(i-2,\,j+2)}\\ \mathbf{x}^{(i-1,\,j-2)}&\mathbf{x}^{(i-1,\,j-1)}&\dots&\mathbf{x}^{(i-1,\,j+2)}\\ \vdots&\vdots&\ddots&\vdots\\ \mathbf{x}^{(i+2,\,j-2)}&\mathbf{x}^{(i+2,\,j-1)}&\dots&\mathbf{x}^{(i+2,\,j+2)}\end{matrix}\right),\@add@centering$ (9)
where $32\leq i\leq\mathcal{H}-$ 3, $32\leq j\leq\mathcal{W}-$ 3, and $\mathbf{X}^{(i,\,j)}\in\mathbb{R}^{p\times p\times\mathcal{C}_{0}}$. The center pixel $\mathbf{X}^{(i,\,j)}[2,2]$ is $\mathbf{x}^{(i,\,j)}$.
The existing centralized scanning method, SS2D, as utilized in [54], employs traditional Raster flattening across the spatial domain of images. This approach can lead to misalignment of image contents, particularly through jump connections that occur when the scan meets image boundaries. Furthermore, SS2D does not fully embrace the two-dimensional structure of the image, limiting its ability to analyze from multiple perspectives.
Consequently, we have opted for the Snake scan method (e.g., U-Turn scan), which has been endorsed in various studies [43, 44] and serves as the foundation for our proposed MCS methodology. The Snake scan method preserves spatial continuity between pixels or partitioned windows when transforming them into sequences, reducing the incidence of jump connections at image boundaries. This method’s effectiveness is further supported by the multi-scanning strategy concept [45], which demonstrates that using sequences scanned from multiple directions can enhance sequential model training.
Leveraging this background, we have developed four types of centralized MCS that optimize the transformation of images into sequences from various starting and ending points, ensuring both potential maximization and system efficiency. The fundamental distinction between the Transformer’s self-attention and our MCS mechanism is illustrated in Fig. 4.
As depicted in Fig. 5, each MCS type initiates from two opposing positions and converges at the center pixel. This method is reminiscent of the bi-directional processing typical of RNNs but incorporates a centralized design, providing a novel approach to image sequence processing.
Before feeding the $\mathbf{X}^{(i,\,j)}$ into MCS, we first deploy the depth-wise convolution (DWconv) [^3] https://github.com/rosinality/depthwise-conv-pytorch and point-wise convolution (PWconv) [^4] https://github.com/anshulpaigwar/Pointwise-Convolutional-Neural-Network on $\mathbf{X}^{(i,\,j)}$ for local feature aggregation and dynamic positional embedding.
$\mathbf{X}^{(i,\,j)}\leftarrow PWconv\Big{(}DWconv\Big{(}\mathbf{X}^{(i,\,j)}\Big{)}\Big{)}$ (10)
III-B1 MCS Type-1
In details, as shown in Fig. 5(a), the type-1 MCS starts from the two positions, $\mathbf{x}^{(i-2,\,j-2)}$ and $\mathbf{x}^{(i+2,\,j+2)}$. Instead of directly preparing two directional sub-sequences, we firstly flatten the whole patch into a complete sequence by snake scanning, starting from $\mathbf{x}^{(i-2,\,j-2)}$, as:
$\begin{array}[]{ccccccccc}&\overbrace{{\color[rgb]{0,1,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,1,0}\mathbf{\mathbf{x}}^{(i-2,j-2)}}}^{Start}&\rightarrow&\mathbf{\mathbf{x}}^{(i-2,j-1)}&\rightarrow&\cdots&\rightarrow\mathbf{\mathbf{x}}^{(i-2,j+2)}\\ &&&&&&\downarrow\\ &\mathbf{\mathbf{x}}^{(i-1,j-2)}&\leftarrow&\mathbf{\mathbf{x}}^{(i-1,j-1)}&\leftarrow&\cdots&\leftarrow\mathbf{\mathbf{x}}^{(i-1,j+2)}\\ &\downarrow&&&&&\\ &\mathbf{\mathbf{x}}^{(i,j-2)}&\rightarrow&\mathbf{\mathbf{x}}^{(i,j-1)}&\rightarrow&\cdots&\rightarrow\mathbf{\mathbf{x}}^{(i,j+2)}\\ &&&&&&\downarrow\\ &\mathbf{\mathbf{x}}^{(i+1,j-2)}&\leftarrow&\mathbf{\mathbf{x}}^{(i+1,j-1)}&\leftarrow&\cdots&\leftarrow\mathbf{\mathbf{x}}^{(i+1,j+2)}\\ &\downarrow&&&&&\\ &\mathbf{\mathbf{x}}^{(i+2,j-2)}&\rightarrow&\mathbf{\mathbf{x}}^{(i+2,j-1)}&\rightarrow&\cdots&\rightarrow\underbrace{{\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}\mathbf{\mathbf{x}}^{(i+2,j+2)}}}_{End}\\ \end{array}$ (11)
Therefore, the generated complete sequence $\mathbf{S}_{1}^{(i,\,j)}\in\mathbb{R}^{p^{2}\times C}$ can be represented as:
$\begin{split}\mathbf{S}^{(i,j)}_{1}=&[{\color[rgb]{0,1,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,1,0}\mathbf{x}^{(i-2,j-2)}},\mathbf{x}^{(i-2,j-1)},\cdots,\mathbf{x}^{(i-2,j+2)},\cdots\\ &\cdots,\mathbf{x}^{(i+2,j-2)},\mathbf{x}^{(i+2,j-1)},\cdots,{\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}\mathbf{x}^{(i+2,j+2)}}]^{\top},\end{split}$ (12)
where each element of the $\mathbf{S}_{1}^{(i,\,j)}$ is a spectral vector, in which the $\mathbf{S}_{1}^{(i,\,j)}[\frac{p^{2}-1}{2}]$ is the center step $\mathbf{x}^{(i,j)}$ of the sequence.
Subsequently, we split the $\mathbf{S}_{1}^{(i,\,j)}$ into two sub-sequences, $\mathbf{S}_{1,\,f}^{(i,\,j)}$ and $\mathbf{S}_{1,\,b}^{(i,\,j)}$. In this case, the $\mathbf{S}_{1,\,f}^{(i,\,j)}$ starts from $\mathbf{x}^{(i-2,\,j-2)}$, and ends at $\mathbf{x}^{(i,\,j)}$; meanwhile, the $\mathbf{S}_{1,\,b}^{(i,\,j)}$ starts from $\mathbf{x}^{(i+2,\,j+2)}$, and ends at $\mathbf{x}^{(i,\,j)}$, as follows:
$\mathbf{S}^{(i,j)}_{1,\,f},\,\mathbf{S}^{(i,j)}_{1,\,b}\leftarrow{Split}\Big{(}\mathbf{S}_{1}^{(i,\,j)}\Big{)},$ (13)
where $\mathbf{S}_{1,\,f}^{(i,\,j)}$ and $\mathbf{S}_{1,\,b}^{(i,\,j)}$ are arranged as:
$\mathbf{S}^{(i,j)}_{1,\,f}$ $=\Big{[}\underbrace{{\color[rgb]{0,1,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,1,0}\mathbf{x}^{(i-2,j-2)}},\mathbf{x}^{(i-1,j-2)},\mathbf{x}^{(i,j-2)},\cdots,{\color[rgb]{1,.5,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,.5,0}\mathbf{x}^{(i,j)}}}_{{length}=(p^{2}+1)/2}\Big{]}^{\top},$ (14a)
$\mathbf{S}^{(i,j)}_{1,\,b}$ $=\Big{[}\underbrace{{\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}\mathbf{x}^{(i+2,j+2)}},\mathbf{x}^{(i+1,j+2)},\mathbf{x}^{(i,j+2)},\cdots,{\color[rgb]{1,.5,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,.5,0}\mathbf{x}^{(i,j)}}}_{{length}=(p^{2}+1)/2}\Big{]}^{\top},$ (14b)
where two sub-sequences termed as forward-direction and backward-direction, both include the center step $\mathbf{x}^{(i,j)}$ of the complete sequence and assign it at the last step in the sub-sequence. In this case, when patch size $p$ equals 5, the length of sub-sequence is 13, as shown in Fig. 5(a)
Consequently, take $\mathbf{S}^{(i,j)}_{1,\,f}$ as an example, we can feed it into Mamba model respectively, referring Eqs. 3a and 3b as:
$\mathbf{h}^{(i,j)}_{1,\,f}[t]$ $=\mathbf{\overline{A}}\,\mathbf{h}^{(i,j)}_{1,\,f}[t-1]+\mathbf{\overline{B}}\,\mathbf{S}^{(i,j)}_{1,\,f}[t],$ (15a)
$\mathbf{y}^{(i,j)}_{1,\,f}[t]$ $=\mathbf{C}\,\mathbf{h}^{(i,j)}_{1,\,f}[t]+\mathbf{D}\,\mathbf{S}^{(i,j)}_{1,\,f}[t],$ (15b)
where $t\in\mathbb{Z}$, and $10\leq t\leq\frac{p^{2}+1}{2}-$ 1, denoting the element index (step) in the sub-sequence. Therefore, the last step, $\mathbf{x}^{(i,j)}$, center of the patch, can iterate all previous steps’ features. By doing the same operation for the another sub-sequence, $\mathbf{S}^{(i,j)}_{1,\,b}$, the whole image patch’s feature can be merged by two of them.
$\mathbf{y}^{(i,j)}_{1}={Merge}\Big{(}\mathbf{y}^{(i,j)}_{1,\,f},\,\mathbf{y}^{(i,j)}_{1,\,b}\Big{)},$ (16)
where $\mathbf{y}^{(i,j)}_{1,\,f}$ and $\mathbf{y}^{(i,j)}_{1,\,b}$ are the output sequence for two sub-sequences. And $\mathbf{y}^{(i,j)}_{1}$ is the merged feature from type-1 MCS for representing whole image feature. The ${Merge}()$ operation will be detailed in the next section.
Figure 5: Four types of MCS with a $5\times 5$ patch. (a) Type-1, (b) Type-2, (c) Type-3, and (d) Type-4. They cover the all the possible continuous snake scanning for a patch.
Figure 6: The T-Mamba encoder. Take inputs of size $\mathbb{R}^{(5^{2})\times C_{2}}$ as an example. Initially, the input is split into forward and backward sequences $\mathbf{s}_{f}$ and $\mathbf{s}_{b}$ by MCS, each processed independently through a Conv1d and an S6 model. A Gaussian decay mask then statically adjusts the features of each sequence. Two processed feature sequences are then merged for feature aggregation. A downsample-driven Semantic Token Learner (STL) subsequently generates representative semantic tokens, which are correlated with $\mathbf{z}$ by an attention-driven Semantic Token Fuser (STF) to implement feature enhancement. Finally, a residual connection from the original input is used to enhance feature integration for output to the next stage.
III-B2 MCS Type-2
Hence, following the same description of type-1 MCS, the type-2 MCS with vertical manner, referring Fig. 5(b), can be formulated as:
$\mathbf{S}^{(i,j)}_{2,\,f}$ $=\Big{[}{\color[rgb]{0,1,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,1,0}\mathbf{x}^{(i-2,j-2)}},\mathbf{x}^{(i-2,j-1)},\mathbf{x}^{(i-2,j)},\cdots,{\color[rgb]{1,.5,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,.5,0}\mathbf{x}^{(i,j)}}\Big{]}^{\top},$ (17a)
$\mathbf{S}^{(i,j)}_{2,\,b}$ $=\Big{[}{\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}\mathbf{x}^{(i+2,j+2)}},\mathbf{x}^{(i+2,j+1)},\mathbf{x}^{(i+2,j)},\cdots,{\color[rgb]{1,.5,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,.5,0}\mathbf{x}^{(i,j)}}\Big{]}^{\top}.$ (17b)
In this case, the two sub-sequences $\mathbf{S}^{(i,j)}_{2,\,f}$ and $\mathbf{S}^{(i,j)}_{2,\,b}$ starts at $\mathbf{x}^{(i-2,\,j-2)}$ and $\mathbf{x}^{(i+2,\,j+2)}$, respectively and ends at $\mathbf{x}^{(i,j)}$ both with a vertical scanning manner. The length of them is same with the previous description.
III-B3 MCS Type-3
Different from the type-1 and type-2 which starts from $\mathbf{x}^{(i-2,\,j-2)}$ and $\mathbf{x}^{(i+2,\,j+2)}$ positions, the type-3 cross Mamba scan is starting from $\mathbf{x}^{(i-2,\,j+2)}$ and $\mathbf{x}^{(i-2,\,j+2)}$ with a horizontal manner, shown in Fig. 5(c). The formulated two sub-sequences are listed below:
$\mathbf{S}^{(i,j)}_{3,\,f}$ $=\Big{[}{\color[rgb]{0,1,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,1,0}\mathbf{x}^{(i-2,j+2)}},\mathbf{x}^{(i-1,j+2)},\mathbf{x}^{(i,j+2)},\cdots,{\color[rgb]{1,.5,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,.5,0}\mathbf{x}^{(i,j)}}\Big{]}^{\top},$ (18a)
$\mathbf{S}^{(i,j)}_{3,\,b}$ $=\Big{[}{\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}\mathbf{x}^{(i+2,j-2)}},\mathbf{x}^{(i+1,j-2)},\mathbf{x}^{(i,j-2)},\cdots,{\color[rgb]{1,.5,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,.5,0}\mathbf{x}^{(i,j)}}\Big{]}^{\top}.$ (18b)
III-B4 MCS Type-4
Similar with the type-3 MCS, the type-4 MCS starts from $\mathbf{x}^{(i-2,\,j-2)}$ and $\mathbf{x}^{(i+2,\,j+2)}$ positions with a vertical manner, shown in Fig. 5(d). Two sub-sequences are formulated below:
$\mathbf{S}^{(i,j)}_{4,\,f}$ $=\Big{[}{\color[rgb]{0,1,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,1,0}\mathbf{x}^{(i-2,j+2)}},\mathbf{x}^{(i-2,j+1)},\mathbf{x}^{(i-2,j)},\cdots,{\color[rgb]{1,.5,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,.5,0}\mathbf{x}^{(i,j)}}\Big{]}^{\top},$ (19a)
$\mathbf{S}^{(i,j)}_{4,\,b}$ $=\Big{[}{\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}\mathbf{x}^{(i+2,j+2)}},\mathbf{x}^{(i+2,j+1)},\mathbf{x}^{(i+2,j)},\cdots,{\color[rgb]{1,.5,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,.5,0}\mathbf{x}^{(i,j)}}\Big{]}^{\top},$ (19b)
Consequently, we have generate four types of MCS with totally eight sub-sequences. Each type of them will be input for the following proposed T-Mamba encoder with bi-directional design for feature generation.
Different from the other works, our proposed MCS is designed as a centralized manner with the snake scanning method, which is more suitable for patch-wise HSI classifier where the center pixel’s label is assigned for the patch for training. Moreover, the lightweight design of using two sub-sequences for representing whole image patch can largely save computational burden and complexity than other works which use two complete sequences for whole image interpretation.
III-C Tokenized Mamba (T-Mamba)
After obtaining two sub-sequences from each type of MCS, we introduce the novel T-Mamba encoder, which processes them in a bi-directional manner. An overview of the T-Mamba is shown in Fig. 6.
III-C1 Whole Flowchart of T-Mamba
Taking Eq. 12 as an example, for a type-1 MCS, we start with an embedded sequence $\mathbf{S}_{1}^{(i,\,j)}\in\mathbb{R}^{5^{2}\times\mathcal{C}_{2}}$. This sequence is first normalized and then projected into a sequence as follows:
$\mathbf{z}={Linear}\Big{(}{Norm}\Big{(}\mathbf{S}_{1}^{(i,\,j)}\Big{)}\Big{)},$ (20)
where $\mathbf{z}$ is of size $\mathbb{R}^{5^{2}\times\mathcal{C}_{3}}$.
Next, $\mathbf{S}_{1}^{(i,\,j)}$ is split into two directional sub-sequences, $\mathbf{s}_{f}$ and $\mathbf{s}_{b}$ by MCS, where $f$ denotes the forward sub-sequence, and $b$ denotes the backward sub-sequence with each sub-sequence having dimensions $\mathbb{R}^{13\times\mathcal{C}_{3}}$, referring Eq. 13:
$\mathbf{s}_{f},\,\mathbf{s}_{b}\leftarrow{MCS_{1}}\Big{(}\mathbf{S}_{1}^{(i,\,j)}\Big{)},$ (21)
where $MCS_{1}()$ denotes the split operation under type-1 MCS.
Both $\mathbf{s}_{f}$ and $\mathbf{s}_{b}$ are then processed through a $SiLU()$ activation function [^5] https://pytorch.org/docs/stable/generated/torch.nn.SiLU.html and a 1D convolution to prepare the features:
$\overline{\mathbf{s}_{f}}$ $={SiLU}\Big{(}{Conv1d}_{f}\Big{(}\mathbf{s}_{f}\Big{)}\Big{)},$ (22a)
$\overline{\mathbf{s}_{b}}$ $={SiLU}\Big{(}{Conv1d}_{b}\Big{(}\mathbf{s}_{b}\Big{)}\Big{)}.$ (22b)
Subsequently, these processed sub-sequences are input to the Mamba model as described in Section. II:
$\widetilde{\mathbf{s}_{f}}$ $={Mamba}_{f}\Big{(}\overline{\mathbf{s}_{f}}\Big{)},$ (23a)
$\widetilde{\mathbf{s}_{b}}$ $={Mamba}_{b}\Big{(}\overline{\mathbf{s}_{b}}\Big{)}.$ (23b)
A Gaussian decay mask (GDM) is then applied to softly select features from the outputs of the Mamba models:
$\widehat{\mathbf{s}_{f}}$ $=\widetilde{\mathbf{s}_{f}}\otimes{GDM}_{f}\Big{(}\widetilde{\mathbf{s}_{f}}\Big{)},$ (24a)
$\widehat{\mathbf{s}_{b}}$ $=\widetilde{\mathbf{s}_{b}}\otimes{GDM}_{b}\Big{(}\widetilde{\mathbf{s}_{b}}\Big{)},$ (24b)
where ${GDM}()$ applies the Gaussian decay mask and will be detailed later, and $\otimes$ denotes the Hadamard multiplication.
The resultant features, $\widehat{\mathbf{s}_{f}}$ and $\widehat{\mathbf{s}_{b}}$, are merged to form a complete feature sequence:
$\widehat{\mathbf{s}}={Merge}\Big{(}\widehat{\mathbf{s}_{f}},\,\widehat{\mathbf{s}_{b}}\Big{)},$ (25)
where $\widehat{\mathbf{s}}\in\mathbb{R}^{5^{2}\times C_{3}}$ and the ${Merge}()$ operation details will be provided later.
The merged sequence $\widehat{\mathbf{s}}$ is then processed by a Semantic Token Learner (STL) module to automatically learn representative semantic tokens and perform down-sampling:
$\mathbf{u}={STL}\left(\widehat{\mathbf{s}}\right),$ (26)
where $\mathbf{u}\in\mathbb{R}^{3^{2}\times C_{3}}$. It is crucial to maintain an odd spatial patch size to keep a center pixel.
These down-sampled semantic tokens $\mathbf{u}$ are then multiplied with the previously prepared $\mathbf{z}$ for analogous attention application:
$\widehat{\mathbf{u}}=STF\Big{(}\mathbf{u},\,{AdaptivePool}\Big{(}{SiLU}\Big{(}\mathbf{z}\Big{)}\Big{)}\Big{)},$ (27)
where $\widehat{\mathbf{u}}\in\mathbb{R}^{3^{2}\times C_{3}}$. ${AdaptivePool}()$ is an adaptive pooling operation that reduces an odd patch size to a smaller odd patch size. $STF()$ represents the operation in Semantic Token Fuser (STF), detailed later.
Finally, the residual connection from the original input is added as follows:
$\mathbf{o}^{(i,\,j)}_{1}={Linear}\Big{(}\widehat{\mathbf{u}}\Big{)}+{Linear}\Big{(}{AdaptivePool}\Big{(}\mathbf{S}_{1}^{(i,\,j)}\Big{)}\Big{)}.$ (28)
These operations at type-1 MCS results in semantic tokens representing the entire image patch, as illustrated in Fig. 6.
III-C2 Gaussian Decay Mask (GDM)
To well distribute the influence of features from each step in the sequence, it is essential to apply nuanced weighting to each step’s features. Rather than relying on binary values of 0 and 1 to decide which features to retain or eliminate, we introduce the GDM as a soft mask. This GDM assigns weights based on spectral and index distances within the sequence, focusing around the central pixel.
For instance, consider a forward feature sequence $\widetilde{\mathbf{s}_{f}}$ outputted by the Mamba model. It can be represented by:
$\widetilde{\mathbf{s}_{f}}=\Big{[}\underbrace{\widetilde{\mathbf{s}_{f}}[0],\,\widetilde{\mathbf{s}_{f}}[1],\,\widetilde{\mathbf{s}_{f}}[2],\,\cdots,\,\widetilde{\mathbf{s}_{f}}[T]}_{length=(p^{2}+1)/2}\Big{]}^{\top},$ (29)
where $t\in\mathbb{Z}$, $0\leq t\leq T$, and $T=\frac{p^{2}+1}{2}-1$, indexing the elements (steps) in the sub-sequence. The last step of the sub-sequence represents the center pixel in the patch. Therefore, the spatial GDM is calculated by evaluating the index difference between each step and the last step.
Specifically, the $\sigma_{idx}$, representing the average index distance to the last index, is computed as:
$\sigma_{idx}=\frac{1}{T}\sum_{t=0}^{T}\Big{|}t-T\Big{|}$ (30)
The weights for each index $t$ in the sub-sequence are then calculated using a Gaussian distribution centered at $T$:
$w_{t}=\exp\left(-\frac{1}{2}\left(\frac{t-T}{\sigma_{idx}}\right)^{2}\right)$ (31)
These weights are normalized to ensure they sum to one, forming the final spatial GDM for the sub-sequence:
$\mathbf{W}_{idx}[t]=\frac{w_{t}}{\sum_{t=0}^{T}w_{t}},$ (32)
where index-based weights $\mathbf{W}_{idx}$ are applied to each element in the sub-sequence $\widetilde{\mathbf{s}_{f}}$.
Additionally, the feature-based GDM is calculated based on the Euclidean distance between each step (feature vector) and the last step, where the $\sigma_{spe}$, representing the average feature distance to the last step, is computed as:
$\sigma_{fea}=\frac{1}{T}\sum_{t=0}^{T}\Big{|}\Big{|}\widetilde{\mathbf{s}_{f}}[t]-\widetilde{\mathbf{s}_{f}}[T]\Big{|}\Big{|}^{2}$ (33)
The weights are then calculated by:
$v_{t}=\exp\left(-\frac{1}{2}\left(\frac{\Big{|}\Big{|}\widetilde{\mathbf{s}_{f}}[t]-\widetilde{\mathbf{s}_{f}}[T]\Big{|}\Big{|}^{2}}{\sigma_{fea}}\right)^{2}\right),$ (34)
where ||.$||.||^{2}$ represents the L2-distance (Euclidean distance).
Finally, spectral-based weights $\mathbf{W}_{fea}$ can be obtained by:
$\mathbf{W}_{fea}[t]=\frac{v_{t}}{\sum_{t=0}^{T}v_{t}}$ (35)
Consequently, two kinds of GDMs are designed to softly select features from the sequence as:
$\mathbf{W}_{idx},\mathbf{W}_{fea}\leftarrow{GDM}(\widetilde{\mathbf{s}_{f}})$ (36)
Therefore, the Eq. 24a can be rewritten as:
$\widehat{\mathbf{s}_{f}}=\widetilde{\mathbf{s}_{f}}\otimes{Norm}\Big{(}\mathbf{W}_{idx}\otimes\mathbf{W}_{fea}\Big{)}$ (37)
The same operation is applied on the $\widetilde{\mathbf{s}_{b}}$ as well.
This approach may allow for a more refined control over feature selection, enhancing the model’s ability to capture and emphasize relevant spatial and spectral information.
III-C3 Cross Scan Merge
To effectively merge features from two half-directional sequences and restore them to the original sequence length, we have designed the ${Merge}()$ function. This function is crafted to seamlessly integrate two sub-sequences, honoring their temporal dynamics and optimally preparing them for subsequent processing layers. The computational steps involved in this merging process are detailed below, as referenced in Eq. 25.
Given two sub-sequences, $\widehat{\mathbf{s}_{f}}$ and $\widehat{\mathbf{s}_{b}}$, each comprising $T$ vectors plus a final vector in $\mathbb{R}^{\mathcal{C}_{3}}$, we first partition these sequences as follows:
$\widehat{\mathbf{q}_{f}}$ $=\Big{[}\widetilde{\mathbf{s}_{f}}[0],\,\widetilde{\mathbf{s}_{f}}[1],\,\widetilde{\mathbf{s}_{f}}[2],\,\cdots,\,\widetilde{\mathbf{s}_{f}}[T-1]\Big{]}^{\top},$ (38a)
$\widehat{\mathbf{q}_{b}}$ $=\Big{[}\widetilde{\mathbf{s}_{b}}[T-1],\,\widetilde{\mathbf{s}_{b}}[T-2],\,\widetilde{\mathbf{s}_{b}}[T-3],\,\cdots,\,\widetilde{\mathbf{s}_{b}}[0]\Big{]}^{\top},$ (38b)
$\mathbf{c}$ $=\frac{\widetilde{\mathbf{s}_{f}}[T]+\widetilde{\mathbf{s}_{b}}[T]}{2},$ (38c)
where $\widehat{\mathbf{q}_{f}}\in\mathbb{R}^{T\times\mathcal{C}_{3}}$ and $\widehat{\mathbf{q}_{b}}\in\mathbb{R}^{T\times\mathcal{C}_{3}}$ represent the initial segments of the two sub-sequences, and $\mathbf{c}\in\mathbb{R}^{1\times\mathcal{C}_{3}}$ denotes the averaged feature from the final steps of both sub-sequences.
The ${Merge}()$ function then concatenates these elements:
$\widehat{\mathbf{s}}=Linear\Big{(}{Concat}\Big{(}\widehat{\mathbf{q}_{f}},\,\mathbf{c},\,\widehat{\mathbf{q}_{b}}\Big{)}\Big{)},$ (39)
where $\widehat{\mathbf{s}}$ resumes the original length of sequence, $\mathbb{R}^{p^{2}\times\mathcal{C}_{3}}$.
This approach effectively combines the two sub-sequences into a unified sequence while preserving a logical progression of features, which is particularly beneficial in models that process sequential data requiring a contextual understanding of feature evolution.
Figure 7: The visualization process of STL. The merged sequence feature is denoted by $\widehat{\mathbf{s}}$. After applying sequential attention to $\widehat{\mathbf{s}}$, the resulting features are represented as $\widehat{\mathbf{m}}$. The learnable transformation matrices $\mathbf{U}_{1}$ and $\mathbf{U}_{2}$ are utilized to extract length-level and dimension-level semantic features, denoted by $\mathbf{A}$ and $\mathbf{V}$, respectively. Finally, the down-sampled semantic tokens $\mathbf{u}$ are derived through the inner product of $\mathbf{A}$ and $\mathbf{V}$.
III-C4 Semantic Token Learner (STL)
Efficiently condensing coarse features into fine and dynamic semantic features within a sequence remains a question. Drawing inspiration from recent advancements [26, 65], we introduce the STL in our study. STL is tailored to abstract and condense sequence input into a set of representative semantic tokens, utilizing a sequential attention mechanism to enhance the focus and relevance of features. Fig. 7 depicts an example of the transformation process facilitated by the STL.
For this process, the input feature sequence is denoted as $\widehat{\mathbf{s}}\in\mathbb{R}^{p^{2}\times\mathcal{C}_{3}}$, where $p$ is the patch size, and $\mathcal{C}_{3}$ is the dimensionality.
Initially, we adopt a sequential attention module, adapted for $\widehat{\mathbf{s}}$, utilizes a 1D convolution. This module emphasizes salient features within the $\widehat{\mathbf{s}}$, allowing for a focused analysis and feature extraction without 2D spatial dimensions.
It operates as follows:
$\widehat{\mathbf{s}}_{max}$ $=max(\widehat{\mathbf{s}}),$ (40a)
$\widehat{\mathbf{s}}_{ave}$ $=mean(\widehat{\mathbf{s}}),$ (40b)
where $\widehat{\mathbf{s}}_{max}\in\mathbb{R}^{p^{2}\times 1}$ and $\widehat{\mathbf{s}}_{ave}\in\mathbb{R}^{p^{2}\times 1}$ represent the max and average pooling features along the sequence, respectively.
Then, these features are concatenated and subjected to a 1D convolution followed by a sigmoid activation $Sigmoid()$ to produce a feature weighting map:
$\mathbf{m}=Sigmoid\Big{(}Conv1d\Big{(}Concat\Big{(}\widehat{\mathbf{s}}_{max},\,\widehat{\mathbf{s}}_{ave}\Big{)}\Big{)}\Big{)},$ (41)
where $\mathbf{m}\in\mathbb{R}^{p^{2}\times 1}$.
The weighting map is then applied to the original input:
$\widehat{\mathbf{m}}=\mathbf{m}\otimes\widehat{\mathbf{s}},$ (42)
yielding $\widehat{\mathbf{m}}\in\mathbb{R}^{p^{2}\times\mathcal{C}_{3}}$.
Subsequently, semantic tokens are initialized as $\mathbf{u}\in\mathbb{R}^{p_{2}^{2}\times\mathcal{C}_{3}}$, with $p_{2}$ being an odd number smaller than $p$ to indicate a reduced number of semantic tokens. The transformation of the weighted feature sequence $\widehat{\mathbf{m}}$ into semantic tokens is performed as follows:
$\mathbf{u}=\underbrace{Softmax\left(\left(\widehat{\mathbf{m}}\odot\mathbf{U}_{1}\right)^{\top}\right)}_{\mathbf{A}}\odot\underbrace{\Big{(}\widehat{\mathbf{m}}\odot\mathbf{U}_{2}\Big{)}}_{\mathbf{V}},$ (43)
where $\mathbf{U}_{1}\in\mathbb{R}^{\mathcal{C}_{3}\times p_{2}^{2}}$ and $\mathbf{U}_{2}\in\mathbb{R}^{\mathcal{C}_{3}\times\mathcal{C}_{3}}$ are weight matrix initialized using a Gaussian distribution. $\odot$ denotes the inner-product multiplication. This configuration allows $\widehat{\mathbf{m}}$ to map into two semantic groups effectively, with $\mathbf{A}\in\mathbb{R}^{p_{2}^{2}\times p_{2}^{2}}$ and $\mathbf{V}\in\mathbb{R}^{p^{2}\times\mathcal{C}_{3}}$ representing the resultant matrices in length-level and dimension-level which are then emphasized by a $Softmax()$ function for the most relevant semantic components. Finally, $\mathbf{u}$, the semantic tokens, is derived by inner-product from $\mathbf{A}$ and $\mathbf{V}$.
This STL aids in downsampling and captures crucial high-level features for downstream tasks.
Figure 8: The visualization process of STF. It illustrates the semantic feature fusion of $\mathbf{z}$ and $\mathbf{u}$. To ensure compatibility in size with $\mathbf{u}$, $\mathbf{z}$ is first adaptively pooled into $\overline{\mathbf{z}}$. Subsequently, the pooled feature $\overline{\mathbf{z}}$ is processed through a learnable transformation matrix $\mathbf{Z}$ to produce $\widehat{\mathbf{z}}$, which captures the influence score. This allows the correlation between $\overline{\mathbf{z}}$ and $\mathbf{u}$ to be assessed through the inner product of $\widehat{\mathbf{z}}$ and the semantic token $\mathbf{u}$. Finally, to integrate the token-influenced data and the attention-modified data, $\overline{\mathbf{u}}$ and $\overline{\mathbf{m}}$ are summed to yield the final output $\widehat{\mathbf{u}}$.
III-C5 Semantic Token Fuser (STF)
To gate with the original input, previously prepared the $\mathbf{z}$ serves as the residual information for incorporating with encoded semantic features $\mathbf{u}$, as shown in Fig. 6. Instead of simply multiplying them together, we aim to fuse them by a STF operation to merge the information contained in a set of semantic tokens back into the initial sequence data, effectively fusing learned abstract representations with original sequence features. This is particularly useful in scenarios where you want to reinforce or adjust the sequence data using higher-level learned representations, which could be more informative or refined than raw features. Fig. 8 illustrates a sample of STF transformation process.
Given two features $\mathbf{u}\in\mathbb{R}^{p_{2}^{2}\times\mathcal{C}_{3}}$ and $\mathbf{z}\in\mathbb{R}^{p^{2}\times\mathcal{C}_{3}}$, we first pool the $\mathbf{z}$ into the same sequential length with $\mathbf{u}$:
$\overline{\mathbf{z}}={AdaptivePool}\Big{(}{SiLU}\Big{(}\mathbf{z}\Big{)}\Big{)}.$ (44)
After that, we compute the influence scores for each token based on the features in $\overline{\mathbf{z}}\in\mathbb{R}^{p_{2}^{2}\times\mathcal{C}_{3}}$:
$\widehat{\mathbf{z}}=Sigmoid\Big{(}\overline{\mathbf{z}}\odot\mathbf{Z}\Big{)}$ (45)
where $\widehat{\mathbf{z}}\in\mathbb{R}^{p_{2}^{2}\times p_{2}^{2}}$ and $\mathbf{Z}\in\mathbb{R}^{\mathcal{C}_{3}\times p_{2}^{2}}$ is a relevant parameter matrix. $Sigmoid()$ function ensures the scores range between 0 and 1.
Then, the tokens $\mathbf{u}$ are weighted by the influence scores $\widehat{\mathbf{z}}$ and summed over the tokens to blend the token information into the sequence:
$\overline{\mathbf{u}}=\widehat{\mathbf{z}}\odot\mathbf{u},$ (46)
where $\overline{\mathbf{u}}\in\mathbb{R}^{p_{2}^{2}\times\mathcal{C}_{3}}$, effectively spreading the token features across the sequence dimensions.
Next, the sequential attention $SeqAttn()$, mentioned in Eqs. 40a-42 is re-applied to $\overline{\mathbf{z}}$ to emphasize significant features:
$\overline{\mathbf{m}}=SeqAttn(\overline{\mathbf{z}}),$ (47)
where $\overline{\mathbf{m}}\in\mathbb{R}^{p_{2}^{2}\times\mathcal{C}_{3}}$ represents the attention-weighted sequence features.
Finally, The final sequence output integrates both the token-influenced data and the attention-modified data:
$\widehat{\mathbf{u}}=\overline{\mathbf{u}}+\overline{\mathbf{m}},$ (48)
where $\widehat{\mathbf{u}}\in\mathbb{R}^{p_{2}^{2}\times\mathcal{C}_{3}}$, providing a rich, detail-enhanced sequence for downstream tasks.
These operations ensure that the fused feature $\widehat{\mathbf{u}}$ is well-suited for downstream tasks, capturing both the detailed and global contextual information necessary for effective model performance. The addition of the attention mechanism particularly enhances the model’s ability to focus on relevant features dynamically based on the given input context.
Figure 9: Four datasets, (a) Indian Pines (IP), (b) Pavia University (PU), (c) Houston 2013 (HU2013), and (d) WHU-Hi-HongHu (HongHu) are used for experiments. The false-color images, fixed and disjoint training sets and testing sets are shown above.
TABLE I: Description of Land-Cover Types and Number of Train-Test Samples for Four Datasets
    Indian Pines (IP)     Houston 2013 (HU2013))     Pavia University (PU)     WHU-Hi-HongHu (HongHu)
 No. Color     Class Name Train Test     Class Name Train Test     Class Name Train Test     Class Name Train Test
1     Corn-n. 50 1,384     Grass-h. 198 1,053     Asphalt 548 6,304     Red-r. 25 14,016
2     Corn-m. 50 784     Grass-s. 190 1,064     Meadow 540 18,146     Road 25 3,487
3     Corn 50 184     Grass-s. 192 505     Gravel 392 1,815     Bare-s. 25 21,796
4     Grass-p. 50 447     Tree 188 1,056     Tree 524 2,912     Cotton 25 163,260
5     Grass-t. 50 697     Soil 186 1,056     Metal-s. 265 1,113     Cotton-f. 25 6,193
6     Hay-w. 50 439     Water 182 143     Bare-s. 532 4,572     Rape 25 44,532
7     Soybean-n. 50 918     Residential 196 1,072     Bitumen 375 981     Chinese-c. 25 24,078
8     Soybean-m. 50 2,418     Commercial 191 1,053     Brick 514 3,364     Pakchoi 25 4,029
9     Soybean-c. 50 564     Road 193 1,059     Shadow 231 795     Cabbage 25 10,794
10     Wheat 50 162     Highway 191 1,036         Tuber-m. 25 12,369
11     Wood 50 1,244     Railway 181 1,054         Brassica-p. 25 10,990
12     Buildings-g.-t. 50 330     Park-l-1 192 1,041         Brassica-c. 25 8,929
13     Stone-s.-t. 50 45     Park-l-2 184 285         Small-b.-c. 25 22,482
14     Alfalfa 15 39     Tennis-g. 181 247         Lactuca-s. 25 7,331
15     Grass-p.-m. 15 11     Running-t. 187 473         Celtuce 25 977
16     Oat 15 5             Film-c.-l. 25 7,237
17                 Romaine-l. 25 2,985
18                 Carrot 25 3,192
19                 White-r. 25 8,687
20                 Garlic-s. 25 3,461
21                 Broad-b. 25 1,303
22                 Tree 25 4,015
Total     695 9,671     2,832 12,197     3,921 40,002     550 386,143
Ratio     6.7% 93.3%     18.8% 81.2%     8.9% 91.1%     0.14% 99.86%
TABLE II: The Hyperparameter settings and size of notations for Four datasets.
Definition     Notation     IP     HU2013     PU     HongHu
 Image Size:     ×$\mathcal{H}\times\mathcal{W}\times\mathcal{C}_{0}$     $145\times 145\times 200$     $349\times 1905\times 144$     $610\times 340\times 103$     $940\times 475\times 270$
Dimension after PCA:     $\mathcal{C}_{1}$     60     30     30     100
Initial Patch Size:     $p\times p$     $7\times 7$     $9\times 9$     $11\times 1$ 11     $9\times 9$
Feature Dimension Setting:     $\mathcal{C}_{2},\,\mathcal{C}_{3},\,\mathcal{C}_{4}$     64     64     32     128
Number of T-Mamba Layers:         4     4     2     3
Number of MiM Layers:         4     5     6     5
Learning Rate:         0.0005     0.0005     0.001     0.001
Feature Dropout:         0.1     0.1     0.2     0.1
III-D Weighted MCS Fusion (WMF)
After obtaining four tokenized features, $\mathbf{o}^{(i,\,j)}_{1}$, $\mathbf{o}^{(i,\,j)}_{2}$, $\mathbf{o}^{(i,\,j)}_{3}$, and $\mathbf{o}^{(i,\,j)}_{4}$, from different T-Mamba encoders, we recognize that each feature set varies depending on the scan order of the sequence. This variability is critical in the context of our proposed MCS, which involves multiple scan orders, each providing a unique output representation. To effectively integrate these features, a thoughtful fusion strategy is essential, as indiscriminate merging could dilute their distinct contributions.
In response, we introduce the WMF, designed to assign attention weights to each feature set according to its relevance and importance in the overall context. The fusion process is formulated as follows:
$\mathbf{O}^{(i,\,j)}_{p}=k_{1}\mathbf{o}^{(i,\,j)}_{1}+k_{2}\mathbf{o}^{(i,\,j)}_{2}+k_{3}\mathbf{o}^{(i,\,j)}_{3}+k_{4}\mathbf{o}^{(i,\,j)}_{4},$ (49)
where $k_{o}\in\mathbb{R}^{1}$ (for $o\in\{1,2,3,4\}$) represents the weight assigned to each encoded feature, and $k_{1}+k_{2}+k_{3}+k_{4}=1$. $\mathbf{O}^{(i,\,j)}_{p}$ denotes the fused feature at scale $p$. These weights are dynamically integrated into the network architecture, allowing for optimization through learning, ensuring that each feature set contributes optimally to the final model output.
Following this feature fusion, the weighted feature $\mathbf{O}^{(i,\,j)}_{p}$ is processed through a decoder, specifically a Multi-Layer Perceptron (MLP), essential for refining and interpreting the fused features for model training:
$\mathbf{Y}^{(i,\,j)}_{p}=MLP\Big{(}Tanh\Big{(}\mathbf{O}^{(i,\,j)}_{p}\Big{)}\Big{)},$ (50)
where $Tanh()$ denotes the Hyperbolic Tangent activation function [^6] https://pytorch.org/docs/stable/generated/torch.nn.Tanh.html.
The training of the model is governed by the cross-entropy loss function, commonly used for classification tasks. This loss function measures the discrepancy between the predicted and actual labels of the training samples and is defined as:
$\mathcal{L}_{p}=-\frac{1}{N_{tr}}\sum^{N_{tr}}_{tr=1}\Big{[}\mathbf{L}_{tr}\log(\mathbf{Y}_{p})+(1-\mathbf{L}_{tr})\log(1-\mathbf{Y}_{p})\Big{]},$ (51)
where $N_{tr}$ is the total number of training samples, $\mathbf{L}_{tr}$ denotes the true label, and $\mathbf{Y}_{p}$ the predicted label for the $tr$-th sample. This function effectively minimizes the error between predicted and true labels, driving the training process towards better accuracy.
TABLE III: DIFFERENT CLASSIFICATION METHODS IN TERMS OF OA, AA, AND Kappa AS WELL AS THE ACCURACIES FOR EACH CLASS ON THE INDIAN PINES DATASET. THE BEST ONE IS HIGHLIGHTED AS RED. THE SECOND BEST IS HIGHLIGHTED AS BLUE.
No.     Color.     vanilla ViT[18]     SpeFormer[19]     MAEST[21]     HiT[22]     SSFTT[26]     3DViT[27]     HUSST[28]     ViM[52]     MiM (Ours)
 1         64.480     70.538     76.721     87.948     84.825     81.328     87.137     74.787     90.291
2         61.374     65.570     68.970     80.650     84.251     73.540     82.510     68.233     90.269
3         69.919     86.352     89.600     82.353     93.583     93.557     90.515     91.153     97.082
4         84.032     90.566     88.602     82.629     89.825     89.070     85.414     91.815     93.537
5         93.333     96.379     97.787     98.501     96.591     96.847     97.951     98.860     98.722
6         94.348     99.773     98.731     95.897     99.658     96.512     99.312     98.848     100
7         69.006     77.595     79.433     88.548     89.398     82.875     82.641     76.452     91.243
8         71.266     78.037     85.752     87.729     90.015     87.600     88.812     81.906     93.109
9         68.959     72.697     76.994     86.124     82.871     80.681     86.925     79.850     88.794
10         99.690     99.690     99.962     97.885     99.387     100     99.083     99.692     100
11         95.215     97.756     96.940     98.355     96.634     97.438     96.226     96.870     98.396
12         65.172     74.237     71.945     70.753     76.562     71.958     76.923     73.851     81.198
13         66.176     92.784     93.750     84.906     88.235     90.909     90.909     90.000     84.906
14         60.000     97.500     78.000     90.698     98.734     69.474     95.000     88.636     97.500
15         60.606     71.667     52.381     84.615     78.571     64.286     73.333     64.706     75.862
16         52.632     35.714     35.714     83.333     58.824     35.714     27.778     33.333     76.923
 OA         74.0875     79.9607     83.0731     86.0511     88.7498     85.1411     87.3333     81.9667     92.0794
AA         77.6380     82.2704     84.7051     89.5578     90.9978     87.9868     89.0293     84.8289     94.1145
Kappa         0.7079     0.7733     0.8074     0.8420     0.8715     0.8305     0.8556     0.7952     0.9096
Figure 10: The classification maps with different methods for the IP dataset. (a) vanilla ViT (OA=74.0875%). (b) SpeFormer (OA=79.9607%). (c) MAEST (OA= 83.0731%) (d) HiT (OA=86.0511%). (e) SSFTT (OA=88.7498%). (f) 3DViT (OA=85.1411%). (g) HUSST (OA=87.3333%). (h) ViM (OA= 81.9667%). (i) MiM (OA=92.0794%).
TABLE IV: DIFFERENT CLASSIFICATION METHODS IN TERMS OF OA, AA, AND Kappa AS WELL AS THE ACCURACIES FOR EACH CLASS ON THE PAVIA UNIVERSITY DATASET. THE BEST ONE IS HIGHLIGHTED AS RED. THE SECOND BEST IS HIGHLIGHTED AS BLUE.
No.     Color.     vanilla ViT[18]     SpeFormer[19]     MAEST[21]     HiT[22]     SSFTT[26]     3DViT[27]     HUSST[28]     ViM[52]     MiM (Ours)
 1         87.245     89.487     90.084     90.559     96.721     95.985     96.443     78.878     97.255
2         74.871     90.862     92.186     91.322     89.994     88.007     87.622     85.758     92.947
3         60.796     78.886     79.862     66.773     83.342     88.233     75.083     46.655     87.009
4         76.601     88.610     90.488     94.595     94.804     97.830     95.025     90.010     96.251
5         99.194     100     100     98.670     100     100     100     99.326     100
6         64.266     75.571     82.845     87.232     73.704     68.196     74.182     71.951     80.032
7         95.694     80.966     89.988     84.953     98.616     99.217     94.418     77.917     99.128
8         82.894     83.960     84.277     90.554     88.452     90.886     90.376     76.343     91.252
9         94.375     99.949     99.310     95.719     92.159     93.173     99.619     95.831     100
 OA         76.0313     86.2784     88.0433     87.9183     88.8707     86.5258     86.2483     78.1811     91.5756
AA         84.4339     89.8503     90.5291     90.7538     91.4693     91.3783     90.6566     78.5284     92.7601
Kappa         0.7086     0.8288     0.8341     0.8449     0.8608     0.8581     0.8572     0.7145     0.8983
Figure 11: The classification maps with different methods for the PU dataset. (a) vanilla ViT (OA=76.0313%). (b) SpeFormer (OA=86.2784%). (c) MAEST (OA=88.0433%) (d) HiT (OA=87.9183%). (e) SSFTT (OA=88.8707%). (f) 3DViT (OA=86.5258%). (g) HUSST (OA=86.2483%). (h) ViM (OA=78.1811%). (i) MiM (OA=91.5756%).
III-E Multi-Scale Loss Design
To account for the multi-scale features potentially present in an image patch for different objects, we propose integrating multiple scale features into the decoding process.
In our approach, each MiM model performs a down-sampling operation on the input image patch, which is initially processed by the STL in the T-Mamba encoder.
Upon obtaining four distinct features, $\mathbf{o}^{(i,\,j)}_{1}$, $\mathbf{o}^{(i,\,j)}_{2}$, $\mathbf{o}^{(i,\,j)}_{3}$, and $\mathbf{o}^{(i,\,j)}_{4}$, from the first MiM model at the initial patch scale $p$, each feature is activated and subsequently fed into the next MiM model for further feature tokenization and down-sampling, as shown in the Fig. 3.
This process allows us to capture multi-scale features in a cascaded and iterative manner, which can be analogous as:
$\mathbf{O}^{(i,\,j)}_{p_{n}}=\underbrace{MiM(MiM(MiM(\cdots MiM(\mathbf{O}^{(i,,j)}_{p}))))}_{n\,times},$ (52)
where $p_{n}<p$, recording the patch size at $n$-th $MiM()$ iteration. It must be noted that all patch sizes should be an odd number. The iteration is ending at $p_{n}=1$. For example, $p=11$, $p_{2}=9$, $p_{3}=7$, $p_{4}=5$, $p_{5}=3$, and $p_{6}=1$ by 6 times of operation under MiM model.
To effectively utilize these multi-scale features, we employ an output-level based decoding strategy, where each scale feature is independently integrated into the loss function. Consequently, as per Eq. 51, we formulate the total loss function as:
$\mathcal{L}_{total}=\frac{1}{n}(\mathcal{L}_{p}+\mathcal{L}_{p_{2}}+\mathcal{L}_{p_{3}}+\cdots+\mathcal{L}_{p_{n}}).$ (53)
The iterative process concludes when the final patch size $p_{n}$ is reduced to 1. Thus, our total loss function effectively incorporates the multi-scale features from a single image patch to enhance training effectiveness.
IV Experimental results and analysis
In this section, we first describe four well-known HSI datasets. We then outline the implementation details and the methods used for comparison. Finally, we conduct extensive experiments to evaluate the performance of the proposed method both quantitatively and qualitatively.
IV-A Datasets Description
Four HSI datasets are illustrated in Fig. 9, along with their corresponding false-color images, training samples and testing samples. The numbers of training and testing samples for each class within these datasets are listed in Table I.
The Indian Pines (IP) dataset is a mixed vegetation site with $145\times 145$ pixels and 220 spectral bands. To remove water absorption bands, the number of spectral bands was reduced to 200, resulting in 16 different land-cover classes.
The Pavia University (PU) dataset consists of $610\times 340$ pixels and initially has 115 bands. In the experiment, 12 of the noisy bands were removed, leaving 103 bands to be used. The dataset has 9 different ground cover types.
The Houston 2013 (HU2013) dataset [^7] http://www.grss-ieee.org/community/technical-committees/data-fusion was used in the 2013 IEEE GRSS Data Fusion Contest. It features 144 spectral bands and spans an area of $349\times 1905$ pixels, covering 15 different land-cover classes.
The WHU-Hi-HongHu (HongHu) dataset [^8] http://rsidea.whu.edu.cn/resource_WHUHi_sharing.htm was acquired by the UAV platform on November 20, 2017, in Honghu City, Hubei Province, China. The experimental area is a complex agricultural scene with many classes of crops, and different cultivars of the same crop are also planted in the region. It consists of $940\times 475$ pixels and initially has 270 bands. This dataset has 22 different ground cover classes.
For fair comparison, all the training and testing samples are manually selected in a fixed and disjoint manner. Details can be referred to [19] and corresponding website.
IV-B Experimental Evaluation Indicators
To quantitatively assess the effectiveness of our proposed method and compare it with other methods, we rely on four key performance indicators: overall accuracy (OA), average accuracy (AA), the kappa coefficient (kappa), and accuracy within each class. These indicators provide a comprehensive view of the classification performance, with higher values indicating superior classification results.
All experiments are conducted using Pytorch 2.3 with a Geforce RTX 3090 GPU and Ubuntu 22.04 system.
TABLE V: DIFFERENT CLASSIFICATION METHODS IN TERMS OF OA, AA, AND Kappa AS WELL AS THE ACCURACIES FOR EACH CLASS ON THE HOUSTON 2013 DATASET. THE BEST ONE IS HIGHLIGHTED AS RED. THE SECOND BEST IS HIGHLIGHTED AS BLUE.
No.     Color.     vanilla ViT[18]     SpeFormer[19]     MAEST[21]     HiT[22]     SSFTT[26]     3DViT[27]     HUSST[28]     ViM[52]     MiM (Ours)
 1         89.129     90.815     89.644     89.616     91.085     95.123     92.073     87.271     94.887
2         94.852     93.950     91.537     92.643     93.056     97.550     94.473     93.035     98.273
3         96.188     97.087     100     99.901     98.232     100     100     97.527     100
4         93.285     96.843     95.083     97.016     94.711     97.681     96.279     87.680     96.866
5         96.445     94.575     96.748     97.959     96.925     97.823     96.378     93.160     99.858
6         74.011     93.464     93.770     91.374     84.713     99.306     97.260     94.595     99.652
7         82.431     91.276     90.431     91.541     87.516     90.009     91.864     84.735     91.423
8         74.693     68.837     75.459     81.531     88.889     72.029     76.727     84.537     87.167
9         73.527     86.340     85.249     84.658     88.835     85.297     86.209     85.089     84.309
10         72.931     71.959     86.763     88.628     91.663     82.154     81.541     84.514     92.380
11         69.631     75.279     86.429     84.187     87.778     87.748     92.630     83.675     90.028
12         70.809     78.114     75.945     82.155     88.132     78.154     81.899     79.158     89.581
13         63.805     90.429     91.667     91.501     90.203     92.632     90.657     87.896     93.080
14         96.509     97.436     100     93.690     97.041     95.367     99.597     99.190     97.436
15         98.943     95.075     96.728     96.629     91.245     97.399     95.122     97.612     98.333
 OA         82.9548     86.4557     88.8415     90.0303     91.4241     89.7024     90.1205     87.5871     92.8917
AA         84.0284     88.3335     89.8294     91.7086     93.0019     90.1268     93.0002     89.1248     94.2108
Kappa         0.8151     0.8530     0.8787     0.8918     0.9069     0.8882     0.8928     0.8653     0.9228
Figure 12: The region of interest (ROI) in classification maps with different methods for the HU2013 dataset. (a) vanilla ViT (OA=82.9548%). (b) SpeFormer (OA=86.4557%). (c) MAEST (OA=88.8415%) (d) HiT (OA=90.0303%). (e) SSFTT (OA=91.4241%). (f) 3DViT (OA=89.7024%). (g) HUSST (OA=90.1205%). (h) ViM (OA=87.5871 %). (i) MiM (OA=92.8917%).
IV-C Brief Description and Settings of Compared Methods
To evaluate the effectiveness of our proposed method, we conducted comparative analyses against a range of baseline and state-of-the-art models. These included vanilla ViT (ViT) [18], SpeFormer [19], MAEST [21], HiT [22], SSFTT [26], 3DViT [27], HUSST [28], and ViM [52]. All models were employed as patch-wise classifiers, with parameter settings adhering to their respective references, especially the initial patch size for different datasets. The comparison is grounded in the same number of the training samples without data augmentation.
  • 1)
    ViT is configured with a feature dimension of 64 and incorporates 4 multi-heads self-attention. The network depth is set to 4 layers. Across all datasets, the initial size of the cropped patch is $9\times 9$. We use a sub-patch size of $3\times 3$, partitioned in a non-overlapping manner.
  • 2)
    SpeFormer integrates group-wise spectral embedding into the Transformer encoder and includes cross-layer adaptive fusion modules to boost the encoder’s learning capabilities. It is configured with an initial patch size of $7\times 7$ across all datasets in our experiments.
  • 3)
    MAEST integrates a reconstruction path with a classification path for HSI classification, employing masking auto-encoding (MAE). The initial patch size is set at $7\times 7$.
  • 4)
    HiT embeds convolution operations into the Transformer structure to capture the subtle spectral discrepancies and convey the local spatial context information. For all datasets, the initial patch size is set at $15\times 15$ for HiT.
  • 5)
    SSFTT combines 3D-CNNs and 2D-CNNs, featuring a Gaussian-weighted feature tokenizer to prepare extracted features for the Transformer encoder. In our experiments, patch sizes vary by dataset: $13\times 13$ for the IP and PU dataset and $9\times 9$ for both HU2013 and HongHu dataset.
  • 6)
    3DViT integrates spectral and spatial features into sub-cubes using a 3D configuration within the ViT framework. It implements 3D coordinate positional embedding and a local-global feature fusion mechanism, merging local details with global context. For all datasets, the initial patch size for the 3DViT model is set at $9\times 9$.
  • 7)
    HUSST implements a cascaded approach, combining a spectral Transformer with a spatial local-global Transformer hierarchically. It introduces cross-scale local-global feature fusion for multi-scale feature aggregation. The initial patch size set at $11\times 1$ 11 for all datasets.
  • 8)
    ViM, the first Mamba model tailored for vision tasks, is applied to HSI classification to assess its viability. The hyperparameters are aligned with those of our proposed MiM model, detailed later.
  • 9)
    Our MiM model standardizes all feature dimensions at 64, with T-Mamba encoders having a depth of 2. The initial patch sizes are set at $7\times 7$ for the IP dataset, $11\times 1$ 11 for the PU dataset, and $9\times 9$ for the HU2013 and HongHu dataset. Others can be found in Algorithm 1.
TABLE VI: DIFFERENT CLASSIFICATION METHODS IN TERMS OF OA, AA, AND Kappa AS WELL AS THE ACCURACIES FOR EACH CLASS ON THE WHU-Hi-HongHu DATASET. THE BEST ONE IS HIGHLIGHTED AS RED. THE SECOND BEST IS HIGHLIGHTED AS BLUE.
No.     Color.     vanilla ViT[18]     SpeFormer[19]     MAEST[21]     HiT[22]     SSFTT[26]     3DViT[27]     HUSST[28]     ViM[52]     MiM (Ours)
 1         83.308     92.414     93.150     92.397     94.062     93.913     93.392     90.774     91.922
2         51.588     77.377     69.992     83.118     80.321     84.327     82.783     78.092     83.239
3         66.129     87.049     87.753     86.075     86.447     88.706     86.066     85.647     89.772
4         81.673     92.768     95.519     95.807     92.688     93.738     94.840     90.939     95.993
5         30.789     42.216     52.979     50.280     39.060     45.475     59.836     40.896     59.971
6         80.313     93.287     94.380     94.295     93.865     93.009     94.936     92.830     96.003
7         55.159     69.221     81.213     66.949     84.649     74.366     79.445     81.871     83.873
8         29.825     48.326     45.361     47.745     46.911     64.771     45.650     54.258     64.170
9         87.712     95.937     95.455     93.624     96.133     94.108     95.912     95.570     96.399
10         50.244     78.034     85.437     78.864     82.318     85.436     83.367     77.348     87.094
11         32.714     68.521     80.407     64.692     76.228     82.055     72.825     75.274     88.261
12         38.737     60.166     72.607     64.964     76.981     65.600     69.342     68.752     77.643
13         41.373     66.599     75.962     60.533     72.911     70.993     75.059     66.550     76.611
14         53.612     75.796     88.200     85.221     85.478     90.432     86.554     74.109     89.186
15         48.922     80.839     84.730     83.587     84.911     70.855     72.665     37.689     87.631
16         85.969     96.110     94.660     94.505     95.649     95.244     96.007     93.233     95.948
17         71.222     89.171     74.695     94.118     85.529     87.956     87.207     85.676     92.016
18         49.649     84.597     83.587     85.643     83.333     84.548     86.010     89.097     86.875
19         59.138     83.529     86.927     82.922     90.572     86.385     88.418     86.205     91.872
20         34.296     71.906     80.832     32.043     79.922     80.556     82.181     74.823     88.583
21         20.711     50.048     52.946     52.932     46.456     62.122     44.932     56.133     71.989
22         32.600     66.921     80.489     87.111     68.578     83.324     67.961     53.190     83.057
 OA         65.2688     82.9934     87.3104     84.2750     85.2845     85.8091     86.9473     82.7789     89.5749
AA         67.4351     84.6977     85.7987     83.7510     87.4042     88.7126     87.9006     85.7671     90.9751
Kappa         0.5873     0.7911     0.8427     0.8048     0.8196     0.8254     0.8387     0.7897     0.8705
Figure 13: The classification maps with different methods for the HongHu dataset. (a) vanilla ViT (OA=65.2688%). (b) SpeFormer (OA=82.9934%). (c) MAEST (OA=87.3104%) (d) HiT (OA=84.2750%). (e) SSFTT (OA=85.2845%). (f) 3DViT (OA=85.8091%). (g) HUSST (OA=86.9473%). (h) ViM (OA=82.7789%). (i) MiM (OA=89.5749%).
Figure 14: Visualization of T-SNE results on the IP test samples by different methods. (a) original test samples’ T-SNE results. (b) vanilla ViT. (c) SpeFormer. (d) MAEST. (e) HiT. (f) SSFTT. (g) 3DViT. (h) HUSST. (i) ViM. (j) MiM.
IV-D Experimental Results
IV-D1 Quantitative Results and Classification Maps
Tables III, IV, V, and VI present average classification metrics under 5 times of running, such as OA, AA, Kappa coefficient, and class-wise accuracy for the IP, PU, HU2013, and HongHu datasets, respectively. Corresponding classification maps are displayed in Figs. 10, 11, 12, and 13.
From these results, it’s evident that baseline methods like vanilla ViT and ViM do not yield satisfactory outcomes, especially the vanilla ViT, which struggles due to its purely parallel and simultaneous processing approach (e.g., self-attention). This often results in noise-like classification artifacts, particularly noticeable in smooth regions and at the boundaries of land-covers, as the classification map for vanilla ViT exhibits much sharper visual artifacts. Conversely, ViM, which integrates bidirectional sequence processing with Mamba encoders, typically shows better performance. This improvement could be attributed to an ordering bias, akin to that found in RNNs, which is advantageous when training samples are limited.
SpeFormer and MAEST, enhancements over vanilla ViT, achieve significantly better results. SpeFormer introduces local feature aggregation within spatial and spectral domains and incorporates cross-layer connections to enhance the robustness of its encoder. Compared to vanilla ViT, SpeFormer produces smoother classifications in land-cover regions. MAEST integrates a masked auto-encoder into SpeFormer, adding a noise mask to the original data and reconstructing it, which enhances generative performance on unseen data, especially when training and testing samples are uncorrelated. MAEST presents better improvements on SpeFormer.
On another front, 3DViT extends ViT into a 3D sequential model across spatial and spectral domains, effectively linking spatial and spectral features to enhance classification accuracy. Despite being a pure Transformer-based model without hybrid elements like RNNs or CNNs, it shows promising results.
Moreover, HiT, SSFTT, and HUSST combine CNNs with Transformers to leverage local and global feature awareness, generally yielding superior results. These models provide clearer distinctions in flat land-cover regions and at boundaries. However, their larger model sizes contribute to higher computational demands and are less hardware-friendly.
TABLE VII: TOTAL RUNNING TIME (MINUTES) AND THE NUMBER OF PARAMETERS REQUIRED BY DIFFERENT METHODS ON THE INDIAN PINES DATASET.
vanilla ViT[18]     SpeFormer[19]     MAEST[21]     HiT[22]     SSFTT[26]     3DViT[27]     HUSST[28]     ViM[52]     MiM (Ours)
 Train (minutes.) 2.1     3.5     4.7     5.6     5.5     4.5     3.5     2.5     3.6
Test (seconds.) 2.0     4.0     5.0     11.0     5.0     8.0     3.0     7.0     5.0
Parameters 416K     644K     953K     45.8M     2.9M     2.4M     9.1M     357K     1.1M
Our method, MiM, introduces a MCS mechanism with a T-Mamba encoder, facilitating lightweight, multi-directional feature generation. Compared to ViM, our specialized T-Mamba encoder and extended bidirectional scan improve classification performance significantly. Our approach, notably lighter than HiT, SSFTT, and HUSST, produces classification maps with smoother regions and more distinct land-cover boundaries. The class-wise accuracies are competitive and generally superior to other methods.
This exploration paves the way for our future research into adapting pure Mamba models and MCS into a 3D configuration, similar to 3DViT which extends ViT into 3D version.
IV-D2 Learned Feature Visualizations by T-SNE
Deep learning-based methods excel in creating new feature representations that replace the original inputs through representation learning. Typically, these methods train a discriminative model to transform input data into a distinct output representation for tasks such as HSI classification, subsequently utilizing these new features for clustering.
To evaluate the representational capabilities of various methods, we employed t-distributed stochastic neighbor embedding (T-SNE) [^9] https://en.wikipedia.org/wiki/T-distributed_stochastic_neighbor_embedding, a tool that visualizes the distribution of learned features from test samples.
Using the IP dataset as an example, Fig. 14 illustrates the visualization outcomes of different methods. From this, two key points emerge. Firstly, while some clusters represented by different colors may appear closely positioned in T-SNE results, this proximity should not be interpreted as indicative of similarity due to T-SNE’s internal mechanics. Secondly, the presence of same-color clusters at considerable distances from each other highlights T-SNE’s focus on preserving local structure while optimizing a non-convex function, which may lead to local minima.
The visualization clearly demonstrates that our MiM model effectively discriminates between learned feature representations. Through T-SNE, the new feature representations not only cluster but also show increased discriminative properties, with several clusters unifying under the same color and appearing smoother.
IV-D3 Comparison of Computational Costs
To evaluate computational efficiency across all methods, we conducted a detailed statistical analysis, summarized in Table VII. This analysis adheres to previously described experimental setups and includes metrics such as total training time (minutes), testing time (seconds), and model size (number of parameters) for the IP dataset.
Our method showcases competitive processing times, significantly outperforming advanced models like HiT, MAEST, 3DViT, and SSFTT in terms of efficiency. However, it records slightly longer times compared to simpler models such as vanilla ViT, SpeFormer, HUSST, and ViM. When it comes to the number of trainable parameters, our approach proves to be more efficient, requiring fewer parameters than other leading methods, which underscores its parameter efficiency.
In summary, our method strikes an effective balance between running time and model complexity, measured by parameter count. While it doesn’t achieve the fastest training or testing times, it offers lower operational times than several sophisticated models and maintains a manageable parameter count. This combination of efficiency and simplicity makes our method particularly suitable for scenarios where both performance and model manageability are priorities.
IV-E Other Analysis and Ablation Study
In this part, a comprehensive analysis, experimental evaluation, and ablation studies have been presented.
TABLE VIII: IMPACT OF DIFFERENT INSERTABLE COMPONENTS IN T-MAMBA ENCODER ON THE CLASSIFICATION RESULTS (%). THE BEST ONE IS HIGHLIGHTED AS RED.
$STL$     $GDM$     $STF$ IP     PU     HU2013     HongHu
 ✗     ✗     82.559     80.999     88.112     84.123
    ✗     86.129     85.698     90.995     86.889
✗         84.006     82.159     89.459     86.114
✗     ✗     84.561     83.226     90.642     85.013
        87.021     86.111     91.324     88.546
    ✗     91.969     89.123     91.452     88.982
✗         87.159     87.661     91.003     88.002
        92.079     91.576     92.892     89.575
IV-E1 Ablation on Proposed Different Components
In the T-Mamba encoder, we integrated several key components: the STL, GDM, and STF. To assess their effectiveness, we conducted an ablation study, the results of which are summarized in Table VIII. The ablation study details how each component was tested. In scenarios where STL and STF were not used, we substituted them with general average pooling and simple summation, respectively, to gauge their relative performance.
STL outperforms average pooling in merging MCS features, suggesting its superior ability to capture contextual information within the merged features, significantly enhancing accuracy.
GDM refines feature focus after Mamba encoding by reweighting the sequence based on the proximity to the central pixel, which is the last index in a sub-sequence. This targeted approach enhances the central pixel’s prominence, reducing the impact of interfering pixels and yielding modest improvements.
STF fuses enhanced features more effectively than simple summation, leveraging higher-level learned representations to enrich the raw feature set. Incorporating STF alone leads to better outcomes than summation, underscoring its utility in strengthening feature representations.
Additionally, combining STL and STF yielded the most substantial improvements, likely due to their complementary mechanisms both utilizing sequential attention.
Overall, this ablation study demonstrates that the integration of STL, GDM, and STF significantly boosts performance by synergistically refining the handling of image band features in a sequential model, particularly with the continuous scanning features critical in our proposed MiM model.
TABLE IX: IMPACT OF DIFFERENT INITIAL PATCH SIZE FOR THE OVERALL ACCURACY (%) ON FOUR DATASETS. THE BEST ONE IS HIGHLIGHTED AS RED.
Patch Size ($p$) IP     PU     HU2013     HongHu
 $3\times 3$ 89.001 $\uparrow$     86.331 $\uparrow$     89.123 $\uparrow$     83.127 $\uparrow$
$5\times 5$ 91.332 $\uparrow$     89.126 $\uparrow$     91.028 $\uparrow$     87.269 $\uparrow$
$7\times 7$ 92.079 $\uparrow$     90.214 $\uparrow$     $9$uparrow     89.113 $\uparrow$
$9\times 9$ 91.554 $\downarrow$     91.002 $\uparrow$     92.892 $\uparrow$     89.575 $\uparrow$
$11\times 1$ 11 90.667 $\downarrow$     91.576 $\uparrow$     92.091 $\downarrow$     88.990 $\downarrow$
$13\times 13$ 89.214 $\downarrow$     90.885 $\downarrow$     90.451 $\downarrow$     87.317 $\downarrow$
$15\times 15$ 87.332 $\downarrow$     88.213 $\downarrow$     88.519 $\downarrow$     86.451 $\downarrow$
IV-E2 Analysis on Initial Patch Size
We investigated the effect of initial patch size on classification accuracy. For instance, with the IP dataset, the optimal starting patch size is 7 $\times$ 7. This setup allows us to derive four scale features—scale 7, 5, 3, and 1—each contributing to the loss function. Therefore, the final result is obtained by training four loss functions together.
As shown in Table IX, we evaluated a range of initial patch sizes (3, 5, 7, 9, 11, 13, 15). The results, specifically OA for four datasets, are detailed in the same table.
An appropriately large initial patch size can improve model performance by capturing more spatial context. However, excessively large patch sizes can be detrimental, introducing noise from interfering pixels.
For the relatively compact IP dataset, the ideal initial patch size is small, confirmed at 7 $\times$ 7. In contrast, for the HU2013 dataset, which features a dense spatial distribution but closely spaced land-covers, a slightly larger patch size of 9 $\times$ 9 is optimal. The PU dataset, characterized by its flat and smooth terrain, warrants an even larger optimal initial patch size of 11 $\times$ 11. Finally, the HongHu dataset gets the ideal initial patch size at 9 $\times$ 9.
IV-E3 Investigation of Scan Design
In the proposed centralized MCS, selecting the appropriate scan manner is essential for optimal feature allocation and positioning within sequential models. It is critical to choose a cross-scan design that effectively spans spatial domains.
Given that random and casual scanning may lead to significant redundancy and poor feature arrangement, we established a certain rule: spatially, the scanning should keep continuous as much as possible. Thus, through investigation, we identified several scanning designs, as shown in Fig. 15. For each scan design, it can be extended into four types, similar to those shown in Fig. 5.
The Mamba-scan achieves the greatest continuity when converting an image into a sequence. The Raster-scan, a common method, linearizes the image into a sequence but introduces jump connections at the boundaries of the image. The Diagonal-scan approaches this by scanning images diagonally, selecting positions along the diagonals. Conversely, the Zig-zag-scan modifies the diagonal approach by maintaining partial continuity at image boundaries.
Figure 15: Investigated scan designs. (a) proposed Mamba-scan design. (b) general Raster-scan design. (c) Diagonal-scan design. (d) Zig-Zag-scan design. Each design can be extended into four types from different start and end positions, similar to those shown in Fig. 5.
TABLE X: THE COMPARISON OF USING DIFFERENT SCAN DESIGNS FOR OA (%). THE BEST ONE IS HIGHLIGHTED AS RED. THE SECOND BEST IS HIGHLIGHTED AS BLUE.
Num. of used IP     PU     HU2013     HongHu
  1 89.123     89.559     90.123     87.315
Mamba-scan 2 90.156     90.452     91.029     88.569
3 91.221     91.001     91.993     89.226
4 92.079     91.576     92.892     89.575
  1 88.326     87.009     89.163     85.429
Raster-scan 2 89.555     88.322     90.442     86.129
3 90.122     89.125     91.105     88.226
4 91.612     90.761     91.986     89.125
  1 88.039     86.002     89.013     83.664
Diagonal-scan 2 89.196     87.569     89.669     86.548
3 90.236     88.555     90.139     87.841
4 91.002     89.571     91.004     88.221
  1 88.129     87.563     87.125     84.003
Zig-Zag-scan 2 90.009     87.993     89.699     86.443
3 91.011     88.961     90.861     87.231
4 91.536     91.299     91.834     87.992
We tested these scanning methods to evaluate their performance, with results summarized in Table X. Each scan design was tested using one to four types. The most effective results were achieved using the Mamba-scan with four directional types, confirming that a continuous sequence arrangement enhances content recognition. This finding is further supported by comparing the Diagonal-scan and the Zig-zag-scan, where the latter showed slight improvements due to its semi-continuous positioning. The commonly used Raster-scan, despite its prevalence in image transformation tasks, tends to disrupt spatial continuity, particularly when spatial context is significant in RS land-cover analysis, resulting in inferior performance compared to the Mamba-scan.
Additionally, employing four types of scans enhances each scanning approach, demonstrating the viability of multi-directional feature learning.
IV-E4 Exploring the Feature Fusion Strategy in T-Mamba
We developed our bi-directional feature fusion strategy in T-Mamba encoder with two ways, as shown in Fig. 16: 1) Post-Merge strategy, and 2) Pre-Merge strategy. To evaluate the effectiveness of these two strategies, we conducted experiments across four datasets. The results of these experiments are detailed in Table XI.
TABLE XI: IMPACT OF DIFFERENT FEATURE FUSION STRATEGY IN T-MAMBA ENCODER FOR THE OA (%) ON FOUR DATASETS. THE BEST ONE IS HIGHLIGHTED AS RED.
IP     PU     HU2013     HongHu
 Post-Merge 91.326     90.332     91.229     88.450
Pre-Merge 92.079     91.576     92.892     89.575
Figure 16: Designed bi-directional feature fusion manner in T-Mamba encoder. (a) Post-Merge strategy, and (b) Pre-Merge strategy.
We observed that the Pre-Merge strategy tends to be more effective for feature condensation and aggregation. This is likely because the STL module is applied to the merged features, which encapsulate the entirety of the image’s features. Consequently, the STL can comprehensively analyze the whole image’s contents, facilitating the condensation and abstraction of features, akin to a ’global’ awareness of the image. Subsequently, the STF module leverages the pooled feature of the entire image, $\mathbf{z}$, to achieve a holistic understanding of image features. In contrast, the Post-Merge strategy, applies the STL and STF to each branch independently without integrating semantic information across the branches. This approach results in less effective feature integration compared to the Pre-Merge strategy.
Therefore, rather than directly adopting the ViM model, separating two individual branches, for our HSI classification task, we have developed a more suitable and innovative approach that enhances performance through strategic feature merging and processing.
Figure 17: The comparison between using different sequential networks (i.e., RNN, GRU, LSTM, Transformer encoder, RT, Mamba, and T-Mamba). We have to note that all networks are not sharing parameters for four types of MCS in MiM model. Experimental results are listed in (a) IP dataset, (b) PU dataset, (c) HU2013 dataset, and (d) HongHu dataset. Meanwhile, the (e) shows the training time, OA(%) and model size by using those networks.
IV-E5 Comparison with Other Sequential Networks
Following the development of the T-Mamba, we conducted comparative experiments against other established sequential networks: RNN [^10] https://pytorch.org/docs/stable/generated/torch.nn.RNN.html, GRU [^11] https://pytorch.org/docs/stable/generated/torch.nn.GRU.html, LSTM [^12] https://pytorch.org/docs/stable/generated/torch.nn.LSTM.html, general Transformer encoder [^13] https://pytorch.org/docs/stable/generated/torch.nn.TransformerEncoder.html, RT (a hybrid of LSTM and Transformer) [45], and a general Mamba model [47]. These models all process vector inputs at each timestep, allowing us to integrate them directly into our system for a straightforward comparison. We have to note that all networks are not sharing parameters for four types of MCS in MiM model. All tests were performed under identical conditions, with the results shown in Fig. 17.
The vanilla RNN displayed the lowest performance, constrained by its basic approach to processing sequential data. Both GRU and LSTM showed similar results, but the GRU was more efficient due to its simpler structure with one fewer gate than the LSTM. The RT model, which merges LSTM’s sequential processing with the Transformer’s self-attention, exhibited enhanced performance, underscoring the benefits of combining these two approaches. Although RT outperformed the standalone LSTM and GRU, it requires a significantly larger parameter set.
Our newly designed T-Mamba, which enhances the basic Mamba encoder with tokenization capabilities, does add some parameters but achieves considerable improvements in classification performance. In comparison to the other models, T-Mamba strikes an optimal balance between model size and efficacy. It surpasses the general Transformer encoder in efficiency, offering a more lightweight and robust solution. While RT provides competitive performance, its larger model size and higher resource consumption are less than ideal, making T-Mamba the more advantageous choice in terms of efficiency and effectiveness.
TABLE XII: DIFFERENT CLASSIFICATION METHODS IN TERMS OF OA ON THE INDIAN PINES DATASET WITH DATA AUGMENTATION. THE BEST ONE IS HIGHLIGHTED AS RED. THE SECOND BEST IS HIGHLIGHTED AS BLUE.
Num. Train vanilla ViT[18]     SpeFormer[19]     MAEST[21]     HiT[22]     SSFTT[26]     3DViT[27]     HUSST[28]     ViM[52]     MiM
 Scena. 1 1024 85.163     94.123     97.323     97.211     97.002     97.328     96.892     94.317     97.332
Scena. 2 4170 90.669     96.128     99.002     98.386     98.991     98.001     97.669     95.659     98.239
IV-E6 Further Discussion on Mamba and Transformer under Data Augmentation
Mamba was introduced as an alternative to the Transformer for efficient and lightweight sequential feature encoding. However, unlike the Transformer, which employs global attention, Mamba processes sequences in a linear-like fashion. The Transformer’s self-attention mechanism has gained widespread adoption across various applications due to its ability to generalize well with sufficient data, making extensive training datasets a necessity for optimal performance. Against this backdrop, we conducted further experiments to evaluate Mamba-based methods against Transformer-based approaches.
These experiments centered on the augmentation of training samples. We established two scenarios: 1) using a general sample of 10% randomly selected from the ground truth, a common practice in HSI classification, and 2) augmenting the data through geometric transformations—45°, 90°, and 135° rotations, and vertical and horizontal flips—as outlined in [63].
Taking the IP dataset as an example, where labeled pixels across land-cover categories are typically imbalanced and scarce, we initially selected 1024 training samples in the scenario-1, significantly more than in previous experiments with disjoint and fixed training sets. For the augmented scenario-2, we applied geometric transformations to disjoint samples, yielding a substantially increased pool of 4170 training samples (695 multiplied by 6).
The results, detailed in Table XII, indicate that increasing the training sample size to nearly double improves the performance of all models, including our Mamba-based model, although the accuracy gap between methods narrows considerably. This is further evidenced by the shrinking difference in performance between SpeFormer and ViM, two purely model-driven approaches.
However, when the number of training samples are expanded six-fold to 4170, our MiM model was no longer the top performer. Instead, MAEST, which integrates the MAE with prior masks, achieved the best results. Transformer-based models generally performed well. This improvement is likely due to the Transformer’s mechanism, which effectively utilizes ample training samples to enhance its recognition and generalization capabilities, as supported by literature suggesting that Transformers can outperform CNNs when sufficient data is available.
Thus, while our MiM model excels in scenarios with limited, disjoint, and fixed samples—conditions that mimic real-world HSI applications. The Transformer-based methods tend to perform better under conditions of abundant training data.
V Conclusion
In this study, we introduced a novel and pioneering approach to HSI classification utilizing the newly developed Mamba architecture. Specifically, we implemented a centralized MCS for multi-directional feature aggregation, creating a lighter and more streamlined framework for visual cognition alongside the Mamba sequential model. To further enhance this architecture, we designed a T-Mamba encoder, incorporating bi-directional GDM and a STL to refine feature quality and support downsampling-driven semantic token representation. Meanwhile, porposed STF is effectively fusing learned abstract representations with original sequence features. Additionally, we integrated a WMF module and a multi-scale loss function to optimize decoding efficiency. As the pioneering application of a vision Mamba-based model for HSI classification, our results have significantly outperformed other baselines and state-of-the-art methods with limited and disjoint data samples. However, with augmented training samples, Transformer-based models tend to be better. Moving forward, we plan to extend our idea to include a 3D version of the Mamba model where the scanning consistency on spectral information may work.