### v4

# Mamba-in-Mamba: Centralized Mamba-Cross-Scan in Tokenized Mamba Model for Hyperspectral Image Classification

Abstract
Hyperspectral image (HSI) classification is pivotal in the remote sensing (RS) field, particularly with the advancement of deep learning techniques. Sequential models, such as Recurrent Neural Networks (RNNs) and Transformers, have been tailored to this task, offering unique viewpoints. However, they face several challenges: 1) RNNs struggle with aggregating central features and are sensitive to interfering pixels; 2) Transformers require extensive computational resources and often underperform with limited HSI training samples. To address these issues, recent advances have introduced State Space Models (SSM) and Mamba, known for their lightweight and parallel scanning capabilities within linear sequence processing, providing a balance between RNNs and Transformers. However, deploying Mamba as a backbone for HSI classification has not been fully explored. Although there are improved Mamba models for visual tasks, such as Vision Mamba (ViM) and Visual Mamba (VMamba), directly applying them to HSI classification encounters problems. For example, these models do not effectively handle land-cover semantic tokens with multi-scale perception for feature aggregation when implementing patch-based HSI classifiers for central pixel prediction. Hence, the suitability of these models for this task remains an open question. In response, this study introduces the innovative Mamba-in-Mamba (MiM) architecture for HSI classification, marking the pioneering deployment of Mamba in this field. The MiM model includes: 1) A novel centralized Mamba-Cross-Scan (MCS) mechanism for transforming images into efficient-pair sequence data; 2) A Tokenized Mamba (T-Mamba) encoder that incorporates a Gaussian Decay Mask (GDM), a Semantic Token Learner (STL), and a Semantic Token Fuser (STF) for enhanced feature generation and concentration; and 3) A Weighted MCS Fusion (WMF) module coupled with a Multi-Scale Loss Design to improve model training efficiency. Experimental results from four public HSI datasets with fixed and disjoint training-testing samples demonstrate that our method outperforms existing baselines and is competitive with state-of-the-art approaches, highlighting its efficacy and potential in HSI applications.
keywords:
Hyperspectral image classification, deep learning, state space model, Mamba.
[…]
{graphicalabstract}
{highlights}
We explore the potential design of Mamba architecture, Mamba-in-Mamba model, for HSI classification task thereby achieving the better competitive performances and satisfying efficiency.
To fully explore the ability of Mamba in visual task, we propose a novel Tokenized Mamba encoder with Gaussian Decay Mask, Semantic Token Learner, and Semantic Token Fuser to improve its suitability, aligning with multi-scale features for the HSI classification.
We propose a centralized Mamba-Cross-Scan to comprehensively and continuously scan the image patch into several pair-sequences diversely. It can fully and lightly understand the HSI patch within the concentration on the center pixel for patch-wise HSI classifiers.
This study shows very competitive and even SOTA performance on four datasets, Indian Pines, Pavia University, Houston 2013, and WHU-Hi-HongHu, with fixed and disjoint training-testing samples, demonstrating the feasibility and efficiency.
1 Introduction
Hyperspectral imaging (HSI) is a sophisticated remote sensing (RS) technique that captures images across numerous dense and continuous spectral bands. This method spans the electromagnetic spectrum from visible to infrared wavelengths, often utilizing satellites or unmanned aerial vehicles (UAVs) [1]. Unlike traditional imaging methods, HSI provides detailed spectral information for each pixel, enabling a more precise characterization of materials [2] and objects on the Earth’s surface [3]. The comprehensive data obtained from HSI makes it invaluable in various applications including agriculture [4], environmental monitoring [5], geology [6], urban planning [7], and defense and security [8]. Given the broad utility of this technology, accurate HSI classification, which assigns a land-cover label to each pixel, is critical. This classification task is not only essential for enhancing the effectiveness of these applications but has also become a focus of research within the RS community [9].
In recent years, the deployment of deep learning-based models for HSI classification has seen extensive growth [10]. These models typically employ a ‘patch-wise’ learning framework, where a local patch around each pixel is processed to assign semantic labels [11, 12]. Notably, sequential models like Recurrent Neural Networks (RNNs) [13, 14, 15, 16, 17] and Transformers [18, 19, 20, 21, 22, 23, 24, 25, 26, 27, 28, 29, 30, 31, 32, 33], originally developed for natural language processing (NLP), have been adapted to HSI classification. They have demonstrated impressive results by utilizing many-to-one or many-to-many schemes to capture contextual features within pixels or partitioned windows of one HSI patch, resulting in a comprehensive feature representation. Despite their success, both types of models face shortcomings that warrant further investigation to enhance their practicability and efficiency.
First, RNNs process an HSI patch by flattening it into a sequence and employing a many-to-one learning framework. This approach selects the output at the last step as the representative feature for the entire patch. However, RNNs inherently favor the latter steps of the input sequence, resulting in a feature representation that is biased towards the final processed steps. This bias can adversely affect model training, particularly when using the cross-entropy criterion to compare the predicted label with the true label of the central pixel. Although some strategies use attention-based weighted summations to balance the outputs across all steps, the attention values often remain skewed towards later pixels, diminishing the influence of the central pixel. Furthermore, RNNs also struggle to handle large HSI patches effectively due to their sequential processing nature, which limits parallel computation across pixels. Larger patches convert into longer sequences, increasing the model’s exposure to interfering pixels with labels differing from the central pixel. This discrepancy can significantly degrade training performance as the model becomes more sensitive to peripheral pixel variations rather than focusing on the central context.
Second, to address the inflexibility of RNNs, Transformer-based methods have been introduced, utilizing parallel self-attention to enhance performance [34]. While these methods overcome some of RNNs’ limitations through parallel processing and self-attention mechanisms, they hold their own set of shortcomings: 1) The self-attention mechanism scales quadratically with the size of HSI patches, leading to significant computational overhead. This becomes especially problematic when larger patches are used to capture more spatial or contextual information around the center pixel, making the approach resource-intensive and less suited for hardware-constrained environments; 2) Transformers often have a large number of parameters, increasing the risk of overfitting, especially with the limited training data typically available in HSI classification; 3) Unlike RNNs or convolutional neural networks (CNNs) [35, 36, 37], which inherently capture recurrent features or local spatial features, Transformers do not have this inductive bias. While positional embeddings are used to introduce some spatial context, their effectiveness is limited without substantial training samples, potentially reducing the model’s ability to accurately capture spatial contexts in HSIs. Although some research [38, 39, 40, 41] has explored hybrid approaches that integrate RNNs or CNNs with Transformers to leverage complementary strengths. However, these methods may still increase the computational burden and complexity, potentially leading to suboptimal performance.
Third, the use of sequential models for image interpretation presents challenges, particularly in how images are arranged into sequences due to the non-causal nature of visual data: 1) Many approaches rely on a simple ‘Raster’ scan to flatten the image, which can result in discontinuities, or ‘jump connections,’ when the scan reaches the boundaries of the image; 2) Previous works [42, 43] have demonstrated that a continuous scan approach yields better results in HSI classification. It can maintain the highest spatial correlation by ensuring continuity in scanning across pixels. It also preserves multi-directional information inherent in the image’s direction-sensitive properties. Despite these advantages, existing one [44] often processes scans one by one in a unidirectional manner, then aggregate the results. This method does not necessarily prioritize the central pixel’s features during model training and then increase the feature redundancy and computational burden.
In response to three scenarios discussed above, two critical needs emerge: 1) Alternative Base Model: There is a need for an alternative model that enhances parallel computing or incorporates attention mechanisms, offering greater flexibility than RNNs. This model should also demand fewer resources and exhibit less complexity, proving lighter and more efficient than Transformers, and 2) Enhanced Scanning Method: A refined scanning method is required that supports continuous and multi-directional processing. This method should provide image-aware cognition without being redundant, effectively capturing the spatial correlations and dynamics inherent in image data.
Recently, structured state space sequence (S4) models [45] and the Mamba architecture [46, 47, 48, 49] are brought into the spotlight. These models are adept at handling long-sequence data and can scale linearly with sequence length. The Mamba architecture, in particular, builds on S4 by adding a selective-scan mechanism that adaptively attentions on relevant inputs, showing superior computational efficiency compared to Transformers. Although the Mamba-based methods has shown promising results in some visual tasks [50, 51, 53, 52, 54, 55] (e.g. Vision Mamba (ViM), Visual Mamba (VMamba), etc), its potential for HSI classification has yet to be fully explored. Directly applying these advanced Mamba-based models on HSI classification task may encounter unsuitability and problems. Therefore, there is a critical need to develop a suitable vision-based Mamba model specifically designed for HSI classification.
Figure 1: Conceptual diagram of the Mamba-in-Mamba (MiM) model. Initially, four distinct types of centralized Mamba-Cross-Scan (MCS) are employed to convert a cropped HSI patch into sequential data. It is then fed into four shared Tokenized Mamba (T-Mamba) encoders designed to process sequences effectively. To adjust the impact of each MCS type, a Weighted MCS Fusion (WMF) is conducted, ensuring a balanced integration of the features extracted from each MCS type. Finally, the outputs from cascaded down-sampling-driven MiM layers are fed into a decoder for classification.
To this end, in this study, we introduce a innovative approach—the Mamba-in-Mamba (MiM) architecture—specifically designed for HSI classification. This method marks the pioneering application of the ViM in this task. The MiM model comprises several innovative components: 1) Centralized Mamba-Cross-Scan (MCS): This new scanning mechanism transforms images into diverse half-directional sequences, offering a more efficient, lightweight, and rational approach, 2) Tokenized Mamba (T-Mamba) Encoder: Integrates a Gaussian Decay Mask (GDM), a Semantic Token Learner (STL), and a Semantic Token Fuser (STF) to significantly enhance representative feature generation and concentration, and 3) Weighted MCS Fusion (WMF) module: Works in conjunction with a Multi-Scale Loss Design to boost decoding efficiency and overall model performance. The conceptual framework of this study is illustrated in Fig. 1.
The contributions of this study are summarized as follows:
  • 1)
    We explore the potential applications of the improved ViM architecture, multi-scale MiM model, for HSI classification task thereby achieving the better performances and efficiency.
  • 2)
    To fully explore the ability of ViM, we propose a novel T-Mamba encoder with GDM, STL, and STF to make it more suitable for the HSI classification.
  • 3)
    We propose centralized MCS to comprehensively and continuously scan the HSI patch into several pair-sequences diversely, which can fully and lightly understand the HSI patch within the highlights on the center pixel for patch-wise HSI classifiers.
  • 4)
    The proposal shows very competitive and even SOTA performance on four datasets with fixed and disjoint training-testing samples, demonstrating the feasibility and efficiency.
The rest of the paper is organized as follows: Section 2 presents the preliminary knowledge of Mamba, Section 3 details the methodology, Section 4 discusses the experimental setup and results, and Section 5 concludes the study.
2 Preliminary
This section provides a foundational concepts associated with State Space Models (SSM) and the Mamba model, presenting the knowledge necessary to proceed this study.
2.1 Structured State Space Sequence Model (S4)
The SSM-based models, represent a recent class of sequence models within the field of deep learning. They are broadly related to RNNs, CNNs, and classical state space models.
Inspired by continuous system theory [56], SSMs are commonly considered linear time-invariant systems that map a signal $\mathbf{x}(t)\in\mathbb{R}^{1\times\mathcal{D}}$ to $\mathbf{y}(t)\in\mathbb{R}^{1\times\mathcal{D}}$ through a latent state $\mathbf{h}(t)\in\mathbb{R}^{1\times\mathcal{N}}$. Here, $t$ represents the time step, $\mathcal{D}$ is the initial dimension, and $\mathcal{N}$ is the dimension of the state.
Mathematically, SSMs are typically formulated as linear ordinary differential equations (ODEs) [57], as follows:
$\mathbf{h}^{\prime}(t)$ $=\mathbf{A}\,\mathbf{h}(t)+\mathbf{B}\,\mathbf{x}(t),$ (1a)
$\mathbf{y}(t)$ $=\mathbf{C}\,\mathbf{h}(t)+\mathbf{D}\,\mathbf{x}(t),$ (1b)
where $\mathbf{A}\in\mathbb{R}^{\mathcal{N}\times\mathcal{N}}$ is the evolution parameter, $\mathbf{B}\in\mathbb{R}^{\mathcal{N}\times\mathcal{D}}$ correctly aligns dimensions for input-to-state transformation, and $\mathbf{C}\in\mathbb{R}^{\mathcal{D}\times\mathcal{N}}$ is the projection parameter. $\mathbf{D}\in\mathbb{R}^{1\times\mathcal{D}}$, which typically supports skip connections, ensures that $\mathbf{y}(t)\in\mathbb{R}^{1\times\mathcal{D}}$.
S4 is a discrete version of a continuous system, developed to address the significant challenges posed when integrating continuous systems into deep learning algorithms. A timescale parameter $\mathbf{\Delta}\in\mathbb{R}^{1\times\mathcal{D}}$ is used to transform the continuous parameters $\mathbf{A}$ and $\mathbf{B}$ into discrete parameters $\mathbf{\overline{A}}\in\mathbb{R}^{\mathcal{N}\times\mathcal{N}}$ and $\mathbf{\overline{B}}\in\mathbb{R}^{\mathcal{N}\times\mathcal{D}}$. The transformation commonly employs a zero-order hold (ZOH) method [^1] https://en.wikipedia.org/wiki/Zero-order_hold, defined as:
$\mathbf{\overline{A}}$ $=\exp(\mathbf{\Delta}\mathbf{A}),$ (2a)
$\mathbf{\overline{B}}$ $=(\mathbf{\Delta}\mathbf{A})^{-1}(\exp(\mathbf{\Delta}\mathbf{A})-\mathbf{I})\odot\mathbf{\Delta}\mathbf{B},$ (2b)
Figure 2: The graphic illustration of S4 model. $\mathbf{x}(t)$ is a input feature (i.e. scalar or vector) at time step $t$. $\mathbf{h}(t)$ represents the latent state feature. $\mathbf{\overline{A}}$, $\mathbf{\overline{B}}$, $\mathbf{C}$, and $\mathbf{D}$ represents the transformation matrices.
Therefore, the discretized versions of Eqs. 1a and 1b, can be reformulated not as a function-to-function mapping, but rather as a sequence-to-sequence model, as shown in Fig. 2:
$\mathbf{h}(t)$ $=\mathbf{\overline{A}}\,\mathbf{h}(t-1)+\mathbf{\overline{B}}\,\mathbf{x}(t),$ (3a)
$\mathbf{y}(t)$ $=\mathbf{C}\,\mathbf{h}(t)+\mathbf{D}\,\mathbf{x}(t).$ (3b)
Finally, the models compute the output through a global convolution operation as follows:
$\mathbf{\overline{K}}$ $=\Big{(}\mathbf{C}\,\mathbf{\overline{B}},\mathbf{C}\,\mathbf{\overline{A}}\,\mathbf{\overline{B}},\,\ldots,\,\mathbf{C}\,\mathbf{\overline{A}}^{L-1}\,\mathbf{\overline{B}}\Big{)},$ (4a)
$\mathbf{y}$ $=\mathbf{x}\star\mathbf{\overline{K}},$ (4b)
where $L$ is the length of the input sequence $\mathbf{x}$, and $\mathbf{\overline{K}}\in\mathbb{R}^{\mathcal{L}}$ is a structured SSM kernel. $\star$ denotes the convolutional operation.
The notable work [58] states that the matrices $\mathbf{A}$, $\mathbf{B}$, and $\mathbf{C}$ in SSM share a crucial attribute known as Linear Time Invariance (LTI). It means that they are constant and remain unchanged for each token generated by the SSM. In essence, regardless of the input sequence presented to the SSM, the values of $\mathbf{A}$, $\mathbf{B}$, and $\mathbf{C}$ remain fixed, resulting in a static representation that is not content-aware. Additionally, the matrices $\mathbf{A}$, $\mathbf{B}$, and $\mathbf{C}$ are independent of the input, as their dimensions $\mathcal{N}$ and $\mathcal{D}$ do not vary, further emphasizing their time-invariant nature.
2.2 Selective Scan-Based S4 Model (S6)
To address the limitations identified in the S4 model, Mamba enhances the performance of the S4 model by introducing a selective scan mechanism as the core S4 operator, termed as S6 model. It allows the continuous parameters to vary with the input, thereby enhancing selective information processing across sequences.
It modifies the S4 model by introducing adaptive parameterization mechanisms for the matrices $\mathbf{B}$, $\mathbf{C}$, and the timescale parameter $\mathbf{\Delta}$ as follows:
$\mathbf{B}$ $={Projection}_{B}(\mathbf{x}),$ (5a)
$\mathbf{C}$ $={Projection}_{C}(\mathbf{x}),$ (5b)
$\mathbf{\Delta}$ $=\tau_{\Delta}({Parameter}+{Projection}_{\Delta}(\mathbf{x})),$ (5c)
where $\mathbf{x}\in\mathbb{R}^{\mathcal{B}\times\mathcal{L}\times\mathcal{D}}$. Here, $\mathcal{B}$ represents the batch size, $\mathcal{L}$ denotes the sequence length, and $\mathcal{D}$ is the sequence dimension. The functions ${Projection}_{B}()$ and ${Projection}_{C}()$ perform linear projections from the $\mathcal{D}$-dimensional space to the $\mathcal{N}$-dimensional space, resulting in $\mathbf{B}\in\mathbb{R}^{\mathcal{B}\times\mathcal{L}\times\mathcal{N}}$ and $\mathbf{C}\in\mathbb{R}^{\mathcal{B}\times\mathcal{L}\times\mathcal{N}}$. Additionally, ${Projection}_{\Delta}()$ transforms $\mathbf{x}$ to a size of ××$\mathbb{R}^{\mathcal{B}\times\mathcal{L}\times 1}$, which is then broadcast to the $\mathcal{D}$-dimension. $\tau_{\Delta}$, typically a softplus activation [59], adjusts the scale, resulting in a new $\mathbf{\Delta}$ of size ××$\mathbb{R}^{\mathcal{B}\times\mathcal{L}\times\mathcal{D}}$.
In a summary, Mamba enhances the S4 model by making the matrices $\mathbf{B}$, $\mathbf{C}$, and the step size $\mathbf{\Delta}$ input-dependent, incorporating the sequence length and batch size to ensure that each input token is processed with distinct matrices. Moreover, these dynamic matrices selectively filter which components to maintain or omit in the hidden state, drawing an analogy to the ‘recurrence’ mechanism in RNNs. This selective processing is implemented through a ‘parallel scan’ operation [^2] https://developer.nvidia.com/gpugems/gpugems3/part-vi-gpu-computing/chapter-39-parallel-prefix-sum-scan-cuda in CUDA, which segments the sequences, processes them iteratively, and combines them efficiently. This integration of dynamic matrices and parallel processing exemplifies how Mamba captures the fast and adaptable nature of sequence handling, thereby enhancing the model’s adaptability.
2.3 Mamba-Based Works
As a cutting-edge foundational model, Mamba has recently gained popularity. There have been various implementations exploring its applicability to diverse tasks.
Zhu et al. [51] pioneered the ViM architecture for general vision tasks such as object detection and semantic segmentation, employing bidirectional scans for precise image patch positioning prior to processing with the Mamba encoder. Building on this framework, Chen et al. [52] enhanced the Mamba with a dynamic multi-path activation mechanism, significantly improving performance across various RS image classification datasets. Concurrently, Liu et al. [53] developed VMamba, which utilizes 2D-Selective-Scan (SS2D) to transform non-causal visual images into ordered patch sequences for applications in image classification and object detection. Further extending this methodology, Chen et al. [54] applied VMamba to RS change detection tasks, encompassing binary change detection, semantic change detection, and building damage assessment, with VMamba serving as the sequential encoder to model spatio-temporal relationships. Additionally, Zhao et al. [60] introduced RS-Mamba for dense prediction in extensive RS images, incorporating the Omni-directional Selective Scan Module (OSSM) for feature aggregation. In the HSI classification domain, Jia et al. [61] adapted the standard SSM-based model, SpectralMamba, for spectral domain processing, while Yang et al. [62] combined Mamba’s spectral processing with CNN-based spatial feature extraction for HSI classification. Moreover, Sheng et al. [63] proposed a lightweight spectral-spatial Mamba-convolution model for HSI classification, utilizing spatial unidirectional and spectral bidirectional scanning for enhanced feature alignment.
Despite these advancements, to our knowledge, the application of ViM with multi-directional scanning for HSI classification has not yet been fully developed or explored. Meanwhile, directly using ViM can not well handle the refined semantic token features and miss the multi-scale features existed in the HSI patch. Hence, developing an optimized ViM architecture for HSI remains an open question. Thus, we aim to explore this gap and propose a tailored HSI-driven ViM architecture.
Figure 3: The overall structure. It begins with a cropped HSI patch after PCA and then goes through a depth-wise convolution and a point-wise convolution for local feature generation and dynamic positional embedding. It is the reshaped by snake flattening before MiM model. In the MiM model, flattened image is first proposed by MCS, as shown in Fig. 5, to obtain four types of pair-sequences. Each pair of sequences processed by a specially designed T-Mamba encoder with sharing parameters, as detailed in Fig. 6, to extract contextual features. The outputs from the four T-Mamba encoders are combined through a WMF module with dynamic weighted scan fusion. Concurrently, fused feature at current scale is recorded for final multi-scale decoding. This process is repeated, with the feature sizes being progressively down-sampled until they reach a spatial size of $1$. Finally, the accumulated multi-scale features from each MiM iteration are compiled to feed into a decoder for model training and decoding.
3 Methodology
This section firstly presents (A) methodological overview briefly. Then, it presents the detailed contents: (B) Centralized MCS, (C) T-Mamba Encoder, (D) WMF Module, (E) Multi-Scale Loss Design. The overall structure and the pseudo algorithm is shown in Fig. 3 and Algorithm 1, respectively. Due to the paper formatting, the summary of algorithm is placed at end of the article.
3.1 Methodological Overview
Our method is composed of several following steps:
  1. 1)
    We introduce a centralized MCS mechanism that transforms an HSI patch into four types of sequences, each featuring two directional scans.
  2. 2)
    The T-Mamba encoder is engineered to efficiently manage two directional scans. It utilizes the GDM, a downsample-driven STL, and an attention-driven STF to facilitate the effective learning of representative tokens.
  3. 3)
    The MiM model processes the four types of sequences generated by the MCS across four T-Mamba encoders, setting the foundation for subsequent feature fusion.
  4. 4)
    A WMF is developed to dynamically assign different weights to the outputs from the MiM model. Meanwhile, a scale feature, derived from the fused outputs, is used to represent this phase for later decoding.
  5. 5)
    Due to the down-sampling-like operation by the MiM model, we iterate the aforementioned procedures until the spatial size of the feature reduces to 1.
  6. 6)
    The accumulated multi-scale features, are prepared for the decoder. Each feature is linked to a specific loss function to optimize performance during training.
Detailed explanations for each component are provided in the subsequent subsections later.
3.2 Centralized Mamba-Cross-Scan (MCS)
The original HSI symbolized as $\mathbf{\mathbf{X}}\in\mathbb{R}^{\mathcal{H}\times\mathcal{W}\times\mathcal{C}_{0}}$ can be defined as follows:
$\centering\mathbf{\mathbf{X}}=\Big{\{}\mathbf{\mathbf{x}}^{(i,j)}\,\Big{|}\,i=0,1,...,\mathcal{H}-1,\,j=0,1,...,\mathcal{W}-1\Big{\}},\@add@centering$ (6)
where $\mathbf{\mathbf{x}}^{(i,j)}\in\mathbb{R}^{\mathcal{C}_{0}}$ signifies the spectral signature at the spatial position $(i,j)$. Consequently,
$\centering\mathbf{\mathbf{X}}=\left(\begin{matrix}\mathbf{\mathbf{x}}^{(0,0)}&\mathbf{\mathbf{x}}^{(0,1)}&\dots&\mathbf{\mathbf{x}}^{(0,\mathcal{W}-1)}\\ \mathbf{\mathbf{x}}^{(1,0)}&\mathbf{\mathbf{x}}^{(1,1)}&\dots&\mathbf{\mathbf{x}}^{(1,\mathcal{W}-1)}\\ \vdots&\vdots&\ddots&\vdots\\ \mathbf{\mathbf{x}}^{(\mathcal{H}-1,0)}&\mathbf{\mathbf{x}}^{(\mathcal{H}-1,1)}&\dots&\mathbf{\mathbf{x}}^{(\mathcal{H}-1,\mathcal{W}-1)}.\end{matrix}\right).\@add@centering$ (7)
Designed as a patch-wise framework, the input for the neural model is a cropped patch from the original HSI. For one cropped HSI patch $\mathbf{X}^{(i,\,j)}$ centered at $\mathbf{x}^{(i,\,j)}$ with patch size $p$, an odd number, the patch can be represented as:
$\mathbf{X}^{(i,\,j)}=\Big{\{}\,\mathbf{x}^{(i+\alpha,\,j+\beta)}\,\Big{|}\,\alpha,\beta=-\frac{p-1}{2},\ldots,\frac{p-1}{2}\Big{\}},$ (8)
where $\alpha,\beta\in\mathbb{Z}$ and $i+\alpha$ and $j+\beta$ record the position.
Taking $p=5$ as an example, the cropped HSI patch $\mathbf{X}^{(i,\,j)}$ can be represented as:
$\centering\mathbf{X}^{(i,\,j)}=\left(\begin{matrix}\mathbf{x}^{(i-2,\,j-2)}&\mathbf{x}^{(i-2,\,j-1)}&\dots&\mathbf{x}^{(i-2,\,j+2)}\\ \mathbf{x}^{(i-1,\,j-2)}&\mathbf{x}^{(i-1,\,j-1)}&\dots&\mathbf{x}^{(i-1,\,j+2)}\\ \vdots&\vdots&\ddots&\vdots\\ \mathbf{x}^{(i+2,\,j-2)}&\mathbf{x}^{(i+2,\,j-1)}&\dots&\mathbf{x}^{(i+2,\,j+2)}\end{matrix}\right),\@add@centering$ (9)
where $32\leq i\leq\mathcal{H}-$ 3, $32\leq j\leq\mathcal{W}-$ 3, and $\mathbf{X}^{(i,\,j)}\in\mathbb{R}^{p\times p\times\mathcal{C}_{0}}$. The center pixel $\mathbf{X}^{(i,\,j)}[2,2]$ is $\mathbf{x}^{(i,\,j)}$.
The existing centralized scanning method, SS2D, as utilized in [53], employs traditional Raster flattening across the spatial domain of images. This approach can lead to misalignment of image contents, particularly through jump connections that occur when the scan meets image boundaries. Furthermore, SS2D does not fully embrace the two-dimensional structure of the image, limiting its ability to analyze from multiple perspectives.
Consequently, we have opted for the Snake scan method (e.g., U-Turn scan), which has been endorsed in various studies [42, 43] and serves as the foundation for our proposed MCS methodology. The Snake scan method preserves spatial continuity between pixels or partitioned windows when transforming them into sequences, reducing the incidence of jump connections at image boundaries. This method’s effectiveness is further supported by the multi-scanning strategy concept [44], which demonstrates that using sequences scanned from multiple directions can enhance sequential model training.
Leveraging this background, we have developed four types of centralized MCS that optimize the transformation of images into sequences from various starting and ending points, ensuring both potential maximization and system efficiency. The fundamental distinction between the Transformer’s self-attention and our MCS mechanism is illustrated in Fig. 4.
As depicted in Fig. 5, each MCS type initiates from two opposing positions and converges at the center pixel. This method is reminiscent of the bi-directional processing typical of RNNs but incorporates a centralized design, providing a novel approach to image sequence processing.
Figure 4: The conceptual difference between (a) self-attention mechanism in Transformers, and (b) proposed MCS mechanism with T-Mamba in overall structure for recognizing global contextual information in the patch.
Before feeding the $\mathbf{X}^{(i,\,j)}$ into MCS, we first deploy the principal component analysis (PCA), depth-wise convolution (DWconv) [^3] https://github.com/rosinality/depthwise-conv-pytorch and point-wise convolution (PWconv) [^4] https://github.com/anshulpaigwar/Pointwise-Convolutional-Neural-Network on $\mathbf{X}^{(i,\,j)}$ for dimension reduction, local feature aggregation and dynamic positional embedding, as:
$\mathbf{X}^{(i,\,j)}\leftarrow PWconv\Big{(}DWconv\Big{(}PCA\Big{(}\mathbf{X}^{(i,\,j)}\Big{)}\Big{)}\Big{)}$ (10)
where the PCA preserves the number of $\mathcal{C}_{1}$ components. The PWconv and DWconv outputs number of $\mathcal{C}_{2}$ feature dimensions, where $\mathcal{C}_{2}>\mathcal{C}_{1}$. Hence, the resultant $\mathbf{X}^{(i,\,j)}\in\mathbb{R}^{p\times p\times\mathcal{C}_{2}}$.
3.2.1 MCS Type-1
In details, as shown in Fig. 5(a), the type-1 MCS starts from the two positions, $\mathbf{x}^{(i-2,\,j-2)}$ and $\mathbf{x}^{(i+2,\,j+2)}$. Instead of directly preparing two directional sub-sequences, we firstly flatten the whole patch into a complete sequence by snake scanning, starting from $\mathbf{x}^{(i-2,\,j-2)}$, as:
$\begin{array}[]{ccccccccc}&\overbrace{{\color[rgb]{0,1,0}\mathbf{\mathbf{x}}^{(i-2,j-2)}}}^{Start}&\rightarrow&\mathbf{\mathbf{x}}^{(i-2,j-1)}&\rightarrow&\cdots&\rightarrow\mathbf{\mathbf{x}}^{(i-2,j+2)}\\ &&&&&&\downarrow\\ &\mathbf{\mathbf{x}}^{(i-1,j-2)}&\leftarrow&\mathbf{\mathbf{x}}^{(i-1,j-1)}&\leftarrow&\cdots&\leftarrow\mathbf{\mathbf{x}}^{(i-1,j+2)}\\ &\downarrow&&&&&\\ &\mathbf{\mathbf{x}}^{(i,j-2)}&\rightarrow&\mathbf{\mathbf{x}}^{(i,j-1)}&\rightarrow&\cdots&\rightarrow\mathbf{\mathbf{x}}^{(i,j+2)}\\ &&&&&&\downarrow\\ &\mathbf{\mathbf{x}}^{(i+1,j-2)}&\leftarrow&\mathbf{\mathbf{x}}^{(i+1,j-1)}&\leftarrow&\cdots&\leftarrow\mathbf{\mathbf{x}}^{(i+1,j+2)}\\ &\downarrow&&&&&\\ &\mathbf{\mathbf{x}}^{(i+2,j-2)}&\rightarrow&\mathbf{\mathbf{x}}^{(i+2,j-1)}&\rightarrow&\cdots&\rightarrow\underbrace{{\color[rgb]{0,0,1}\mathbf{\mathbf{x}}^{(i+2,j+2)}}}_{End}\\ \end{array}$ (11)
Therefore, the generated complete sequence $\mathbf{S}_{1}^{(i,\,j)}\in\mathbb{R}^{p^{2}\times\mathcal{C}_{2}}$ can be represented as:
$\begin{split}\mathbf{S}^{(i,j)}_{1}=&[{\color[rgb]{0,1,0}\mathbf{x}^{(i-2,j-2)}},\mathbf{x}^{(i-2,j-1)},\cdots,\mathbf{x}^{(i-2,j+2)},\cdots\\ &\cdots,\mathbf{x}^{(i+2,j-2)},\mathbf{x}^{(i+2,j-1)},\cdots,{\color[rgb]{0,0,1}\mathbf{x}^{(i+2,j+2)}}]^{\top},\end{split}$ (12)
where each element of the $\mathbf{S}_{1}^{(i,\,j)}$ is a spectral vector, in which the $\mathbf{S}_{1}^{(i,\,j)}[\frac{p^{2}-1}{2}]$ is the center step $\mathbf{x}^{(i,j)}$ of the sequence.
Figure 5: Four types of MCS with a $5\times 5$ patch. (a) Type-1, (b) Type-2, (c) Type-3, and (d) Type-4. They cover the all the possible continuous snake scanning for a patch.
Subsequently, we split the $\mathbf{S}_{1}^{(i,\,j)}$ into two sub-sequences, $\mathbf{S}_{1,\,f}^{(i,\,j)}$ and $\mathbf{S}_{1,\,b}^{(i,\,j)}$. In this case, the $\mathbf{S}_{1,\,f}^{(i,\,j)}$ starts from $\mathbf{x}^{(i-2,\,j-2)}$, and ends at $\mathbf{x}^{(i,\,j)}$; meanwhile, the $\mathbf{S}_{1,\,b}^{(i,\,j)}$ starts from $\mathbf{x}^{(i+2,\,j+2)}$, and ends at $\mathbf{x}^{(i,\,j)}$, as follows:
$\mathbf{S}^{(i,j)}_{1,\,f},\,\mathbf{S}^{(i,j)}_{1,\,b}\leftarrow{Split}\Big{(}\mathbf{S}_{1}^{(i,\,j)}\Big{)},$ (13)
where $\mathbf{S}_{1,\,f}^{(i,\,j)}$ and $\mathbf{S}_{1,\,b}^{(i,\,j)}$ are arranged as:
$\mathbf{S}^{(i,j)}_{1,\,f}$ $=\Big{[}\underbrace{{\color[rgb]{0,1,0}\mathbf{x}^{(i-2,j-2)}},\mathbf{x}^{(i-1,j-2)},\mathbf{x}^{(i,j-2)},\cdots,{\color[rgb]{1,.5,0}\mathbf{x}^{(i,j)}}}_{{length}=(p^{2}+1)/2}\Big{]}^{\top},$ (14a)
$\mathbf{S}^{(i,j)}_{1,\,b}$ $=\Big{[}\underbrace{{\color[rgb]{0,0,1}\mathbf{x}^{(i+2,j+2)}},\mathbf{x}^{(i+1,j+2)},\mathbf{x}^{(i,j+2)},\cdots,{\color[rgb]{1,.5,0}\mathbf{x}^{(i,j)}}}_{{length}=(p^{2}+1)/2}\Big{]}^{\top},$ (14b)
where two sub-sequences termed as forward-direction and backward-direction, both include the center step $\mathbf{x}^{(i,j)}$ of the complete sequence and assign it at the last step in the sub-sequence. In this case, when patch size $p$ equals 5, the length of sub-sequence is 13, as shown in Fig. 5(a).
Consequently, take $\mathbf{S}^{(i,j)}_{1,\,f}$ as an example, we can feed it into Mamba model respectively, referring Eqs. 3a and 3b as:
$\mathbf{h}^{(i,j)}_{1,\,f}[t]$ $=\mathbf{\overline{A}}\,\mathbf{h}^{(i,j)}_{1,\,f}[t-1]+\mathbf{\overline{B}}\,\mathbf{S}^{(i,j)}_{1,\,f}[t],$ (15a)
$\mathbf{y}^{(i,j)}_{1,\,f}[t]$ $=\mathbf{C}\,\mathbf{h}^{(i,j)}_{1,\,f}[t]+\mathbf{D}\,\mathbf{S}^{(i,j)}_{1,\,f}[t],$ (15b)
where $t\in\mathbb{Z}$, and $10\leq t\leq\frac{p^{2}+1}{2}-$ 1, denoting the element index (step) in the sub-sequence. Therefore, the last step, $\mathbf{x}^{(i,j)}$, center of the patch, can iterate all previous steps’ features. By doing the same operation for the another sub-sequence, $\mathbf{S}^{(i,j)}_{1,\,b}$, the whole image patch’s feature can be merged by two of them.
$\mathbf{y}^{(i,j)}_{1}={Merge}\Big{(}\mathbf{y}^{(i,j)}_{1,\,f},\,\mathbf{y}^{(i,j)}_{1,\,b}\Big{)},$ (16)
where $\mathbf{y}^{(i,j)}_{1,\,f}$ and $\mathbf{y}^{(i,j)}_{1,\,b}$ are the output sequence for two sub-sequences. And $\mathbf{y}^{(i,j)}_{1}$ is the merged feature from type-1 MCS for representing whole image feature. The ${Merge}()$ operation will be detailed in the next section.
3.2.2 MCS Type-2
Hence, following the same description of type-1 MCS, the type-2 MCS with vertical manner, referring Fig. 5(b), can be formulated as:
$\mathbf{S}^{(i,j)}_{2,\,f}$ $=\Big{[}{\color[rgb]{0,1,0}\mathbf{x}^{(i-2,j-2)}},\mathbf{x}^{(i-2,j-1)},\mathbf{x}^{(i-2,j)},\cdots,{\color[rgb]{1,.5,0}\mathbf{x}^{(i,j)}}\Big{]}^{\top},$ (17a)
$\mathbf{S}^{(i,j)}_{2,\,b}$ $=\Big{[}{\color[rgb]{0,0,1}\mathbf{x}^{(i+2,j+2)}},\mathbf{x}^{(i+2,j+1)},\mathbf{x}^{(i+2,j)},\cdots,{\color[rgb]{1,.5,0}\mathbf{x}^{(i,j)}}\Big{]}^{\top}.$ (17b)
In this case, the two sub-sequences $\mathbf{S}^{(i,j)}_{2,\,f}$ and $\mathbf{S}^{(i,j)}_{2,\,b}$ starts at $\mathbf{x}^{(i-2,\,j-2)}$ and $\mathbf{x}^{(i+2,\,j+2)}$, respectively and ends at $\mathbf{x}^{(i,j)}$ both with a vertical scanning manner. The length of them is same with the previous description.
3.2.3 MCS Type-3
Different from the type-1 and type-2 which starts from two positions of $\mathbf{x}^{(i-2,\,j-2)}$ and $\mathbf{x}^{(i+2,\,j+2)}$, the type-3 cross Mamba scan is starting from $\mathbf{x}^{(i-2,\,j+2)}$ and $\mathbf{x}^{(i-2,\,j+2)}$ with a horizontal manner, shown in Fig. 5(c). The formulated two sub-sequences are listed below:
$\mathbf{S}^{(i,j)}_{3,\,f}$ $=\Big{[}{\color[rgb]{0,1,0}\mathbf{x}^{(i-2,j+2)}},\mathbf{x}^{(i-1,j+2)},\mathbf{x}^{(i,j+2)},\cdots,{\color[rgb]{1,.5,0}\mathbf{x}^{(i,j)}}\Big{]}^{\top},$ (18a)
$\mathbf{S}^{(i,j)}_{3,\,b}$ $=\Big{[}{\color[rgb]{0,0,1}\mathbf{x}^{(i+2,j-2)}},\mathbf{x}^{(i+1,j-2)},\mathbf{x}^{(i,j-2)},\cdots,{\color[rgb]{1,.5,0}\mathbf{x}^{(i,j)}}\Big{]}^{\top}.$ (18b)
3.2.4 MCS Type-4
Similar with the type-3 MCS, the type-4 MCS starts from $\mathbf{x}^{(i-2,\,j-2)}$ and $\mathbf{x}^{(i+2,\,j+2)}$ positions with a vertical manner, shown in Fig. 5(d). Two sub-sequences are formulated below:
$\mathbf{S}^{(i,j)}_{4,\,f}$ $=\Big{[}{\color[rgb]{0,1,0}\mathbf{x}^{(i-2,j+2)}},\mathbf{x}^{(i-2,j+1)},\mathbf{x}^{(i-2,j)},\cdots,{\color[rgb]{1,.5,0}\mathbf{x}^{(i,j)}}\Big{]}^{\top},$ (19a)
$\mathbf{S}^{(i,j)}_{4,\,b}$ $=\Big{[}{\color[rgb]{0,0,1}\mathbf{x}^{(i+2,j+2)}},\mathbf{x}^{(i+2,j+1)},\mathbf{x}^{(i+2,j)},\cdots,{\color[rgb]{1,.5,0}\mathbf{x}^{(i,j)}}\Big{]}^{\top},$ (19b)
Consequently, we have generate four types of MCS with totally eight sub-sequences. Each type of them will be input for the following proposed T-Mamba encoder with bi-directional design for feature generation.
Different from the other works, our proposed MCS is designed as a centralized manner with the snake scanning method, which is more suitable for patch-wise HSI classifier where the center pixel’s label is assigned for the patch for training. Moreover, the lightweight design of using two sub-sequences for representing whole image patch can largely save computational burden and complexity than other works which use two complete sequences for whole image interpretation.
Figure 6: The T-Mamba encoder. Take inputs of size $\mathbb{R}^{(5^{2})\times C_{2}}$ as an example. Initially, the input is split into forward and backward sequences $\mathbf{s}_{f}$ and $\mathbf{s}_{b}$ by MCS, each processed independently through a Conv1d and an S6 model. A Gaussian decay mask then statically adjusts the features of each sequence. Two processed feature sequences are then merged for feature aggregation. A downsample-driven Semantic Token Learner (STL) subsequently generates representative semantic tokens, which are correlated with $\mathbf{z}$ by an attention-driven Semantic Token Fuser (STF) to implement feature enhancement. Finally, a residual connection from the original input is used to enhance feature integration for output to the next stage.
3.3 Tokenized Mamba (T-Mamba)
After obtaining two sub-sequences from each type of MCS, we introduce the novel T-Mamba encoder, which processes them in a bi-directional manner. An overview of the T-Mamba is shown in Fig. 6.
3.3.1 Whole Flowchart of T-Mamba
Taking Eq. 12 as an example, for a type-1 MCS, we start with an embedded sequence $\mathbf{S}_{1}^{(i,\,j)}\in\mathbb{R}^{5^{2}\times\mathcal{C}_{2}}$. This sequence is first normalized and then projected into a sequence as follows:
$\mathbf{z}={Linear}\Big{(}{Norm}\Big{(}\mathbf{S}_{1}^{(i,\,j)}\Big{)}\Big{)},$ (20)
where $\mathbf{z}$ is of size $\mathbb{R}^{5^{2}\times\mathcal{C}_{3}}$.
Next, $\mathbf{S}_{1}^{(i,\,j)}$ is split into two directional sub-sequences, $\mathbf{s}_{f}$ and $\mathbf{s}_{b}$ by MCS, where $f$ denotes the forward sub-sequence, and $b$ denotes the backward sub-sequence with each sub-sequence having dimensions $\mathbb{R}^{13\times\mathcal{C}_{3}}$, referring Eq. 13:
$\mathbf{s}_{f},\,\mathbf{s}_{b}\leftarrow{MCS_{1}}\Big{(}\mathbf{S}_{1}^{(i,\,j)}\Big{)},$ (21)
where $MCS_{1}()$ denotes the split operation under type-1 MCS.
Both $\mathbf{s}_{f}$ and $\mathbf{s}_{b}$ are then processed through a $SiLU()$ activation function [^5] https://pytorch.org/docs/stable/generated/torch.nn.SiLU.html and a 1D convolution to prepare the features:
$\overline{\mathbf{s}_{f}}$ $={SiLU}\Big{(}{Conv1d}_{f}\Big{(}\mathbf{s}_{f}\Big{)}\Big{)},$ (22a)
$\overline{\mathbf{s}_{b}}$ $={SiLU}\Big{(}{Conv1d}_{b}\Big{(}\mathbf{s}_{b}\Big{)}\Big{)}.$ (22b)
Subsequently, these processed sub-sequences are input to the Mamba (S6) model as described in Section. 2:
$\widetilde{\mathbf{s}_{f}}$ $={Mamba}_{f}\Big{(}\overline{\mathbf{s}_{f}}\Big{)},$ (23a)
$\widetilde{\mathbf{s}_{b}}$ $={Mamba}_{b}\Big{(}\overline{\mathbf{s}_{b}}\Big{)}.$ (23b)
A Gaussian decay mask (GDM) is then applied to softly select features from the outputs of the Mamba models:
$\widehat{\mathbf{s}_{f}}$ $=\widetilde{\mathbf{s}_{f}}\otimes{GDM}_{f}\Big{(}\widetilde{\mathbf{s}_{f}}\Big{)},$ (24a)
$\widehat{\mathbf{s}_{b}}$ $=\widetilde{\mathbf{s}_{b}}\otimes{GDM}_{b}\Big{(}\widetilde{\mathbf{s}_{b}}\Big{)},$ (24b)
where ${GDM}()$ applies the Gaussian decay mask and will be detailed later, and $\otimes$ denotes the Hadamard multiplication.
The resultant features, $\widehat{\mathbf{s}_{f}}$ and $\widehat{\mathbf{s}_{b}}$, are merged to form a complete feature sequence:
$\widehat{\mathbf{s}}={Merge}\Big{(}\widehat{\mathbf{s}_{f}},\,\widehat{\mathbf{s}_{b}}\Big{)},$ (25)
where $\widehat{\mathbf{s}}\in\mathbb{R}^{5^{2}\times C_{3}}$ and the ${Merge}()$ operation details will be provided later.
The merged sequence $\widehat{\mathbf{s}}$ is then processed by a Semantic Token Learner (STL) module to automatically learn representative semantic tokens and perform down-sampling:
$\mathbf{u}={STL}\left(\widehat{\mathbf{s}}\right),$ (26)
where $\mathbf{u}\in\mathbb{R}^{3^{2}\times C_{3}}$. It is crucial to maintain an odd spatial patch size to keep a center pixel.
These down-sampled semantic tokens $\mathbf{u}$ are then multiplied with the previously prepared $\mathbf{z}$ for analogous attention application:
$\widehat{\mathbf{u}}=STF\Big{(}\mathbf{u},\,{AdaptivePool}\Big{(}{SiLU}\Big{(}\mathbf{z}\Big{)}\Big{)}\Big{)},$ (27)
where $\widehat{\mathbf{u}}\in\mathbb{R}^{3^{2}\times C_{3}}$. ${AdaptivePool}()$ is an adaptive pooling operation that reduces an odd patch size to a smaller odd patch size. $STF()$ represents the operation in Semantic Token Fuser (STF), detailed later.
Finally, the residual connection from the original input is added as follows:
$\mathbf{o}^{(i,\,j)}_{1}={Linear}\Big{(}\widehat{\mathbf{u}}\Big{)}+{Linear}\Big{(}{AdaptivePool}\Big{(}\mathbf{S}_{1}^{(i,\,j)}\Big{)}\Big{)}.$ (28)
These operations at type-1 MCS results in semantic tokens representing the entire image patch, as illustrated in Fig. 6.
3.3.2 Gaussian Decay Mask (GDM)
To well distribute the influence of features from each step in the sequence, it is essential to apply nuanced weighting to each step’s features. Rather than relying on binary values of 0 and 1 to decide which features to retain or eliminate, we introduce the GDM as a soft mask. This GDM assigns weights based on spectral and index distances within the sequence, focusing around the central pixel.
For instance, consider a forward feature sequence $\widetilde{\mathbf{s}_{f}}$ outputted by the Mamba model. It can be represented by:
$\widetilde{\mathbf{s}_{f}}=\Big{[}\underbrace{\widetilde{\mathbf{s}_{f}}[0],\,\widetilde{\mathbf{s}_{f}}[1],\,\widetilde{\mathbf{s}_{f}}[2],\,\cdots,\,\widetilde{\mathbf{s}_{f}}[T]}_{length=(p^{2}+1)/2}\Big{]}^{\top},$ (29)
where $t\in\mathbb{Z}$, $0\leq t\leq T$, and $T=\frac{p^{2}+1}{2}-1$, indexing the elements (steps) in the sub-sequence. The last step of the sub-sequence represents the center pixel in the patch. Therefore, the spatial GDM is calculated by evaluating the index difference between each step and the last step.
Specifically, the $\sigma_{idx}$, representing the average index distance to the last index, is computed as:
$\sigma_{idx}=\frac{1}{T}\sum_{t=0}^{T}\Big{|}t-T\Big{|}$ (30)
The weights for each index $t$ in the sub-sequence are then calculated using a Gaussian distribution centered at $T$:
$w_{t}=\exp\left(-\frac{1}{2}\left(\frac{t-T}{\sigma_{idx}}\right)^{2}\right)$ (31)
These weights are normalized to ensure they sum to one, forming the final spatial GDM for the sub-sequence:
$\mathbf{W}_{idx}[t]=\frac{w_{t}}{\sum_{t=0}^{T}w_{t}},$ (32)
where index-based weights $\mathbf{W}_{idx}$ are applied to each element in the sub-sequence $\widetilde{\mathbf{s}_{f}}$.
Additionally, the feature-based GDM is calculated based on the Euclidean distance between each step (feature vector) and the last step, where the $\sigma_{spe}$, representing the average feature distance to the last step, is computed as:
$\sigma_{fea}=\frac{1}{T}\sum_{t=0}^{T}\Big{|}\Big{|}\widetilde{\mathbf{s}_{f}}[t]-\widetilde{\mathbf{s}_{f}}[T]\Big{|}\Big{|}^{2}$ (33)
The weights are then calculated by:
$v_{t}=\exp\left(-\frac{1}{2}\left(\frac{\Big{|}\Big{|}\widetilde{\mathbf{s}_{f}}[t]-\widetilde{\mathbf{s}_{f}}[T]\Big{|}\Big{|}^{2}}{\sigma_{fea}}\right)^{2}\right),$ (34)
where ||.$||.||^{2}$ represents the L2-distance (Euclidean distance).
Finally, spectral-based weights $\mathbf{W}_{fea}$ can be obtained by:
$\mathbf{W}_{fea}[t]=\frac{v_{t}}{\sum_{t=0}^{T}v_{t}}$ (35)
Consequently, two kinds of GDMs are designed to softly select features from the sequence as:
$\mathbf{W}_{idx},\mathbf{W}_{fea}\leftarrow{GDM}(\widetilde{\mathbf{s}_{f}})$ (36)
Therefore, the Eq. 24a can be rewritten as:
$\widehat{\mathbf{s}_{f}}=\widetilde{\mathbf{s}_{f}}\otimes{Norm}\Big{(}\mathbf{W}_{idx}\otimes\mathbf{W}_{fea}\Big{)}$ (37)
The same operation is applied on the $\widetilde{\mathbf{s}_{b}}$ as well.
This approach may allow for a more refined control over feature selection, enhancing the model’s ability to capture and emphasize relevant spatial and spectral information.
3.3.3 Cross Scan Merge
To effectively merge features from two half-directional sequences and restore them to the original sequence length, we have designed the ${Merge}()$ function. This function is crafted to seamlessly integrate two sub-sequences, honoring their temporal dynamics and optimally preparing them for subsequent processing layers. The computational steps involved in this merging process are detailed below, as referenced in Eq. 25.
Given two sub-sequences, $\widehat{\mathbf{s}_{f}}$ and $\widehat{\mathbf{s}_{b}}$, each comprising $T$ vectors plus a final vector in $\mathbb{R}^{\mathcal{C}_{3}}$, we first partition these sequences as follows:
$\widehat{\mathbf{q}_{f}}$ $=\Big{[}\widetilde{\mathbf{s}_{f}}[0],\,\widetilde{\mathbf{s}_{f}}[1],\,\widetilde{\mathbf{s}_{f}}[2],\,\cdots,\,\widetilde{\mathbf{s}_{f}}[T-1]\Big{]}^{\top},$ (38a)
$\widehat{\mathbf{q}_{b}}$ $=\Big{[}\widetilde{\mathbf{s}_{b}}[T-1],\,\widetilde{\mathbf{s}_{b}}[T-2],\,\widetilde{\mathbf{s}_{b}}[T-3],\,\cdots,\,\widetilde{\mathbf{s}_{b}}[0]\Big{]}^{\top},$ (38b)
$\mathbf{c}$ $=\frac{\widetilde{\mathbf{s}_{f}}[T]+\widetilde{\mathbf{s}_{b}}[T]}{2},$ (38c)
where $\widehat{\mathbf{q}_{f}}\in\mathbb{R}^{T\times\mathcal{C}_{3}}$ and $\widehat{\mathbf{q}_{b}}\in\mathbb{R}^{T\times\mathcal{C}_{3}}$ represent the initial segments of the two sub-sequences, and $\mathbf{c}\in\mathbb{R}^{1\times\mathcal{C}_{3}}$ denotes the averaged feature from the final steps of both sub-sequences.
The ${Merge}()$ function then concatenates these elements:
$\widehat{\mathbf{s}}=Linear\Big{(}{Concat}\Big{(}\widehat{\mathbf{q}_{f}},\,\mathbf{c},\,\widehat{\mathbf{q}_{b}}\Big{)}\Big{)},$ (39)
where $\widehat{\mathbf{s}}$ resumes the original length of sequence, $\mathbb{R}^{p^{2}\times\mathcal{C}_{3}}$.
This approach effectively combines the two sub-sequences into a unified sequence while preserving a logical progression of features, which is particularly beneficial in models that process sequential data requiring a contextual understanding of feature evolution.
Figure 7: The visualization process of STL. The merged sequence feature is denoted by $\widehat{\mathbf{s}}$. After applying sequential attention to $\widehat{\mathbf{s}}$, the resulting features are represented as $\widehat{\mathbf{m}}$. The learnable transformation matrices $\mathbf{U}_{1}$ and $\mathbf{U}_{2}$ are utilized to extract length-level and dimension-level semantic features, denoted by $\mathbf{A}$ and $\mathbf{V}$, respectively. Finally, the down-sampled semantic tokens $\mathbf{u}$ are derived through the inner product of $\mathbf{A}$ and $\mathbf{V}$.
3.3.4 Semantic Token Learner (STL)
Efficiently condensing coarse features into fine and dynamic semantic features within a sequence remains a question. Drawing inspiration from recent advancements [26, 64], we introduce the STL in our study. STL is tailored to abstract and condense sequence input into a set of representative semantic tokens, utilizing a sequential attention mechanism to enhance the focus and relevance of features. Fig. 7 depicts an example of the transformation process facilitated by the STL.
For this process, the input feature sequence is denoted as $\widehat{\mathbf{s}}\in\mathbb{R}^{p^{2}\times\mathcal{C}_{3}}$, where $p$ is the patch size, and $\mathcal{C}_{3}$ is the dimensionality.
Initially, we adopt a sequential attention module, adapted for $\widehat{\mathbf{s}}$, utilizes a 1D convolution. This module emphasizes salient features within the $\widehat{\mathbf{s}}$, allowing for a focused analysis and feature extraction without 2D spatial dimensions.
It operates as follows:
$\widehat{\mathbf{s}}_{max}$ $=max(\widehat{\mathbf{s}}),$ (40a)
$\widehat{\mathbf{s}}_{ave}$ $=mean(\widehat{\mathbf{s}}),$ (40b)
where $\widehat{\mathbf{s}}_{max}\in\mathbb{R}^{p^{2}\times 1}$ and $\widehat{\mathbf{s}}_{ave}\in\mathbb{R}^{p^{2}\times 1}$ represent the max and average pooling features along the sequence, respectively.
Then, these features are concatenated and subjected to a 1D convolution followed by a sigmoid activation $Sigmoid()$ to produce a feature weighting map:
$\mathbf{m}=Sigmoid\Big{(}Conv1d\Big{(}Concat\Big{(}\widehat{\mathbf{s}}_{max},\,\widehat{\mathbf{s}}_{ave}\Big{)}\Big{)}\Big{)},$ (41)
where $\mathbf{m}\in\mathbb{R}^{p^{2}\times 1}$.
The weighting map is then applied to the original input:
$\widehat{\mathbf{m}}=\mathbf{m}\otimes\widehat{\mathbf{s}},$ (42)
yielding $\widehat{\mathbf{m}}\in\mathbb{R}^{p^{2}\times\mathcal{C}_{3}}$.
Subsequently, semantic tokens are initialized as $\mathbf{u}\in\mathbb{R}^{p_{2}^{2}\times\mathcal{C}_{3}}$, with $p_{2}$ being an odd number smaller than $p$ to indicate a reduced number of semantic tokens. The transformation of the weighted feature sequence $\widehat{\mathbf{m}}$ into semantic tokens is performed as follows:
$\mathbf{u}=\underbrace{Softmax\left(\left(\widehat{\mathbf{m}}\odot\mathbf{U}_{1}\right)^{\top}\right)}_{\mathbf{A}}\odot\underbrace{\Big{(}\widehat{\mathbf{m}}\odot\mathbf{U}_{2}\Big{)}}_{\mathbf{V}},$ (43)
where $\mathbf{U}_{1}\in\mathbb{R}^{\mathcal{C}_{3}\times p_{2}^{2}}$ and $\mathbf{U}_{2}\in\mathbb{R}^{\mathcal{C}_{3}\times\mathcal{C}_{3}}$ are weight matrix initialized using a Gaussian distribution. $\odot$ denotes the inner-product multiplication. This configuration allows $\widehat{\mathbf{m}}$ to map into two semantic groups effectively, with $\mathbf{A}\in\mathbb{R}^{p_{2}^{2}\times p_{2}^{2}}$ and $\mathbf{V}\in\mathbb{R}^{p^{2}\times\mathcal{C}_{3}}$ representing the resultant matrices in length-level and dimension-level which are then emphasized by a $Softmax()$ function for the most relevant semantic components. Finally, $\mathbf{u}$, the semantic tokens, is derived by inner-product from $\mathbf{A}$ and $\mathbf{V}$.
This STL aids in downsampling and captures crucial high-level features for downstream tasks.
Figure 8: The visualization process of STF. It illustrates the semantic feature fusion of $\mathbf{z}$ and $\mathbf{u}$. To ensure compatibility in size with $\mathbf{u}$, $\mathbf{z}$ is first adaptively pooled into $\overline{\mathbf{z}}$. Subsequently, the pooled feature $\overline{\mathbf{z}}$ is processed through a learnable transformation matrix $\mathbf{Z}$ to produce $\widehat{\mathbf{z}}$, which captures the influence score. This allows the correlation between $\overline{\mathbf{z}}$ and $\mathbf{u}$ to be assessed through the inner product of $\widehat{\mathbf{z}}$ and the semantic token $\mathbf{u}$. Finally, to integrate the token-influenced data and the attention-modified data, $\overline{\mathbf{u}}$ and $\overline{\mathbf{m}}$ are summed to yield the final output $\widehat{\mathbf{u}}$.
3.3.5 Semantic Token Fuser (STF)
To gate with the original input, previously prepared the $\mathbf{z}$ serves as the residual information for incorporating with encoded semantic features $\mathbf{u}$, as shown in Fig. 6. Instead of simply multiplying them together, we aim to fuse them by a STF operation to merge the information contained in a set of semantic tokens back into the initial sequence data, effectively fusing learned abstract representations with original sequence features. This is particularly useful in scenarios where you want to reinforce or adjust the sequence data using higher-level learned representations, which could be more informative or refined than raw features. Fig. 8 illustrates a sample of STF transformation process.
Given two features $\mathbf{u}\in\mathbb{R}^{p_{2}^{2}\times\mathcal{C}_{3}}$ and $\mathbf{z}\in\mathbb{R}^{p^{2}\times\mathcal{C}_{3}}$, we first pool the $\mathbf{z}$ into the same sequential length with $\mathbf{u}$:
$\overline{\mathbf{z}}={AdaptivePool}\Big{(}{SiLU}\Big{(}\mathbf{z}\Big{)}\Big{)}.$ (44)
After that, we compute the influence scores for each token based on the features in $\overline{\mathbf{z}}\in\mathbb{R}^{p_{2}^{2}\times\mathcal{C}_{3}}$:
$\widehat{\mathbf{z}}=Sigmoid\Big{(}\overline{\mathbf{z}}\odot\mathbf{Z}\Big{)}$ (45)
where $\widehat{\mathbf{z}}\in\mathbb{R}^{p_{2}^{2}\times p_{2}^{2}}$ and $\mathbf{Z}\in\mathbb{R}^{\mathcal{C}_{3}\times p_{2}^{2}}$ is a relevant parameter matrix. $Sigmoid()$ function ensures the scores range between 0 and 1.
Then, the tokens $\mathbf{u}$ are weighted by the influence scores $\widehat{\mathbf{z}}$ and summed over the tokens to blend the token information into the sequence:
$\overline{\mathbf{u}}=\widehat{\mathbf{z}}\odot\mathbf{u},$ (46)
where $\overline{\mathbf{u}}\in\mathbb{R}^{p_{2}^{2}\times\mathcal{C}_{3}}$, effectively spreading the token features across the sequence dimensions.
Next, the sequential attention $SeqAttn()$, mentioned in Eqs. 40a-42 is re-applied to $\overline{\mathbf{z}}$ to emphasize significant features:
$\overline{\mathbf{m}}=SeqAttn(\overline{\mathbf{z}}),$ (47)
where $\overline{\mathbf{m}}\in\mathbb{R}^{p_{2}^{2}\times\mathcal{C}_{3}}$ represents the attention-weighted sequence features.
Finally, The final sequence output integrates both the token-influenced data and the attention-modified data:
$\widehat{\mathbf{u}}=\overline{\mathbf{u}}+\overline{\mathbf{m}},$ (48)
where $\widehat{\mathbf{u}}\in\mathbb{R}^{p_{2}^{2}\times\mathcal{C}_{3}}$, providing a rich, detail-enhanced sequence for downstream tasks.
These operations ensure that the fused feature $\widehat{\mathbf{u}}$ is well-suited for downstream tasks, capturing both the detailed and global contextual information necessary for effective model performance. The addition of the attention mechanism particularly enhances the model’s ability to focus on relevant features dynamically based on the given input context.
Algorithm 1 Summary of procedure for the proposed method
1:  Input: Input HSI data $\mathbf{X}\in\mathbb{R}^{\mathcal{H}\times\mathcal{W}\times\mathcal{C}_{0}}$, ground truth $\mathbf{L}\in\mathbb{R}^{\mathcal{H}\times\mathcal{W}}$; patch size $p$; fixed training samples.
2:  Output: Predicted labels of the testing samples.
3:  Initialization: Optimizer: AdamW; Learning rate: 0.0005; Criterion: CrossEntropy; Epochs: 300; Batch size: 64.
4:  Procedure:
5:  Apply principal component analysis (PCA) on $\mathbf{X}\in\mathbb{R}^{\mathcal{H}\times\mathcal{W}\times\mathcal{C}}$ for dimension reduction, getting $\mathcal{C}_{1}$.
6:  Crop patch $\textbf{X}^{(i,\,j)}$ centered at pixel $\textbf{x}^{(i,\,j)}$ from samples.
7:  Perform depth-wise convolution on $\textbf{X}^{(i,\,j)}$ from feature dimension $\mathcal{C}_{1}$ to $\mathcal{C}_{2}$, generating local features.
8:  Perform point-wise convolution on $\textbf{X}^{(i,\,j)}$ then, serving as dynamic positional embedding.
9:  MiM model:
10:  Apply four types of MCS on $\textbf{X}^{(i,\,j)}$, getting sequences, $\mathbf{S}_{1}^{(i,\,j)}$, $\mathbf{S}_{2}^{(i,\,j)}$, $\mathbf{S}_{3}^{(i,\,j)}$, and $\mathbf{S}_{4}^{(i,\,j)}$ (Eqs. 14a-19a).
11:  for $m=1$ to $4$ in $\mathbf{S}_{m}^{(i,\,j)}$ do
12:     Split $\mathbf{S}_{m}^{(i,\,j)}$ into sub-sequences, $\mathbf{s}_{f}$ and $\mathbf{s}_{b}$ (Eq. 21).
13:     T-Mamba encoder:
14:     for Both forward pass $\mathbf{s}_{f}$ and backward pass $\mathbf{s}_{b}$ do
15:        Feed sub-sequences $\mathbf{s}_{f}$, $\mathbf{s}_{b}$ into respective Conv1d and S6 model, obtaining $\widetilde{\mathbf{s}_{f}}$ and $\widetilde{\mathbf{s}_{b}}$ (Eqs. 22a-23b).
16:        Apply GDMs to $\widetilde{\mathbf{s}_{f}}$ and $\widetilde{\mathbf{s}_{b}}$, resulting in $\widehat{\mathbf{s}_{f}}$ and $\widehat{\mathbf{s}_{b}}$ (Eqs. 24a and 24b).
17:        Merge two features $\widehat{\mathbf{s}_{f}}$ and $\widehat{\mathbf{s}_{b}}$, getting $\widehat{\mathbf{s}}$ (Eq. 25).
18:        Deploy STL on $\widehat{\mathbf{s}}$ to generate down-sampled semantic tokens $\mathbf{u}$ (Eq. 26).
19:        Apply STF on $\mathbf{u}$ and original feature $\mathbf{z}$ to generate enhanced tokens $\mathbf{\widehat{u}}$ (Eq. 27).
20:        Apply a residual connection between $\mathbf{\widehat{u}}$ and pooled $\mathbf{S}_{m}^{(i,\,j)}$, producing current output $\mathbf{o}_{m}^{(i,\,j)}$ (Eq. 28).
21:        Activate $\mathbf{o}_{m}^{(i,\,j)}$ using the Tanh function.
22:     end for
23:  end for
24:  Merge four outputs, $\mathbf{o}_{1}^{(i,\,j)}$, $\mathbf{o}_{2}^{(i,\,j)}$, $\mathbf{o}_{3}^{(i,\,j)}$, and $\mathbf{o}_{4}^{(i,\,j)}$, by WMF for weighted fusion, recorded as current scale ($p$) feature $\mathbf{O}_{p}^{(i,\,j)}$ (Eq. 49).
25:  Iterate MiM models on $\mathbf{o}_{1}^{(i,\,j)}$, $\mathbf{o}_{2}^{(i,\,j)}$, $\mathbf{o}_{3}^{(i,\,j)}$, and $\mathbf{o}_{4}^{(i,\,j)}$, until the patch size $p$ is reduced to $1$ (Eq. 52).
26:  Train the model using a decoder (MLP) and a multi-scale loss function (Eq. 53).
Table 1: Description of land-cover types and number of train-test samples for four datasets
Indian Pines (IP) Houston 2013 (HU2013)) Pavia University (PU) WHU-Hi-HongHu (HongHu)
No. Color Class Name Train Test Class Name Train Test Class Name Train Test Class Name Train Test
1 Corn-n. 50 1,384 Grass-h. 198 1,053 Asphalt 548 6,304 Red-r. 25 14,016
2 Corn-m. 50 784 Grass-s. 190 1,064 Meadow 540 18,146 Road 25 3,487
3 Corn 50 184 Grass-s. 192 505 Gravel 392 1,815 Bare-s. 25 21,796
4 Grass-p. 50 447 Tree 188 1,056 Tree 524 2,912 Cotton 25 163,260
5 Grass-t. 50 697 Soil 186 1,056 Metal-s. 265 1,113 Cotton-f. 25 6,193
6 Hay-w. 50 439 Water 182 143 Bare-s. 532 4,572 Rape 25 44,532
7 Soybean-n. 50 918 Residential 196 1,072 Bitumen 375 981 Chinese-c. 25 24,078
8 Soybean-m. 50 2,418 Commercial 191 1,053 Brick 514 3,364 Pakchoi 25 4,029
9 Soybean-c. 50 564 Road 193 1,059 Shadow 231 795 Cabbage 25 10,794
10 Wheat 50 162 Highway 191 1,036 Tuber-m. 25 12,369
11 Wood 50 1,244 Railway 181 1,054 Brassica-p. 25 10,990
12 Buildings-g.-t. 50 330 Park-l-1 192 1,041 Brassica-c. 25 8,929
13 Stone-s.-t. 50 45 Park-l-2 184 285 Small-b.-c. 25 22,482
14 Alfalfa 15 39 Tennis-g. 181 247 Lactuca-s. 25 7,331
15 Grass-p.-m. 15 11 Running-t. 187 473 Celtuce 25 977
16 Oat 15 5 Film-c.-l. 25 7,237
17 Romaine-l. 25 2,985
18 Carrot 25 3,192
19 White-r. 25 8,687
20 Garlic-s. 25 3,461
21 Broad-b. 25 1,303
22 Tree 25 4,015
Total 695 9,671 2,832 12,197 3,921 40,002 550 386,143
Ratio 6.7% 93.3% 18.8% 81.2% 8.9% 91.1% 0.14% 99.86%
Table 2: The hyperparameter settings and size of notations for four datasets.
Definition Notation IP HU2013 PU HongHu
Image Size: ×$\mathcal{H}\times\mathcal{W}\times\mathcal{C}_{0}$ $145\times 145\times 200$ $349\times 1905\times 144$ $610\times 340\times 103$ $940\times 475\times 270$
Dimension after PCA: $\mathcal{C}_{1}$ 60 30 30 100
Initial Patch Size: $p\times p$ $7\times 7$ $9\times 9$ $11\times 1$ 11 $9\times 9$
Feature Dimension Setting: $\mathcal{C}_{2},\,\mathcal{C}_{3},\,\mathcal{C}_{4}$ 64 64 32 128
Number of T-Mamba Layers: 4 4 2 3
Number of MiM Layers: 4 5 6 5
Learning Rate: 0.0005 0.0005 0.001 0.001
Feature Dropout: 0.1 0.1 0.2 0.1
3.4 Weighted MCS Fusion (WMF)
After obtaining four tokenized features, $\mathbf{o}^{(i,\,j)}_{1}$, $\mathbf{o}^{(i,\,j)}_{2}$, $\mathbf{o}^{(i,\,j)}_{3}$, and $\mathbf{o}^{(i,\,j)}_{4}$, from different T-Mamba encoders, we recognize that each feature set varies depending on the scan order of the sequence. This variability is critical in the context of our proposed MCS, which involves multiple scan orders, each providing a unique output representation. To effectively integrate these features, a thoughtful fusion strategy is essential, as indiscriminate merging could dilute their distinct contributions.
In response, we introduce the WMF, designed to assign attention weights to each feature set according to its relevance and importance in the overall context. The fusion process is formulated as follows:
$\mathbf{O}^{(i,\,j)}_{p}=k_{1}\mathbf{o}^{(i,\,j)}_{1}+k_{2}\mathbf{o}^{(i,\,j)}_{2}+k_{3}\mathbf{o}^{(i,\,j)}_{3}+k_{4}\mathbf{o}^{(i,\,j)}_{4},$ (49)
where $k_{o}\in\mathbb{R}^{1}$ (for $o\in\{1,2,3,4\}$) represents the weight assigned to each encoded feature, and $k_{1}+k_{2}+k_{3}+k_{4}=1$. $\mathbf{O}^{(i,\,j)}_{p}$ denotes the fused feature at scale $p$. These weights are dynamically integrated into the network architecture, allowing for optimization through learning, ensuring that each feature set contributes optimally to the final model output.
Following this feature fusion, the weighted feature $\mathbf{O}^{(i,\,j)}_{p}$ is processed through a decoder, specifically a Multi-Layer Perceptron (MLP), essential for refining and interpreting the fused features for model training:
$\mathbf{Y}^{(i,\,j)}_{p}=MLP\Big{(}Tanh\Big{(}\mathbf{O}^{(i,\,j)}_{p}\Big{)}\Big{)},$ (50)
where $Tanh()$ denotes the Hyperbolic Tangent activation function [^6] https://pytorch.org/docs/stable/generated/torch.nn.Tanh.html.
The training of the model is governed by the cross-entropy loss function, commonly used for classification tasks. This loss function measures the discrepancy between the predicted and actual labels of the training samples and is defined as:
$\mathcal{L}_{p}=-\frac{1}{N_{tr}}\sum^{N_{tr}}_{tr=1}\Big{[}\mathbf{L}_{tr}\log(\mathbf{Y}_{p})+(1-\mathbf{L}_{tr})\log(1-\mathbf{Y}_{p})\Big{]},$ (51)
where $N_{tr}$ is the total number of training samples, $\mathbf{L}_{tr}$ denotes the true label, and $\mathbf{Y}_{p}$ the predicted label for the $tr$-th sample. This function effectively minimizes the error between predicted and true labels, driving the training process towards better accuracy.
Table 3: Different classification methods in terms of OA, AA, and Kappa as well as the accuracies for each class on the Indian Pines dataset. The best one is highlighted as red. The second best is highlighted as blue.
No. Color. vanilla ViT SpeFormer MAEST HiT SSFTT 3DViT HUSST ViM MiM (Ours)
1 64.480 70.538 76.721 87.948 84.825 81.328 87.137 74.787 90.291
2 61.374 65.570 68.970 80.650 84.251 73.540 82.510 68.233 90.269
3 69.919 86.352 89.600 82.353 93.583 93.557 90.515 91.153 97.082
4 84.032 90.566 88.602 82.629 89.825 89.070 85.414 91.815 93.537
5 93.333 96.379 97.787 98.501 96.591 96.847 97.951 98.860 98.722
6 94.348 99.773 98.731 95.897 99.658 96.512 99.312 98.848 100
7 69.006 77.595 79.433 88.548 89.398 82.875 82.641 76.452 91.243
8 71.266 78.037 85.752 87.729 90.015 87.600 88.812 81.906 93.109
9 68.959 72.697 76.994 86.124 82.871 80.681 86.925 79.850 88.794
10 99.690 99.690 99.962 97.885 99.387 100 99.083 99.692 100
11 95.215 97.756 96.940 98.355 96.634 97.438 96.226 96.870 98.396
12 65.172 74.237 71.945 70.753 76.562 71.958 76.923 73.851 81.198
13 66.176 92.784 93.750 84.906 88.235 90.909 90.909 90.000 84.906
14 60.000 97.500 78.000 90.698 98.734 69.474 95.000 88.636 97.500
15 60.606 71.667 52.381 84.615 78.571 64.286 73.333 64.706 75.862
16 52.632 35.714 35.714 83.333 58.824 35.714 27.778 33.333 76.923
OA 74.0875 79.9607 83.0731 86.0511 88.7498 85.1411 87.3333 81.9667 92.0794
AA 77.6380 82.2704 84.7051 89.5578 90.9978 87.9868 89.0293 84.8289 94.1145
Kappa 0.7079 0.7733 0.8074 0.8420 0.8715 0.8305 0.8556 0.7952 0.9096
Table 4: Different classification methods in terms of OA, AA, and Kappa as well as the accuracies for each class on the Pavia University dataset. The best one is highlighted as red. The second best is highlighted as blue.
No. Color. vanilla ViT SpeFormer MAEST HiT SSFTT 3DViT HUSST ViM MiM (Ours)
1 87.245 89.487 90.084 90.559 96.721 95.985 96.443 78.878 97.255
2 74.871 90.862 92.186 91.322 89.994 88.007 87.622 85.758 92.947
3 60.796 78.886 79.862 66.773 83.342 88.233 75.083 46.655 87.009
4 76.601 88.610 90.488 94.595 94.804 97.830 95.025 90.010 96.251
5 99.194 100 100 98.670 100 100 100 99.326 100
6 64.266 75.571 82.845 87.232 73.704 68.196 74.182 71.951 80.032
7 95.694 80.966 89.988 84.953 98.616 99.217 94.418 77.917 99.128
8 82.894 83.960 84.277 90.554 88.452 90.886 90.376 76.343 91.252
9 94.375 99.949 99.310 95.719 92.159 93.173 99.619 95.831 100
OA 76.0313 86.2784 88.0433 87.9183 88.8707 86.5258 86.2483 78.1811 91.5756
AA 84.4339 89.8503 90.5291 90.7538 91.4693 91.3783 90.6566 78.5284 92.7601
Kappa 0.7086 0.8288 0.8341 0.8449 0.8608 0.8581 0.8572 0.7145 0.8983
3.5 Multi-Scale Loss Design
To account for the multi-scale features potentially present in an image patch for different objects, we propose integrating multiple scale features into the decoding process.
In our approach, each MiM model performs a down-sampling operation on the input image patch, which is initially processed by the STL in the T-Mamba encoder.
Upon obtaining four distinct features, $\mathbf{o}^{(i,\,j)}_{1}$, $\mathbf{o}^{(i,\,j)}_{2}$, $\mathbf{o}^{(i,\,j)}_{3}$, and $\mathbf{o}^{(i,\,j)}_{4}$, from the first MiM model at the initial patch scale $p$, each feature is activated and subsequently fed into the next MiM model for further feature tokenization and down-sampling, as shown in the Fig. 3.
This process allows us to capture multi-scale features in a cascaded and iterative manner, which can be analogous as:
$\mathbf{O}^{(i,\,j)}_{p_{n}}=\underbrace{MiM(MiM(MiM(\cdots MiM(\mathbf{O}^{(i,,j)}_{p}))))}_{n\,times},$ (52)
where $p_{n}<p$, recording the patch size at $n$-th $MiM()$ iteration. It must be noted that all patch sizes should be an odd number. The iteration is ending at $p_{n}=1$. For example, $p=11$, $p_{2}=9$, $p_{3}=7$, $p_{4}=5$, $p_{5}=3$, and $p_{6}=1$ by 6 times of operation under MiM model.
To effectively utilize these multi-scale features, we employ an output-level based decoding strategy, where each scale feature is independently integrated into the loss function. Consequently, as per Eq. 51, we formulate the total loss function as:
$\mathcal{L}_{total}=\frac{1}{n}(\mathcal{L}_{p}+\mathcal{L}_{p_{2}}+\mathcal{L}_{p_{3}}+\cdots+\mathcal{L}_{p_{n}}).$ (53)
The iterative process concludes when the final patch size $p_{n}$ is reduced to 1. Thus, our total loss function effectively incorporates the multi-scale features from a single image patch to enhance training effectiveness.
Figure 9: Four experimental datasets, (a) Indian Pines (IP), (b) Pavia University (PU), (c) Houston 2013 (HU2013), and (d) WHU-Hi-HongHu (HongHu) are used for experiments. The false-color images, fixed and disjoint training sets and testing sets are shown above.
4 Experimental results and analysis
In this section, we first describe four well-known HSI datasets. We then outline the implementation details and the methods used for comparison. Finally, we conduct extensive experiments to evaluate the performance of the proposed method both quantitatively and qualitatively. Due to the paper formatting, all evaluation tables are placed at end of the article.
4.1 Datasets Description
Four HSI datasets are illustrated in Fig. 9, along with their corresponding false-color images, training and testing samples. The numbers of training and testing samples for each class within these datasets are listed in Table 1.
The Indian Pines (IP) dataset is a mixed vegetation site with $145\times 145$ pixels and 220 spectral bands. To remove water absorption bands, the number of spectral bands was reduced to 200, resulting in 16 different land-cover classes.
The Pavia University (PU) dataset consists of $610\times 340$ pixels and initially has 115 bands. In the experiment, 12 of the noisy bands were removed, leaving 103 bands to be used. The dataset has 9 different ground cover types.
The Houston 2013 (HU2013) dataset [^7] http://www.grss-ieee.org/community/technical-committees/data-fusion was used in the 2013 IEEE GRSS Data Fusion Contest. It features 144 spectral bands and spans an area of $349\times 1905$ pixels, covering 15 different land-cover classes.
The WHU-Hi-HongHu (HongHu) dataset [^8] http://rsidea.whu.edu.cn/resource_WHUHi_sharing.htm was acquired by the UAV platform on November 20, 2017, in Honghu City, Hubei Province, China. The experimental area is a complex agricultural scene with many classes of crops, and different cultivars of the same crop are also planted in the region. It consists of $940\times 475$ pixels and initially has 270 bands. This dataset has 22 different ground cover classes.
For fair comparison, all the training and testing samples are manually selected in a fixed and disjoint manner. Details can be referred to [19] and corresponding website.
4.2 Experimental Evaluation Indicators
To quantitatively assess the effectiveness of our proposed method and compare it with other methods, we rely on four key performance indicators: overall accuracy (OA), average accuracy (AA), the kappa coefficient (kappa), and accuracy within each class. They provide a comprehensive view of the classification performance, with higher values indicating superior classification results.
All experiments are conducted using Pytorch 2.3 with a Geforce RTX 3090 GPU and Ubuntu 22.04 system.
Table 5: Different classification methods in terms of OA, AA, and Kappa as well as the accuracies for each class on the Houston 2013 dataset. The best one is highlighted as red. The second best is highlighted as blue.
No. Color. vanilla ViT SpeFormer MAEST HiT SSFTT 3DViT HUSST ViM MiM (Ours)
1 89.129 90.815 89.644 89.616 91.085 95.123 92.073 87.271 94.887
2 94.852 93.950 91.537 92.643 93.056 97.550 94.473 93.035 98.273
3 96.188 97.087 100 99.901 98.232 100 100 97.527 100
4 93.285 96.843 95.083 97.016 94.711 97.681 96.279 87.680 96.866
5 96.445 94.575 96.748 97.959 96.925 97.823 96.378 93.160 99.858
6 74.011 93.464 93.770 91.374 84.713 99.306 97.260 94.595 99.652
7 82.431 91.276 90.431 91.541 87.516 90.009 91.864 84.735 91.423
8 74.693 68.837 75.459 81.531 88.889 72.029 76.727 84.537 87.167
9 73.527 86.340 85.249 84.658 88.835 85.297 86.209 85.089 84.309
10 72.931 71.959 86.763 88.628 91.663 82.154 81.541 84.514 92.380
11 69.631 75.279 86.429 84.187 87.778 87.748 92.630 83.675 90.028
12 70.809 78.114 75.945 82.155 88.132 78.154 81.899 79.158 89.581
13 63.805 90.429 91.667 91.501 90.203 92.632 90.657 87.896 93.080
14 96.509 97.436 100 93.690 97.041 95.367 99.597 99.190 97.436
15 98.943 95.075 96.728 96.629 91.245 97.399 95.122 97.612 98.333
OA 82.9548 86.4557 88.8415 90.0303 91.4241 89.7024 90.1205 87.5871 92.8917
AA 84.0284 88.3335 89.8294 91.7086 93.0019 90.1268 93.0002 89.1248 94.2108
Kappa 0.8151 0.8530 0.8787 0.8918 0.9069 0.8882 0.8928 0.8653 0.9228
4.3 Brief Description and Settings of Compared Methods
We conducted comparative analyses against a range of baseline and state-of-the-art models. These included vanilla ViT (ViT) [18], SpeFormer [19], MAEST [21], HiT [22], SSFTT [26], 3DViT [27], HUSST [28], and ViM [51]. All models were employed as patch-wise classifiers, with parameter settings adhering to their respective references, especially the initial patch size for different datasets. The comparison is grounded in the same number of the training samples without data augmentation.
  • 1)
    ViT is configured with a feature dimension of 64 and incorporates 4 multi-heads self-attention. The network depth is set to 4 layers. Across all datasets, the initial size of the cropped patch is $9\times 9$. We use a sub-patch size of $3\times 3$, partitioned in a non-overlapping manner.
  • 2)
    SpeFormer integrates group-wise spectral embedding into the Transformer encoder and includes cross-layer adaptive fusion modules to boost the encoder’s learning capabilities. It is configured with an initial patch size of $7\times 7$ across all datasets in our experiments.
  • 3)
    MAEST integrates a reconstruction path with a classification path for HSI classification, employing masking auto-encoding (MAE). The initial patch size is set at $7\times 7$.
  • 4)
    HiT embeds convolution operations into the Transformer structure to capture the subtle spectral discrepancies and convey the local spatial context information. For all datasets, the initial patch size is set at $15\times 15$ for HiT.
  • 5)
    SSFTT combines 3D-CNNs and 2D-CNNs, featuring a Gaussian-weighted feature tokenizer to prepare extracted features for the Transformer encoder. In our experiments, patch sizes vary by dataset: $13\times 13$ for the IP and PU dataset and $9\times 9$ for both HU2013 and HongHu dataset.
  • 6)
    3DViT integrates spectral and spatial features into sub-cubes using a 3D configuration within the ViT framework. It implements 3D coordinate positional embedding and a local-global feature fusion mechanism, merging local details with global context. For all datasets, the initial patch size for the 3DViT model is set at $9\times 9$.
  • 7)
    HUSST implements a cascaded approach, combining a spectral Transformer with a spatial local-global Transformer hierarchically. It introduces cross-scale local-global feature fusion for multi-scale feature aggregation. The initial patch size set at $11\times 1$ 11 for all datasets.
  • 8)
    ViM, the first Mamba model tailored for vision tasks, is applied to HSI classification to assess its viability. The hyperparameters are aligned with those of our proposed MiM model, detailed later.
  • 9)
    Our MiM model standardizes all feature dimensions at 64, with T-Mamba encoders having a depth of 2. The initial patch sizes are set at $7\times 7$ for the IP dataset, $11\times 1$ 11 for the PU dataset, and $9\times 9$ for the HU2013 and HongHu dataset. Others can be found in Algorithm 1.
Table 6: Different classification methods in terms of OA, AA, and Kappa as well as the accuracies for each class on the WHU-Hi-HongHu dataset. The best one is highlighted as red. The second best is highlighted as blue.
No. Color. vanilla ViT SpeFormer MAEST HiT SSFTT 3DViT HUSST ViM MiM (Ours)
1 83.308 92.414 93.150 92.397 94.062 93.913 93.392 90.774 91.922
2 51.588 77.377 69.992 83.118 80.321 84.327 82.783 78.092 83.239
3 66.129 87.049 87.753 86.075 86.447 88.706 86.066 85.647 89.772
4 81.673 92.768 95.519 95.807 92.688 93.738 94.840 90.939 95.993
5 30.789 42.216 52.979 50.280 39.060 45.475 59.836 40.896 59.971
6 80.313 93.287 94.380 94.295 93.865 93.009 94.936 92.830 96.003
7 55.159 69.221 81.213 66.949 84.649 74.366 79.445 81.871 83.873
8 29.825 48.326 45.361 47.745 46.911 64.771 45.650 54.258 64.170
9 87.712 95.937 95.455 93.624 96.133 94.108 95.912 95.570 96.399
10 50.244 78.034 85.437 78.864 82.318 85.436 83.367 77.348 87.094
11 32.714 68.521 80.407 64.692 76.228 82.055 72.825 75.274 88.261
12 38.737 60.166 72.607 64.964 76.981 65.600 69.342 68.752 77.643
13 41.373 66.599 75.962 60.533 72.911 70.993 75.059 66.550 76.611
14 53.612 75.796 88.200 85.221 85.478 90.432 86.554 74.109 89.186
15 48.922 80.839 84.730 83.587 84.911 70.855 72.665 37.689 87.631
16 85.969 96.110 94.660 94.505 95.649 95.244 96.007 93.233 95.948
17 71.222 89.171 74.695 94.118 85.529 87.956 87.207 85.676 92.016
18 49.649 84.597 83.587 85.643 83.333 84.548 86.010 89.097 86.875
19 59.138 83.529 86.927 82.922 90.572 86.385 88.418 86.205 91.872
20 34.296 71.906 80.832 32.043 79.922 80.556 82.181 74.823 88.583
21 20.711 50.048 52.946 52.932 46.456 62.122 44.932 56.133 71.989
22 32.600 66.921 80.489 87.111 68.578 83.324 67.961 53.190 83.057
OA 65.2688 82.9934 87.3104 84.2750 85.2845 85.8091 86.9473 82.7789 89.5749
AA 67.4351 84.6977 85.7987 83.7510 87.4042 88.7126 87.9006 85.7671 90.9751
Kappa 0.5873 0.7911 0.8427 0.8048 0.8196 0.8254 0.8387 0.7897 0.8705
Figure 10: The classification maps with different methods for the IP dataset. (a) vanilla ViT (OA=74.0875%). (b) SpeFormer (OA=79.9607%). (c) MAEST (OA= 83.0731%) (d) HiT (OA=86.0511%). (e) SSFTT (OA=88.7498%). (f) 3DViT (OA=85.1411%). (g) HUSST (OA=87.3333%). (h) ViM (OA= 81.9667%). (i) MiM (OA=92.0794%).
Figure 11: The classification maps with different methods for the PU dataset. (a) vanilla ViT (OA=76.0313%). (b) SpeFormer (OA=86.2784%). (c) MAEST (OA=88.0433%) (d) HiT (OA=87.9183%). (e) SSFTT (OA=88.8707%). (f) 3DViT (OA=86.5258%). (g) HUSST (OA=86.2483%). (h) ViM (OA=78.1811%). (i) MiM (OA=91.5756%).
Figure 12: The region of interest (ROI) in classification maps with different methods for the HU2013 dataset. (a) vanilla ViT (OA=82.9548%). (b) SpeFormer (OA=86.4557%). (c) MAEST (OA=88.8415%) (d) HiT (OA=90.0303%). (e) SSFTT (OA=91.4241%). (f) 3DViT (OA=89.7024%). (g) HUSST (OA=90.1205%). (h) ViM (OA=87.5871 %). (i) MiM (OA=92.8917%).
Figure 13: The classification maps with different methods for the HongHu dataset. (a) vanilla ViT (OA=65.2688%). (b) SpeFormer (OA=82.9934%). (c) MAEST (OA=87.3104%) (d) HiT (OA=84.2750%). (e) SSFTT (OA=85.2845%). (f) 3DViT (OA=85.8091%). (g) HUSST (OA=86.9473%). (h) ViM (OA=82.7789%). (i) MiM (OA=89.5749%).
4.4 Experimental Results
4.4.1 Quantitative Results and Classification Maps
Tables 3, 4, 5, and 6 present average classification metrics under 5 times of running, such as OA, AA, Kappa coefficient, and class-wise accuracy for the IP, PU, HU2013, and HongHu datasets, respectively. Corresponding classification maps are displayed in Figs. 10, 11, 12, and 13.
From these results, it’s evident that baseline methods like vanilla ViT and ViM do not yield satisfactory outcomes, especially the vanilla ViT, which struggles due to its purely parallel and simultaneous processing approach (e.g., self-attention). This often results in noise-like classification artifacts, particularly noticeable in smooth regions and at the boundaries of land-covers, as the classification map for vanilla ViT exhibits much sharper visual artifacts. Conversely, ViM, which integrates bidirectional sequence processing with Mamba encoders, typically shows better performance. This improvement could be attributed to an ordering bias, akin to that found in RNNs, which is advantageous when training samples are limited.
SpeFormer and MAEST, enhancements over vanilla ViT, achieve significantly better results. SpeFormer introduces local feature aggregation within spatial and spectral domains and incorporates cross-layer connections to enhance the robustness of its encoder. Compared to vanilla ViT, SpeFormer produces smoother classifications in land-cover regions. MAEST integrates a masked auto-encoder into SpeFormer, adding a noise mask to the original data and reconstructing it, which enhances generative performance on unseen data, especially when training and testing samples are uncorrelated. MAEST presents better improvements on SpeFormer.
On another front, 3DViT extends ViT into a 3D sequential model across spatial and spectral domains, effectively linking spatial and spectral features to enhance classification accuracy. Despite being a pure Transformer-based model without hybrid elements like RNNs or CNNs, it shows promising results.
Moreover, HiT, SSFTT, and HUSST combine CNNs with Transformers to leverage local and global feature awareness, generally yielding superior results. These models provide clearer distinctions in flat land-cover regions and at boundaries. However, their larger model sizes contribute to higher computational demands and are less hardware-friendly.
Table 7: Total running time (minutes) and the number of parameters required by different methods on the Indian Pines dataset.
vanilla ViT SpeFormer MAEST HiT SSFTT 3DViT HUSST ViM MiM (Ours)
Train (minutes.) 2.1 3.5 4.7 5.6 5.5 4.5 3.5 2.5 3.6
Test (seconds.) 2.0 4.0 5.0 11.0 5.0 8.0 3.0 7.0 5.0
Parameters 416K 644K 953K 45.8M 2.9M 2.4M 9.1M 357K 1.1M
Our method, MiM, introduces a MCS mechanism with a T-Mamba encoder, facilitating lightweight, multi-directional feature generation. Compared to ViM, our specialized T-Mamba encoder and extended bidirectional scan improve classification performance significantly. Our approach, notably lighter than HiT, SSFTT, and HUSST, produces classification maps with smoother regions and more distinct land-cover boundaries. The class-wise accuracies are competitive and generally superior to other methods.
Key differences in classification performance are highlighted and marked in Figs. 10, and 11 for detailed comparisons.
This exploration paves the way for our future research into adapting pure Mamba models and MCS into a 3D configuration, similar to 3DViT which extends ViT into 3D version.
Figure 14: Visualization of T-SNE results on the IP test samples by different methods. (a) original test samples’ T-SNE results. (b) vanilla ViT. (c) SpeFormer. (d) MAEST. (e) HiT. (f) SSFTT. (g) 3DViT. (h) HUSST. (i) ViM. (j) MiM.
4.4.2 Learned Feature Visualizations by T-SNE
Deep learning-based methods excel in creating new feature representations that replace the original inputs through representation learning. Typically, these methods train a discriminative model to transform input data into a distinct output representation for tasks such as HSI classification, subsequently utilizing these new features for clustering.
To evaluate the representational capabilities of various methods, we employed t-distributed stochastic neighbor embedding (T-SNE) [^9] https://en.wikipedia.org/wiki/T-distributed_stochastic_neighbor_embedding, a tool that visualizes the distribution of learned features from test samples.
Using the IP dataset as an example, Fig. 14 illustrates the visualization outcomes of different methods. From this, two key points emerge. Firstly, while some clusters represented by different colors may appear closely positioned in T-SNE results, this proximity should not be interpreted as indicative of similarity due to T-SNE’s internal mechanics. Secondly, the presence of same-color clusters at considerable distances from each other highlights T-SNE’s focus on preserving local structure while optimizing a non-convex function, which may lead to local minima.
The visualization clearly demonstrates that our MiM model effectively discriminates between learned feature representations. Through T-SNE, the new feature representations not only cluster but also show increased discriminative properties, with several clusters unifying under the same color and appearing smoother.
4.4.3 Comparison of Computational Costs
To evaluate computational efficiency across all methods, we conducted a detailed statistical analysis, summarized in Table 7. This analysis adheres to previously described experimental setups and includes metrics such as total training time (minutes), testing time (seconds), and model size (number of parameters) for the IP dataset.
Our method showcases competitive processing times, significantly outperforming advanced models like HiT, MAEST, 3DViT, and SSFTT in terms of efficiency. However, it records slightly longer times compared to simpler models such as vanilla ViT, SpeFormer, HUSST, and ViM. When it comes to the number of trainable parameters, our approach proves to be more efficient, requiring fewer parameters than other leading methods, which underscores its parameter efficiency.
In summary, our method strikes an effective balance between running time and model complexity, measured by parameter count. While it doesn’t achieve the fastest training or testing times, it offers lower operational times than several sophisticated models and maintains a manageable parameter count. This combination of efficiency and simplicity makes our method particularly suitable for scenarios where both performance and model manageability are priorities.
4.5 Other Analysis and Ablation Study
In this part, a comprehensive analysis, experimental evaluation, and ablation studies have been presented.
Table 8: Impact of different insertable components in T-Mamba encoder on the overall classification results (%). The best one is highlighted as red.
$STL$ $GDM$ $STF$ IP PU HU2013 HongHu
82.559 80.999 88.112 84.123
86.129 85.698 90.995 86.889
84.006 82.159 89.459 86.114
84.561 83.226 90.642 85.013
87.021 86.111 91.324 88.546
91.969 89.123 91.452 88.982
87.159 87.661 91.003 88.002
92.079 91.576 92.892 89.575
4.5.1 Ablation on Proposed Different Components
In the T-Mamba encoder, we integrated several key components: the STL, GDM, and STF. To assess their effectiveness, we conducted an ablation study, the results of which are summarized in Table 8. The ablation study details how each component was tested. In scenarios where STL and STF were not used, we substituted them with general average pooling and simple summation, respectively, to gauge their relative performance.
STL outperforms average pooling in merging MCS features, suggesting its superior ability to capture contextual information within the merged features, significantly enhancing accuracy.
GDM refines feature focus after Mamba encoding by reweighting the sequence based on the proximity to the central pixel, which is the last index in a sub-sequence. This targeted approach enhances the central pixel’s prominence, reducing the impact of interfering pixels and yielding modest improvements.
STF fuses enhanced features more effectively than simple summation, leveraging higher-level learned representations to enrich the raw feature set. Incorporating STF alone leads to better outcomes than summation, underscoring its utility in strengthening feature representations.
Additionally, combining STL and STF yielded the most substantial improvements, likely due to their complementary mechanisms both utilizing sequential attention.
Overall, this ablation study demonstrates that the integration of STL, GDM, and STF significantly boosts performance by synergistically refining the handling of image band features in a sequential model, particularly with the continuous scanning features critical in our proposed MiM model.
Table 9: Impact of different initial patch size for the overall accuracy (%) on four datasets. The best one is highlighted as red.
Patch Size ($p$) IP PU HU2013 HongHu
$3\times 3$ 89.001 $\uparrow$ 86.331 $\uparrow$ 89.123 $\uparrow$ 83.127 $\uparrow$
$5\times 5$ 91.332 $\uparrow$ 89.126 $\uparrow$ 91.028 $\uparrow$ 87.269 $\uparrow$
$7\times 7$ 92.079 $\uparrow$ 90.214 $\uparrow$ $9$uparrow 89.113 $\uparrow$
$9\times 9$ 91.554 $\downarrow$ 91.002 $\uparrow$ 92.892 $\uparrow$ 89.575 $\uparrow$
$11\times 1$ 11 90.667 $\downarrow$ 91.576 $\uparrow$ 92.091 $\downarrow$ 88.990 $\downarrow$
$13\times 13$ 89.214 $\downarrow$ 90.885 $\downarrow$ 90.451 $\downarrow$ 87.317 $\downarrow$
$15\times 15$ 87.332 $\downarrow$ 88.213 $\downarrow$ 88.519 $\downarrow$ 86.451 $\downarrow$
4.5.2 Analysis on Initial Patch Size
We investigated the effect of initial patch size on classification accuracy. For instance, with the IP dataset, the optimal starting patch size is 7 $\times$ 7. This setup allows us to derive four scale features—scale 7, 5, 3, and 1—each contributing to the loss function. Therefore, the final result is obtained by training four loss functions together.
As shown in Table 9, we evaluated a range of initial patch sizes (3, 5, 7, 9, 11, 13, 15). The results, specifically OA for four datasets, are detailed in the same table.
An appropriately large initial patch size can improve model performance by capturing more spatial context. However, excessively large patch sizes can be detrimental, introducing noise from interfering pixels.
For the relatively compact IP dataset, the ideal initial patch size is small, confirmed at 7 $\times$ 7. In contrast, for the HU2013 dataset, which features a dense spatial distribution but closely spaced land-covers, a slightly larger patch size of 9 $\times$ 9 is optimal. The PU dataset, characterized by its flat and smooth terrain, warrants an even larger optimal initial patch size of 11 $\times$ 11. Finally, the HongHu dataset gets the ideal initial patch size at 9 $\times$ 9.
4.5.3 Investigation and Comparison on Scan Design
In the proposed centralized MCS, selecting the appropriate scan manner is essential for optimal feature allocation and positioning within sequential models. It is critical to choose a cross-scan design that effectively spans spatial domains.
Given that random and casual scanning may lead to significant redundancy and poor feature arrangement, we established a certain rule: spatially, the scanning should keep continuous as much as possible. Thus, through investigation, we identified several scanning designs, as shown in Fig. 15. For each scan design, it can be extended into four types, similar to those shown in Fig. 5.
The Mamba-scan achieves the greatest continuity when converting an image into a sequence. The Raster-scan, a common method, linearizes the image into a sequence but introduces jump connections at the boundaries of the image. The Diagonal-scan approaches this by scanning images diagonally, selecting positions along the diagonals. Conversely, the Zig-zag-scan modifies the diagonal approach by maintaining partial continuity at image boundaries.
Figure 15: Investigated scan designs. (a) proposed Mamba-scan design. (b) general Raster-scan design. (c) Diagonal-scan design. (d) Zig-Zag-scan design. Each design can be extended into four types from different start and end positions, similar to those shown in Fig. 5.
Table 10: The comparison of using different scan designs for OA (%). The best one is highlighted as red. The second best is highlighted as blue.
Num. of used IP PU HU2013 HongHu
1 89.123 89.559 90.123 87.315
Mamba-scan 2 90.156 90.452 91.029 88.569
3 91.221 91.001 91.993 89.226
4 92.079 91.576 92.892 89.575
1 88.326 87.009 89.163 85.429
Raster-scan 2 89.555 88.322 90.442 86.129
3 90.122 89.125 91.105 88.226
4 91.612 90.761 91.986 89.125
1 88.039 86.002 89.013 83.664
Diagonal-scan 2 89.196 87.569 89.669 86.548
3 90.236 88.555 90.139 87.841
4 91.002 89.571 91.004 88.221
1 88.129 87.563 87.125 84.003
Zig-Zag-scan 2 90.009 87.993 89.699 86.443
3 91.011 88.961 90.861 87.231
4 91.536 91.299 91.834 87.992
We tested these scanning methods to evaluate their performance, with results summarized in Table 10. Each scan design was tested using one to four types. The most effective results were achieved using the Mamba-scan with four directional types, confirming that a continuous sequence arrangement enhances content recognition. This finding is further supported by comparing the Diagonal-scan and the Zig-zag-scan, where the latter showed slight improvements due to its semi-continuous positioning. The commonly used Raster-scan, despite its prevalence in image transformation tasks, tends to disrupt spatial continuity, particularly when spatial context is significant in RS land-cover analysis, resulting in inferior performance compared to the Mamba-scan.
Additionally, employing four types of scans enhances each scanning approach, demonstrating the viability of multi-directional feature learning.
4.5.4 Exploring the Feature Fusion Strategy in T-Mamba
We developed our bi-directional feature fusion strategy in T-Mamba encoder with two ways, as shown in Fig. 16: 1) Post-Merge strategy, and 2) Pre-Merge strategy. To evaluate the effectiveness of these two strategies, we conducted experiments across four datasets. The results of these experiments are detailed in Table 11.
Table 11: Impact of different feature fusion strategy in T-Mamba encoder for the OA (%) on four datasets. The best one is highlighted as red.
IP PU HU2013 HongHu
Post-Merge 91.326 90.332 91.229 88.450
Pre-Merge 92.079 91.576 92.892 89.575
Figure 16: Designed bi-directional feature fusion manner in T-Mamba encoder. (a) Post-Merge strategy, and (b) Pre-Merge strategy.
We observed that the Pre-Merge strategy tends to be more effective for feature condensation and aggregation. This is likely because the STL module is applied to the merged features, which encapsulate the entirety of the image’s features. Consequently, the STL can comprehensively analyze the whole image’s contents, facilitating the condensation and abstraction of features, akin to a ’global’ awareness of the image. Subsequently, the STF module leverages the pooled feature of the entire image, $\mathbf{z}$, to achieve a holistic understanding of image features. In contrast, the Post-Merge strategy, applies the STL and STF to each branch independently without integrating semantic information across the branches. This approach results in less effective feature integration compared to the Pre-Merge strategy.
Therefore, rather than directly adopting the ViM model, separating two individual branches, for our HSI classification task, we have developed a more suitable and innovative approach that enhances performance through strategic feature merging and processing.
Figure 17: The comparison between using different sequential networks (i.e., RNN, GRU, LSTM, Transformer encoder, RT, Mamba, and T-Mamba). We have to note that all networks are not sharing parameters for four types of MCS in MiM model. Experimental results are listed in (a) IP dataset, (b) PU dataset, (c) HU2013 dataset, and (d) HongHu dataset.
Figure 18: The comparison on the training time, OA(%) and model size by using different sequential networks, aligning with Fig. 17.
4.5.5 Comparison with Other Sequential Networks
Following the development of the T-Mamba, we conducted comparative experiments against other established sequential networks: RNN, GRU, LSTM, general Transformer encoder, RT (a hybrid of LSTM and Transformer) [44], and a general Mamba model [46]. These models all process vector inputs at each timestep, allowing us to integrate them directly into our system for a straightforward comparison. We have to note that all networks are not sharing parameters for four types of MCS in MiM model. All tests were performed under identical conditions, with the results shown in Figs. 17 and 18.
The vanilla RNN displayed the lowest performance, constrained by its basic approach to processing sequential data. Both GRU and LSTM showed similar results, but the GRU was more efficient due to its simpler structure with one fewer gate than the LSTM. The RT model, which merges LSTM’s sequential processing with the Transformer’s self-attention, exhibited enhanced performance, underscoring the benefits of combining these two approaches. Although RT outperformed the standalone LSTM and GRU, it requires a significantly larger parameter set.
Our newly designed T-Mamba, which enhances the basic Mamba encoder with tokenization capabilities, does add some parameters but achieves considerable improvements in classification performance. In comparison to the other models, T-Mamba strikes an optimal balance between model size and efficacy. It surpasses the general Transformer encoder in efficiency, offering a more lightweight and robust solution. While RT provides competitive performance, its larger model size and higher resource consumption are less than ideal, making T-Mamba the more advantageous choice in terms of efficiency and effectiveness.
Table 12: Different classification methods in terms of OA on the Indian Pines dataset with data augmentation. The best one is highlighted as red. The second best is highlighted as blue.
Num. Train vanilla ViT SpeFormer MAEST HiT SSFTT 3DViT HUSST ViM MiM
Scena. 1 1024 85.163 94.123 97.323 97.211 97.002 97.328 96.892 94.317 97.332
Scena. 2 4170 90.669 96.128 99.002 98.386 98.991 98.001 97.669 95.659 98.239
4.5.6 Further Discussion on Mamba and Transformer under Data Augmentation
Mamba was introduced as an alternative to the Transformer for efficient and lightweight sequential feature encoding. However, unlike the Transformer, which employs global attention, Mamba processes sequences in a linear-like fashion. The Transformer’s self-attention mechanism has gained widespread adoption across various applications due to its ability to generalize well with sufficient data, making extensive training datasets a necessity for optimal performance. Against this backdrop, we conducted further experiments to evaluate Mamba-based methods against Transformer-based approaches.
These experiments centered on the augmentation of training samples. We established two scenarios: 1) using a general sample of 10% randomly selected from the ground truth, a common practice in HSI classification, and 2) augmenting the data through geometric transformations—45°, 90°, and 135° rotations, and vertical and horizontal flips—as outlined in [62].
Taking the IP dataset as an example, where labeled pixels across land-cover categories are typically imbalanced and scarce, we initially selected 1024 training samples in the scenario-1, significantly more than in previous experiments with disjoint and fixed training sets. For the augmented scenario-2, we applied geometric transformations to disjoint samples, yielding a substantially increased pool of 4170 training samples (695 multiplied by 6).
The results, detailed in Table 12, indicate that increasing the training sample size to nearly double improves the performance of all models, including our Mamba-based model, although the accuracy gap between methods narrows considerably. This is further evidenced by the shrinking difference in performance between SpeFormer and ViM, two purely model-driven approaches.
However, when the number of training samples are expanded six-fold to 4170, our MiM model was no longer the top performer. Instead, MAEST, which integrates the MAE with prior masks, achieved the best results. Transformer-based models generally performed well. This improvement is likely due to the Transformer’s mechanism, which effectively utilizes ample training samples to enhance its recognition and generalization capabilities, as supported by literature suggesting that Transformers can outperform CNNs when sufficient data is available.
Thus, while our MiM model excels in scenarios with limited, disjoint, and fixed samples—conditions that mimic real-world HSI applications. The Transformer-based methods tend to perform better under conditions of abundant training data.
5 Conclusion
In this study, we introduced a novel and pioneering approach to HSI classification utilizing the newly developed Mamba architecture. Specifically, we implemented a centralized MCS for multi-directional feature aggregation, creating a lighter and more streamlined framework for visual cognition alongside the Mamba sequential model. To further enhance this architecture, we designed a T-Mamba encoder, incorporating bi-directional GDM and a STL to refine feature quality and support downsampling-driven semantic token representation. Meanwhile, porposed STF is effectively fusing learned abstract representations with original sequence features. Additionally, we integrated a WMF module and a multi-scale loss function to optimize decoding efficiency. As the pioneering application of a vision Mamba-based model for HSI classification, our results have significantly outperformed other baselines and state-of-the-art methods with limited and disjoint data samples. However, with augmented training samples, Transformer-based models tend to be better. Moving forward, we plan to extend our idea to include a 3D version of the Mamba model where the scanning consistency on spectral information may work.